### v1

# Scalable Generalized Bayesian Online Neural Network Training for Sequential Decision Making

Abstract
We introduce scalable algorithms for online learning and generalized Bayesian inference of neural network parameters, designed for sequential decision making tasks. Our methods combine the strengths of frequentist and Bayesian filtering, which include fast low-rank updates via a block-diagonal approximation of the parameter error covariance, and a well-defined posterior predictive distribution that we use for decision making. More precisely, our main method updates a low-rank error covariance for the hidden layers parameters, and a full-rank error covariance for the final layer parameters. Although this characterizes an improper posterior, we show that the resulting posterior predictive distribution is well-defined. Our methods update all network parameters online, with no need for replay buffers or offline retraining. We show, empirically, that our methods achieve a competitive tradeoff between speed and accuracy on (non-stationary) contextual bandit problems and Bayesian optimization problems.
1 Introduction
In sequential decision-making problems, such as Bayesian optimization and contextual bandits, uncertainty quantification and efficient belief updates are essential for balancing exploration and exploitation [15]. A standard approach to this is to use Bayesian neural networks (BNNs) [47]. However, exact posterior inference in BNNs, through methods such as Hamiltonian Monte Carlo (HMC), is computationally expensive, limiting its usage in time-sensitive applications like recommender systems, adaptive control, and financial forecasting [56, 43, 45, 8, 62].
A popular alternative to HMC is the use of variational Bayes methods [57, 6, 31]. Most work uses a Gaussian approximation to the posterior. However, this requires $O({D_{\bm{\theta}}}^{2})$ space to store the covariance (or precision matrix), and $O({D_{\bm{\theta}}}^{3})$ time to update it at each step, where ${D_{\bm{\theta}}}$ is the number of parameters in the network. One solution is to approximate the posterior precision matrix with a diagonal-plus-low-rank structure (see, e.g., the LoFi method [10]), which takes $O({D_{\bm{\theta}}}\,d^{2})$ time to update, where $d$ is the low-rank value. Another approach is to use last-layer Bayesian methods, that use a full covariance Gaussian over the final layer weights, but use a point estimate for the hidden layer weights (see e.g., the VBLL method [7]). These methods take $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}})$ time per update, where ${D_{{\boldsymbol{\ell}}}}$ is the number of parameters in the final layer and ${D_{{\bm{h}}}}$ is the number of parameters in the hidden layers. However, they do not capture uncertainty in the hidden layer parameters (which are usually the majority of the network). See Section 6 for further discussion of related work.
In this work, we propose a methodology that combines recursive frequentist inference, for efficient full-rank/low-rank parameter updates, with a generalized Bayesian posterior predictive that we use for sequential decision making. The frequentist part finds estimates of the parameters and quantifies the error variance of these estimates, which is simpler than characterizing densities. More precisely, we do not need to ensure that the updates we make yield a proper target density. However, following the principles of the Martingale posterior [22], we can still construct a valid posterior predictive distribution without invoking the prior-likelihood-posterior update. In summary, our probabilistic method to train neural networks enjoys these properties: (i) efficient closed-form updates (Monte Carlo sampling is not required); (ii) fully online (it avoids replay buffers, which can be problematic in non-stationary problems); (iii) delivers closed-form uncertainty-aware predictions through the posterior predictive.
We evaluate the performance of our methods across various sequential decision-making tasks. In non-stationary neural contextual bandits, our method achieves the highest performance with lowest computational cost compared to baselines. In stationary settings such as Bayesian optimization, our method matches the performance of replay-buffer approaches while achieving a range of Pareto-efficient tradeoffs between runtime and performance. Our code is implemented in Jax and it is publicly available at github.com/gerdm/low-rank-last-layer-inference
2 Problem statement
Consider a sequence of output measurements ${\bm{y}}_{1:T}=({\bm{y}}_{1},\ldots,{\bm{y}}_{T})$, with ${\bm{y}}_{t}\in^{D_{\bm{y}}}$, which are generated conditional on some input features ${\bm{x}}_{1:T}=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$, with ${\bm{x}}_{t}\in^{{D_{\bm{x}}}}$, for $t\in\{1,\dots,T\}$ and $T\in\mathbb{N}$. Let $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ be a datapoint at time $t$ and $\mathcal{D}_{1:T}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{T})$ the dataset at time $T$. We assume that ${\bm{y}}_{t}$, conditional on ${\bm{x}}_{t}$, is generated from a distribution that is characterized with unknown parameters ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$, i.e., ${\bm{y}}_{t}\sim p({f}({\bm{\theta}}_{t},{\bm{x}}_{t}))$, for some observation model $p$ with mean modelled by ${f}$.
[…]
Our goal is to efficiently compute best linear unbiased predictions (BLUPs) under a linearized model of ${f}$, along with an approximate block-diagonal error variance-covariance (EVC) structure that captures uncertainty separately for the hidden and last layer parameters. While we do not model or approximate the full Bayesian posterior $p({\bm{\theta}}_{t}\mid\mathcal{D}_{1:t})$, we treat the BLUP parameter estimates and their EVCs as defining degenerate Gaussian distributions [54], which we use to construct a proper posterior predictive $p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t})$ that supports probabilistic, sample-based sequential decision making. In this sense, we combine frequentist recursive parameter estimation with Bayesian decision making via the posterior predictive.
3 Background
3.1 Tools for sequential decision making
A key challenge in sequential decision making is the exploration-exploitation tradeoff, in which an agent must balance the use of its current knowledge of the environment (exploitation) with acquiring new information to improve future decisions (exploration) [4]. For example consider the case of a $K$-armed contextual bandit. Let the input (context) be ${\bm{x}}_{t}$, and the predicted mean reward for arm $a$ be ${f}({\bm{\theta}}_{t},{\bm{x}}_{t},a)$, where $a\in\{1,\ldots,K\}$. A simple approach — which we use as a baseline — is the $\epsilon$-greedy exploration, in which the next action is $a_{t+1}=\operatorname*{arg\,max}_{a}{f}(\hat{{\bm{\theta}}}_{t},{\bm{x}}_{t},a)$ with probability $1-\epsilon$, and uniformly at random with probability $\epsilon$, where $\epsilon\in(0,1)$ and $\hat{{\bm{\theta}}}_{t}$ is a point estimate.
A more sophisticated approach is known as Thompson sampling (TS) [52], where an agent samples the parameters from the posterior distribution (or approximate posterior [50]) and then makes a greedy decision based on a plugin estimate of the reward given the sampled parameter. This approach is studied in the context of Bayesian neural bandits in [51]. In this work, however, we focus on uncertainty in predictions rather than uncertainty in parameters. For example, for the case of the $K$-armed contextual bandit, our approach works as follows: at each timestep $t$, estimates of the reward are sampled from the posterior predictive $\hat{y}_{a}\sim p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},a,\mathcal{D}_{1:t})$, after which an action is selected as $a_{t+1}=\operatorname*{arg\,max}_{a\in\{1,\ldots,K\}}\hat{y}_{a}$.222 An alternative design is for ${f}$ to output a reward for each action, i.e., ${f}:\mathbb{R}^{D_{\bm{\theta}}}\times\mathbb{R}^{K\,{D_{\bm{x}}}}\to\mathbb{R}^{K}$. In this case, ${\bm{x}}_{t}\in\mathbb{R}^{K,{D_{\bm{x}}}}$ represents the concatenated features for all arms. Both configurations are explored in our experiments.
This predictive-based approach to TS is also considered in the context of Bayesian optimization in [7]. It is particularly relevant in our setting because our parameter estimates resemble degenerate multivariate Gaussians, due to the positive semi-definiteness of the covariance matrices [54]. Despite this, we show that the posterior predictive distribution is a valid density, (cf. [22]).
3.2 Extended Kalman filtering for online learning
We model observations as ${\bm{y}}_{t}={f}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1},{\bm{x}}_{t})+{\bm{e}}_{t}$, where ${f}$ is a neural network and ${\bm{e}}_{t}$ is zero-mean noise with covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$; see Appendix A for notation. Next, assume that the latent parameter vector ${\bm{\theta}}_{t}=[{\boldsymbol{\ell}}_{t}^{\intercal},{\bm{h}}_{t}^{\intercal}]^{\intercal}$ evolves as ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}+{\bm{u}}_{t}$ where ${\bm{u}}_{t}$ is a zero-mean dynamics noise process with variance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$. Following the frequentist interpretation of the Kalman filter [28, 21], we aim to recursively compute the best linear unbiased predictor (BLUP) ${\bm{\theta}}_{t|k}=\mathbf{A}_{t|k}{\bm{y}}_{1:k}$ and its associated error variance-covariance (EVC) matrix $\boldsymbol{\Sigma}_{t|k}={\rm Cov}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|k})$, for $k=t$ and $k=t-1$, under a local linearization of the model. Here, $\mathbf{A}_{t|k}\in{\mathcal{M}_{{D_{\bm{\theta}}}\times(k\,{D_{\bm{y}}})}(\mathbb{R})}$ is the best linear mapping (in $L^{2}$) from the collection of the first $k$ observations to the space of parameters at time $t$.
To obtain closed-form solutions for the BLUP and EVC at time $t$, we linearize the observation model ${f}$ around ${\bm{\theta}}_{t-1|t-1}=({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1})$ to obtain
${\bm{y}}_{t}\approx{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{L}}_{t}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t-1|t-1})+\tilde{\mathbf{H}}_{t}({\bm{h}}_{t}-{\bm{h}}_{t-1|t-1})+{\bm{e}}_{t},$ (1)
where $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$ and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$. It then follows that ${\bm{\theta}}_{t|t}$ and $\boldsymbol{\Sigma}_{t|t}$ can be written as a function of ${\bm{\theta}}_{t|t-1}$ and $\boldsymbol{\Sigma}_{t|t-1}$, which can be written as a function of ${\bm{\theta}}_{t-1|t-1}$ and $\boldsymbol{\Sigma}_{t-1|t-1}$ and so on. As a consequence, these updates enable efficient sequential (single-pass) online learning. This procedure is known as the extended Kalman filter (EKF).
[…]
4 Method
The standard EKF strategy for online learning presented in Section 3.2 does not typically scale to large neural networks because of the memory cost is $O({D_{\bm{\theta}}}^{2})$ and the compute cost is $O({D_{\bm{\theta}}}^{3})$. A number of approaches are proposed to tackle this issue (see Appendix 6 for details). Our method leverages the frequentist perspective of filtering to derive low-rank updates and the Bayesian interpretation of filtering to make sequential decisions through the posterior predictive.
4.1 Algorithm
In this section, we sketch our algorithm. For further details and full derivation, see Appendix D.1.
Consider the linearized model (1), with initial conditions
$\mathbb{E}[{\boldsymbol{\ell}}_{0}]$ $={\boldsymbol{\ell}}_{0|0},$ $\mathbb{E}[{\bm{h}}_{0}]$ $={\bm{h}}_{0|0},$ ${\rm Var}({\boldsymbol{\ell}}_{0})$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0},$ ${\rm Var}({\bm{h}}_{0})$ $=\mathbf{C}_{0}^{\intercal}\,\mathbf{C}_{0}.$ (2)
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. Next, assume that for $t\geq 1$, the last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\$ (3)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables with dynamics covariance matrices $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$.
Our method maintains an approximate belief state for the state-space model represented by the tuple $({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2})$, where ${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ is the BLUP estimate for model parameters in the last layer, ${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ is the BLUP estimate for model parameters in the hidden layers (feature extractor), $\hat{\boldsymbol{\Sigma}}_{t-1}^{1/2}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ is the Cholesky EVC factor for the covariance of the last layer, and $\mathbf{C}_{t}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ is the low-rank EVC factor for the hidden layers. Algorithm 1 shows a single step of HiLoFi.
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\mathbf{C}_{t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  // BLUP update step
15:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
16:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
17:  // EVC low-rank and Cholesky update step
18:  $\mathbf{C}_{t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
19:  $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\leftarrow{\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
20:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\right)$
Algorithm 1 Single step of HiLoFi for online learning. The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$; see Appendix C.2 for details.
The mean update step has the form ${\bm{\theta}}_{t|t}={\bm{\theta}}_{t|t-1}+\mathbf{K}_{t}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}={\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$ is the error or innovation term, and $\mathbf{K}_{t}$ is the Kalman gain matrix. This has roughly the same form as the EKF, except that $\mathbf{K}_{t}$ is computed separately for the last and hidden layers.
The EVC update however, is more complicated, since it requires the approximation of a block-diagonal covariance: the last layer block is full-rank, and the hidden layer blocks are low-rank. To maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers. The Cholesky factor is taken from a surrogate matrix that avoids computing a full ${\mathcal{M}_{{D_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ matrix by neglecting the artificial dynamics covariance $q_{{\boldsymbol{\ell}},t}$. The low-rank component for hidden layers is approximated via a two-step process: (i) a fast surrogate predicted covariance, and (ii) a rank ${d_{{\bm{h}}}}$ projection. Like the Cholesky factor, the surrogate covariance in (i) reduces the computational cost by ignoring the dynamics noise However, we retain some information from the dynamics noise to update the hidden layer parameters. Step (ii) enables a fast update rule.
After every update, the induced posterior predictive distribution is
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})={\cal N}\left({\bm{y}}_{t}\,|\,\tilde{\mathbf{L}}\,{\boldsymbol{\ell}}_{t|t-1}+\tilde{\mathbf{H}}\,{\bm{h}}_{t|t-1},\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\tilde{\mathbf{H}}^{\intercal}+\mathbf{R}_{t}\right),$ (4)
which is a proper density as long as $(\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\tilde{\mathbf{H}}^{\intercal}+\mathbf{R}_{t})\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. This can be easily ensured by taking $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. In (4), $\tilde{\mathbf{L}}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\tilde{\mathbf{L}}^{\intercal}+\tilde{\mathbf{H}}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\tilde{\mathbf{H}}^{\intercal}$ is the epistemic uncertainty and $\mathbf{R}_{t}$ is the aleatoric uncertainty.
The computational complexity of HiLoFi is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$. A full breakdown of the computational complexity and its relationship to other methods is in Appendix D.2.
Whenever the output layer or the input to the last layer is high-dimensional, we can perform a low-rank approximation to the last-layer to reduce computational costs. We call this variant LoLoFi. Its computational complexity is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. See Appendix E.2 for details.
A purely low-rank version of our method, which we call LRKF, is discussed in Appendix E.3. This method uses a single low rank covariance matrix approximation for all parameters in the neural network. In the experiments, we find that this approach has lower performance than HiLoFi, which is not surprising. It also has lower performance than LoLoFi (at least on the contextual bandit problem in section 5.2), presumably because modeling the last layer separately allows us to use a different subspace for the low rank approximation of the output layer weights, compared to use a shared subspace.
4.2 Theoretical properties
Proposition 4.1 (Covariance approximation error).
At time $t$, the per-step approximation error of HiLoFi under the linearized SSM (1) is bounded by
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq q_{{\bm{h}},t}\,\mathbf{E}_{{\bm{h}},{\rm surr}}+q_{{\boldsymbol{\ell}},t}\,\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}+\sqrt{\sum_{k={d_{{\bm{h}}}}+1}^{{D_{{\bm{h}}}}}\lambda_{k}^{2}}+2\|\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}^{2},$ (5)
where $\mathbf{E}_{{\bm{h}},{\rm surr}}=(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2})$, $\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}=\left\|\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\right\|_{\rm F}$, and $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are smallest $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Remark 4.2.
Proposition 4.1, whose proof is in Appendix F.4, shows that the per-step error in the HiLoFi approximation is bounded by: (i) the neglected terms by each surrogate matrix, (ii) the low-rank approximation for the covariance of the hidden layers, and (iii) the cross-terms from the blocks in the off-diagonal. (Note, however, that it does not account for error due to linearization.) One could minimize the above lower bound by setting $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$. However, this may result in lower per-step performance. See Appendix G.1.2 for an example of the main sources of error.
5 Experiments
In this section, we experimentally evaluate our method. We start by considering a 1d example of "in-between" uncertainty, to qualitatively illustrate the behavior of the method. We then quantitatively compare its performance to other methods across the following sequential decision-making tasks:
  1. (I)
    MNIST for classification and bandits: This is an online high-dimensional classification problem, which we also study as a contextual bandit problem.
  2. (II)
    Recommender systems: This is a challenging real-world problem with non-stationary data. Here, tackling exploration-exploitation tradeoff is key.
  3. (III)
    Bayesian optimization: The goal of this task is to find the maximum of an unknown blackbox function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$. This is a static inference problem where sample efficiency and uncertainty quantification are crucial.
For each of the above tasks, we assess the various methods on predictive performance and wall-clock time. All experiments were run on a TPU v4-8.
We consider the following variants of our framework. High-rank low-rank filter (HiLoFi): Our main method which uses full rank for last layer and low rank for hidden layers. Low-rank low-rank filter (LoLoFi): A version of our method that uses low-rank matrices for last and hidden layers. We only consider LoLoFi in task (II), due to the relatively large dimension of the last layer. Low-rank square-root Kalman Filter (LRKF): This is a special case of our method, described in Appendix E.3, that uses a single low-rank covariance for all the parameters. This serves to illustrate the gains from modeling the final layer separately.
As existing baselines, we consider the following. Diagonal plus low-rank precision (LoFi) [10]: A fully online method that models the precision matrix using a low-rank plus diagonal matrix, and uses the same linearization scheme as us. Gaussian processes (GP) [64]: A standard approach to modeling predictive uncertainty. We only include GPs in the low-dimensional tasks in (III) because of scalability limitations. Variational Bayesian last layer (VBLL) [7]: A partially Bayesian method that requires access to the full dataset (or a replay buffer), and performs multiple optimization steps per update. See Appendix E.4 for details. Online variational Bayesian last-layer (OnVBLL): A straightforward modification of the VBLL method with a first-in-first-out (FIFO) buffer and no regularization. Last-layer Laplace approximation with FIFO buffer (LLL) [13]: A method that uses a Laplace approximation on the last layer, with a FIFO replay buffer and multiple inner updates per time step. Unless otherwise stated, we train the parameters of the neural network using the AdamW optimization algorithm [44].
As an even simpler baseline, which does not model uncertainty, we consider Muon (muon) [32], which a special case of the Shampoo optimizer [26]. This is a quasi second-order optimization method that is scalable and shows strong empirical performance. We use the implementation in the Optax library [14]. Unlike the other methods, it only computes a point estimate, so we cannot use it for computing the posterior predictive. We use this method for the $\epsilon$-greedy bandit and as a choice of optimizer for LLL in Section 5.2.
5.1 In-between uncertainty
Problem description.
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [23] after a single pass of the data. We consider the one-dimensional dataset introduced in [61].
Model.
We use an MLP with 4 hidden layers and 128 units per layer, and we employ an ELU activation function.
Results.
Figure 1 shows the evolution of the posterior predictive mean and the posterior predictive variance as a function of the number of seen observations. We see that HiLoFi behaves in an intuitively sensible way, showing more uncertainty away from the data, just like a Gaussian process. See Appendix G.4 for the performance of other methods (specifically LRKF and VBLL) on this problem, as well an ablation study that illustrates the effect of changing rank.
Figure 1: In-between uncertainty induced by HiLoFi as we increase number of observations.
5.2 MNIST
Problem description.
We consider the standard MNIST classification problem, but in an online setting, where each image and its label is presented once, in a single pass over the data. For the classification problem, after predictions are made, the agent is revealed the true label of the data; we refer to this as the complete information case. We also consider the MNIST contextual bandit task introduced by [49], where the agent is presented with an image and must choose one of ten possible classes as an action, and then gets a binary feedback, based on whether its prediction was correct or not; we refer to this as the incomplete information case.
Model.
The predictive model (for the 10 label logits or the per-action rewards) is represented using a modified LeNet5 CNN architecture [39] with ELU activation function.
Exploration Methods.
For the bandit problem, we consider two exploration strategies: TS, which needs access to samples from an approximate posterior, and $\epsilon$-greedy, with $\epsilon=0.05$, which can leverage standard point estimation methods such as SGD or muon.
Inference methods.
We compare LLL, LRKF, LoFi, and HiLoFi. For LLL, we use the online last-layer variant from [13], without hyperparameter re-optimization at each timestep. In the $\epsilon$-greedy setting with LLL, we evaluate both online SGD with the AdamW optimiser [44] and the muon optimiser. Choice for hyperparameters are detailed in Appendix G.1.3.
Online learning results.
Figure 2 shows 5000-step rolling mean of the one-step-ahead classification outcome (one for correct, zero for incorrect) using the various learning algorithms. We observe that there is no clear difference between the methods. This shows that one can tackle this complete information problem without needing to model uncertainty. This is not the case in the incomplete information case that we study next, nor any of the other experiments below.
Figure 2: Rolling one-step-ahead accuracy for the MNIST dataset.
Bandit results.
Figure 3 shows the average cumulative reward over 10 runs of the contextual bandit version of MNIST. We observe that the top three performing methods are HiLoFi, LoLoFi, and LRKF. This outcome is expected because HiLoFi employs a full-rank covariance matrix in the last layer, which is better than a low-rank approximation of the covariance matrix associated with the last layer (LoLoFi), or a low-rank covariance to model all dependencies across layers (LRKF). For a detailed analysis of the trade-off between computation time and reward with error bars see Appendix G.1.3.
Figure 3: Cumulative reward for the bandit MNIST problem.
5.3 Recommender systems
Problem description.
We study the performance of the methods on the Kuairec dataset of [24], which is derived from a real-world, non-stationary, video recommender system. This dataset has been used to study non-stationary contextual bandits in [67].
Dataset.
We group rows in the dataset (for a given user) in blocks of the 10 next videos that the user saw. For the features, we consider the $\log(1+x)$ transform of like count, share count, play count, comment count, complete play count, follow count, reply comment count, and download count. For the target variable, we use the $\log(1+x)$ transform of the watch ratio $[0,\infty)$.
Model.
The reward model is a neural network with an embedding layer (one per video/arm), and a dense layer for all features. We join the embedding layer and the dense layer, and we then consider a three-hidden-layer neural network (50 units per layer, ELU activations), and linear output unit.
Algorithm.
We perform sequential Bayesian inference using the relevant method, and then use TS to choose the action at each step (see Section 3.1). The choice of hyperparameters is in G.2.
Results.
Figure 4 shows the average daily reward (left panel) and the running time of each method (right panel). We observe that HiLoFi has the highest average daily reward with second lowest running time. On average, LRKF and LoFi have similar performance, however, LoFi has a higher running time. Next, VBLL has higher average daily reward than OnVBLL, but it is slower. For an in-depth analysis of the results see Appendix G.2.
Figure 4: Recommender system results. Left: average daily reward. Right: running time.
5.4 Bayesian optimization
Problem description.
Following [7], we evaluate the competing methods on classical Bayesian optimization (BO) tasks, where the goal is to maximize an unknown function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$ using a surrogate model trained sequentially on past observations.
Model.
We employ a three-layer MLP (180 units per layer, ELU activations) as the surrogate for all neural-based methods. For the GP baseline, we use a Matérn-5/2 kernel with lengthscale 0.1 and a 20-point buffer to reduce computational cost. While GP performs better when its hyperparameters are tuned during training [29], this incurs much higher computational costs, so we do not pursue it here. The choice of hyperparameters is detailed in Appendix G.3.
Datasets.
We consider seven benchmark functions commonly used in BO [7]: Ackley (2D, 5D, 10D, 50D), Branin (2D), Hartmann (6D), and DrawNN (50D and 200D). These span a range of dimensionalities and multi-modalities, and are standard in evaluating global optimal performance. For all methods, except for DrawNN, we sample a function from the posterior predictive and evaluate it on a fixed set of candidate points generated using a Sobol sequence [55], in order to find its maximum. For DrawNN functions, we instead optimize the sampled function directly over the continuous domain $[0,1]$ using projected gradient descent, implemented with the Jaxopt library [5].
Algorithm.
We perform sequential Bayesian inference using the relevant method, and then choose the next query point at each step using TS (see Section 3.1). Results using expected improvement [30] are in Figure 11 in the Appendix.
Results.
Figure 5 (left) shows the final best value on the vertical axis (higher is better), and the amount of run time on the horizontal axis (lower is better). The number of steps for each dataset are in the x-axis in Figure 10 in Appendix G.3. Our method is competitive with other methods. Figure 5 (right) shows the tradeoff between performance (measured by rank) and compute time. Our method sits on the Pareto frontier for a range of tradeoffs between time and performance; furthermore, across tasks, the performance and compute time of HiLoFi dominates (better performance and less compute time) those from competing methods except from VBLL (higher performance) and LRKF (lower compute time). Thus, HiLoFi offers strong tradeoffs between compute time and performance. For further analysis of these results see Appendix G.3.
Figure 5: Left panel: Performance across all BO benchmark functions. For time, lower is better. For performance, higher is better. The dashed lines correspond to the results for our method, so methods that are above and to the left are better. Right panel: Time versus performance tradeoff plots.
6 Related work
Here we briefly discuss additional related work, beyond those in Section 1. We focus on deterministic methods for approximate online parameter inference. One line of work approximates the posterior using a diagonal plus low-rank (DLR) representation for the precision matrix, which allows for low-rank updates. This includes the L-RVGA method of [36], which applies a factor-analysis perspective to recursive updates, as well as SLANG [46] and LoFi [10]. Lastly, efficient low-rank Kalman filtering has been explored for state-space models in [53]. However, their method assumes that the dimension of the low-rank component is smaller than the dimension of the measurements, an assumption that is not required in our case. Going beyond low-rank approximations, the FOO-VB method [66] scales online BNN inference by considering a Kronecker product decomposition of the posterior (error-variance) covariance. In the offline setting, scaling BNNs using various low-rank assumptions has been studied in, e.g., [41, 59, 35].
Another class of methods draws on the lottery ticket hypothesis [38, 42], which suggests that sparse subspaces can yield performant models. Techniques such as the subspace neural bandit [17] and PULSE [9] restrict learning to a subset of weights represented as a linear subspace. However, these methods require offline learning of the projection matrix from the linear subspace to the space of weights before online inference begins.
7 Conclusion and future work
We introduced novel methods for efficient online training of neural networks. Our approach is based on linearization and optimal low-rank approximations of error covariance matrices. We proposed three variants: HiLoFi, LoLoFi, and LRKF. These variants offer a good trade-off between performance and runtime in static settings, and achieve highest performance and efficiency in high-dimensional, non-stationary scenarios.
One limitation of our methods is that it does not support fixed-lag smoothing, which can be important for sample efficiency in problems with low signal-to-noise ratio. In addition, our theory does not capture errors due to linearization. Additional limitations include sensitivity to outliers [1, 16, 37] and reliance on hyperparameters. In the future, we aim to extend our approach to fully-online reinforcement learning [20].
We view our contributions as primarily methodological, with no particular ethical concerns.
Acknowledgements
We thank the TPU Research Cloud program for providing the compute resources used in our experiments, and James Harrison for helpful comments.
Appendix A Notation
In what follows, ${\mathcal{M}_{m\times n}(\mathbb{R})}$ stands for the space of $m$-by-$n$ matrices with real values, ${\mathcal{M}^{++}_{m}(\mathbb{R})}$ is the space of positive definite (pd) matrices, and ${\mathcal{M}^{+}_{m}(\mathbb{R})}$ is the space of psd matrices. A vector ${\bm{a}}\in\mathbb{R}^{d}$ is a a column vector, $\mathbf{B}^{1/2}$ is the upper-triangular Cholesky decomposition of $\mathbf{B}\in{\mathcal{M}^{++}_{m}(\mathbb{R})}$, and $\mathbf{I}_{D}$ is the $D$-dimensional identity matrix with $D\in\mathbb{N}$.
The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization (in upper-triangular form) of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is itself an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation (in rectangular form) of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$, where each $\mathbf{A}_{k}\in\mathbb{R}^{d_{k}\times D}$ is a rectangular matrix such that $\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$ is symmetric and of rank, at most, $d_{k}$. That is, $\mathbf{C}^{\intercal}\,\mathbf{C}$ with $\mathbf{C}={\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ approximates the sum with a symmetric rank-$d$ matrix. Similarly, ${\cal P}_{d,+a}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns the best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}+a\,\mathbf{I}_{D}$, with $a>0$; see Appendix C.2 for details. The vector ${\bm{y}}_{1:k}$ is given by $[{\bm{y}}_{1}^{\intercal},\dots,{\bm{y}}_{k}^{\intercal}]^{\intercal}$, which is a vertically-stacked column vector. For vectors, $\|\cdot\|_{2}$ is the $L^{2}$ norm, and for matrices, $\|\cdot\|_{\rm F}$ is the Frobenius norm. Lastly, ${\cal N}({\bm{x}}\,|\,{\bm{m}},\mathbf{S})$ is a multivariate Gaussian density evaluated at ${\bm{x}}$ with mean ${\bm{m}}$ and covariance matrix $\mathbf{S}$.
Appendix B Extended Kalman filtering for online learning
Here, we explain in more detail the background of the EKF for online learning presented in Section 3.2. To make this section self-contained, we repeat parts of the material introduced in that section. Let us assume that ${\bm{y}}_{t}={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})+{\bm{e}}_{t}$, where ${\bm{e}}_{t}$ is a zero-mean random variable with observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. Given a starting vector ${\bm{\theta}}_{0}$, we assume that model parameters ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ and observations ${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ evolve according to the state-space model
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (6)
${\bm{y}}_{t}$ $={f}({\boldsymbol{\ell}}_{t-1},\,{\bm{h}}_{t-1},\,{\bm{x}}_{t})+{\bm{e}}_{t},$
where ${\bm{u}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ is a zero-mean random vector with known dynamics covariance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$, and ${\bm{e}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ is a zero-mean random vector with known observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. As before, we assume ${\rm Cov}({\bm{\theta}}_{0},{\bm{e}}_{t})=\boldsymbol{0}$ for all $t\in\mathbb{N}$ and ${\rm Cov}({\bm{u}}_{t},{\bm{e}}_{k})=\boldsymbol{0}$ for all $t,k\in\mathbb{N}$.
We now consider a first-order approximation of ${f}$ around ${\bm{\theta}}_{t-1|t-1}$, that is,
${\bm{y}}_{t}\approx{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t-1|t-1})+{\bm{e}}_{t}\,,$ (7)
where ${\bm{\theta}}_{t-1|t-1}=\begin{bmatrix}{\boldsymbol{\ell}}_{t-1|t-1}^{\intercal}&{\bm{h}}_{t-1|t-1}^{\intercal}\end{bmatrix}^{\intercal}$ is given, and $\tilde{\mathbf{J}}_{t}=\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$ with $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$ and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$.
Following a frequentist approach to the Kalman filter [28] and given the starting latent random vector ${\bm{\theta}}_{0|0}:={\bm{\theta}}_{0}$, we seek to obtain sequential updates for the best ($L^{2}$) linear estimate of the expected value of ${\bm{\theta}}_{t}$ given data ${\bm{y}}_{1:t}$. We formalize this in the following proposition.
Proposition B.1.
Let $k,t\in\mathbb{N}$ and let ${\bm{y}}_{t}$, ${\bm{\theta}}_{t}$ follow the SSM (7). The solution to the optimization problem
$\operatorname*{arg\,max}_{\mathbf{A}\in{\mathcal{M}_{({D_{{\boldsymbol{\ell}}}}+{D_{{\bm{h}}}})\times(k\,{D_{\bm{y}}})}(\mathbb{R})}}\mathbb{E}\left[||{\bm{\theta}}_{t}-\mathbf{A}\,{\bm{y}}_{1:k}||_{2}^{2}\right],$ (8)
is the matrix $\mathbf{A}_{t|k}^{\star}$ given by
$\mathbf{A}^{\star}_{t|k}={\rm Cov}({\bm{\theta}}_{t},{\bm{y}}_{1:k})\,{\rm Var}({\bm{y}}_{1:k})^{-1}.$ (9)
The best linear unbiased predictor (BLUP) for model parameters ${\bm{\theta}}_{t}$, given observations ${\bm{y}}_{1:k}$ is
${\bm{\theta}}_{t|k}$ $=\mathbf{A}_{t|k}^{\star}\,{\bm{y}}_{1:k},$ (10)
and the error variance covariance (EVC) matrix is defined as
$\boldsymbol{\Sigma}_{t|k}$ $:={\rm Var}({\bm{\theta}}_{t})-\mathbf{A}_{t|k}^{\star}\,{\rm Var}({\bm{y}}_{1:k})\,\mathbf{A}_{t|k}^{\star\intercal}.$ (11)
Proposition B.2.
Under the SSM (7), the BLUP and the EVC can be written in the form of Kalman filtering predict and update equations. Here, the predict equations are
${\bm{\theta}}_{t|t-1}$ $={\bm{\theta}}_{t-1|t-1},$ (12)
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t},$
and the update equations are
${\bm{\theta}}_{t|t}$ $={\bm{\theta}}_{t|t-1}+\mathbf{K}_{t}\,\boldsymbol{\epsilon}_{t},$ (13)
$\boldsymbol{\Sigma}_{t|t}$ $=(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})\,\boldsymbol{\Sigma}_{t|t-1}(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\mathbf{K}_{t}^{\intercal},$
with
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t}),$ (14)
$\mathbf{K}_{t}$ $=\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1},$
$\mathbf{S}_{t}$ $=\tilde{\mathbf{J}}_{t}\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}+\mathbf{R}_{t}.$
Proof.
The proof follows from the linear form of (7), Lemmas 2.1 to 2.3, and Theorem 4.2 in [21].
Remark B.3.
The update equations in (13) are recursive and characterize estimate of the unknown latent state and the estimated error estimation through the first two moments. Furthermore, the update for the covariance in (13) is in Joseph form [65] and is known to be numerically stable. In our method, we leverage these two facts to target a low-rank EVC through a second-step optimization that has numerically-stable updates.
Corollary B.4 (Kalman filter as a Bayesian posterior).
[…]
$p({\bm{\theta}}_{t}\,|\,{\bm{y}}_{1:t})={\cal N}({\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t|t},\boldsymbol{\Sigma}_{t|t}).$ (15)
As a consequence, the posterior predictive for the next observation in the sequence is
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})$ $=\int{\cal N}\left({\bm{y}}_{t}\,|\,{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1}),\mathbf{R}_{t}\right)p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}$ (16)
$={\cal N}({\bm{y}}_{t}\,|\,\tilde{\mathbf{J}}_{t}\,{\bm{\theta}}_{t|t-1},\tilde{\mathbf{J}}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\tilde{\mathbf{J}}_{t}^{\intercal}+\ \mathbf{R}_{t}).$
Appendix C Linear algebra results
C.1 Sum of Cholesky matrices
Here, we generalize the result found in [60] to any set of $K>2$ matrices, which we use to derive the method LRKF, HiLoFi, and LoLoFi.
Proposition C.1 (QR of sum of Cholesky matrices).
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,I\}$ be a collection of $(D\times D)$ positive definite matrices and let $\mathbf{A}_{i}^{1/2}$ be the upper-triangular Cholesky decomposition of $\mathbf{A}_{i}$. Let $\mathbf{M}$ be the stacked $(D\,I\times D)$ matrix given by
$\mathbf{M}=\begin{bmatrix}\mathbf{A}^{1/2}_{1}\\ \mathbf{A}^{1/2}_{2}\\ \vdots\\ \mathbf{A}^{1/2}_{I}\end{bmatrix}.$ (17)
It follows that
$\mathbf{R}^{\intercal}\,\mathbf{R}=\sum_{i=1}^{I}\mathbf{A}_{i}\,.$ (18)
where $\mathbf{R}$ is R component in the QR decomposition of $\mathbf{M}$.
Proof.
We see that
$\mathbf{M}^{\intercal}\,\mathbf{M}=\sum_{i=1}^{I}\mathbf{A}_{i}\,.$ (19)
Consider the QR decomposition of $\mathbf{M}$
$\mathbf{M}=\mathbf{Q}\mathbf{R},$ (20)
with $\mathbf{Q}$ an orthogonal matrix and $\mathbf{R}$ and upper-triangular matrix.
Then,
$\sum_{i=1}^{I}\mathbf{A}_{i}=\mathbf{M}^{\intercal}\,\mathbf{M}=\mathbf{R}^{\intercal}\,\mathbf{Q}^{\intercal}\,\mathbf{Q}\,\mathbf{R}=\mathbf{R}^{\intercal}\,\mathbf{R}.$ (21)
As a consequence, we note that the square root of the matrix $\mathbf{A}_{1}+\dots+\mathbf{A}_{I}$ is the upper-triangular matrix in the QR decomposition of $\mathbf{M}$. ∎
Remark
In what follows, we denote the result of obtaining the $\mathbf{R}$ matrix of the QR decomposition of the row-stacked matrices $\mathbf{A}^{1/2}_{i}$ for $i\in\{1,\dots,I\}$ by
${\cal Q}_{R}(\mathbf{A}^{1/2}_{1},\dots,\mathbf{A}^{1/2}_{I}).$ (22)
C.2 Singular value decomposition given sum of low-rank-matrices
Proposition C.2 (SVD of sum of low-rank matrices).
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,N\}$ be ${\mathcal{M}^{++}_{D}(\mathbb{R})}$ positive semi-definite matrices such that $\mathbf{A}_{i}=\mathbf{W}_{i}^{\intercal}\,\mathbf{W}_{i}$, where $\mathbf{W}_{i}\in\mathbb{R}^{d\times D}$ and $d\ll D$. The best rank-$d$ approximation of the sum is
$\sum_{i=N}^{I}\mathbf{A}_{i}\approx\mathbf{J}^{\intercal}\,\mathbf{J}\,,$ (23)
where $\mathbf{J}=\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D}$, $\mathbf{S}_{:d}$ are the top $d$ singular values and $\mathbf{V}$ contains the right singular vectors of the stacked $(d\,N\times D)$ matrix given by
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\ \mathbf{W}_{2}\\ \vdots\\ \mathbf{W}_{N}\end{bmatrix}.$ (24)
Proof.
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,N\}$ be ${\mathcal{M}^{++}_{D}(\mathbb{R})}$ matrices with $\mathbf{A}_{i}=\mathbf{W}_{i}^{\intercal}\mathbf{W}_{i}$, $\mathbf{W}_{i}\in\mathbb{R}^{d\times D}$, and $d\ll D$. Form the matrix $\mathbf{N}$ given by
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\[6.0pt] \mathbf{W}_{2}\\[3.0pt] \vdots\\[3.0pt] \mathbf{W}_{N}\end{bmatrix}=\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal},$ (25)
where $\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$ is its full singular value decomposition, with $\mathbf{U}\in\mathbb{R}^{dN\times dN}$, $\mathbf{S}\in\mathbb{R}^{dN\times D}$, and $\mathbf{V}\in\mathbb{R}^{D\times D}$. Then
$\sum_{i=1}^{N}\mathbf{A}_{i}$ $=\sum_{i=1}^{N}\mathbf{W}_{i}^{\intercal}\mathbf{W}_{i}=\mathbf{N}^{\intercal}\mathbf{N}$ (26)
$=(\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal})^{\intercal}\,(\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal})$
$=\mathbf{V}\,\mathbf{S}^{\intercal}\,\mathbf{U}^{\intercal}\,\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$
$=\mathbf{V}\,\mathbf{S}^{2}\,\mathbf{V}^{\intercal}.$
Hence $\sum_{i}\mathbf{A}_{i}$ has eigenvectors $\mathbf{V}$ and eigenvalues given by the diagonal entries of $\mathbf{S}^{2}$. By the symmetric-matrix form of the Eckart–Young theorem [19, 12], its best rank-$d$ approximation in Frobenius norm is
$\mathbf{V}\,{\rm diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2},0,\dots)\,\mathbf{V}^{\intercal}.$ (27)
Writing $\mathbf{S}_{:d}={\rm diag}(\sigma_{1},\dots,\sigma_{d})\in\mathbb{R}^{d\times D}$, one checks
$\mathbf{V}\,{\rm diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2},0,\dots)\,\mathbf{V}^{\intercal}=\bigl{(}\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\bigr{)}^{T}\,\bigl{(}\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\bigr{)}=\mathbf{J}^{\intercal}\,\mathbf{J},$ (28)
where
$\mathbf{J}$ $=\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D},$ (29)
$\mathbf{S}_{:d}$ $:={\rm diag}\left(\sigma_{1},\dots,\sigma_{d}\right),$
and $\{\sigma_{k}\}$ are the singular vectors of $\sum_{n=1}^{N}\mathbf{A}_{n}$. ∎
Takeaway.
Proposition C.2 shows that to form the best rank-$d$ approximation of $\sum_{i=1}^{N}\mathbf{A}_{i}$ one does not need to assemble or carry out the SVD of the full $D\times D$ sum. Instead, one stacks the low-rank factors into
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\ \vdots\\ \mathbf{W}_{N}\end{bmatrix}\in\mathbb{R}^{(dN)\times D},$ (30)
and compute a reduced SVD (or symmetric eigendecomposition) of the small $(dN)\times(dN)$ Gram matrix $\mathbf{N}\mathbf{N}^{\top}$. This costs
$\mathcal{O}((dN)^{2}D+(dN)^{3})\approx\mathcal{O}(d^{2}N^{2}D)\text{ when }dN\ll D,$ (31)
versus the $\mathcal{O}(D^{3})$ required for a full SVD on the $D\times D$ sum. Moreover, memory and computation scale with $d$ and $dN$ rather than with the large dimension $D$ because all operations involve only the $(dN)\times D$ matrix $\mathbf{N}$ and the $d\times D$ factor $\mathbf{J}$.
In what follows, we denote the result of obtaining the best rank-$d$ approximation matrix of the SVD decomposition of the row-stacked matrices $\mathbf{W}^{1/2}_{i}$ for $i\in\{1,\dots,N\}$ by
${\cal P}_{d}(\mathbf{W}_{1},\ldots,\mathbf{W}_{N}).$ (32)
Corollary C.3.
If the sum of matrices is of the form
$\sum_{n=1}^{N}\mathbf{A}_{n}+a\,\mathbf{I}_{D}$ (33)
with $a>0$, then the best rank-$d$ approximation is of the form
$\sum_{n=1}^{N}\mathbf{A}_{n}+a\,\mathbf{I}_{D}\approx\mathbf{J}_{+q}^{\intercal}\,\mathbf{J}_{+q}$ (34)
where
$\mathbf{J}_{+q}$ $=\mathbf{S}_{:d,+q}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D},$ (35)
$\mathbf{S}_{:d,+q}$ $:={\rm diag}\left(\sqrt{\sigma_{1}^{2}+q},\dots,\sqrt{\sigma_{d}^{2}+q}\right),$
and $\{\sigma_{k}\}_{k=1}^{d}$ are the top-$d$ singular values of $\sum_{n=1}^{N}\mathbf{A}_{n}$.
Proof.
From (26) in Proposition C.2 we obtain
$\sum_{i=1}^{N}\mathbf{A}_{i}+a\,\mathbf{I}=\mathbf{V}\,\mathbf{S}^{2}\,\mathbf{V}^{\intercal}+a\mathbf{I}=\mathbf{V}\,(\mathbf{S}^{2}+a\mathbf{I}_{D})\,\mathbf{V}^{\intercal}=\mathbf{V}\,\mathbf{S}_{+a}^{2}\,\mathbf{V}^{\intercal},$ (36)
where
$\mathbf{S}_{+a}^{2}={\rm diag}(\sigma_{1}^{2}+a,\ldots,\sigma_{D}^{2}+a).$ (37)
Let
$\mathbf{S}_{+a}={\rm diag}\left(\sqrt{\sigma_{1}^{2}+a},\ldots,\sqrt{\sigma_{D}^{2}+a}\right),$ (38)
then
$\sum_{n=1}^{N}\mathbf{A}_{n}$ $=\mathbf{V}\,\mathbf{S}_{+a}^{2}\,\mathbf{V}^{\intercal}$ (39)
$=(\mathbf{V}\,\mathbf{S}_{+a}\mathbf{V}^{\intercal})^{\intercal}\,(\mathbf{V}\,\mathbf{S}_{+a}\mathbf{V}^{\intercal})$
$=\mathbf{J}_{+a}^{\intercal}\,\mathbf{J}_{+a}.$
Takeaway.
If the sum of matrices can be written as the sum of multiplied low-rank factors plus an additional identity matrix times a constant, then, computation of the best rank $d$-matrix in low-rank factor can be obtained by performing SVD over the low-rank factors and modifying the singular values to include the term $a$.
In what follows, we denote the result of obtaining the best rank-$d$ approximation matrix of the SVD decomposition of the row-stacked matrices $\mathbf{W}^{1/2}_{i}$ for $i\in\{1,\dots,N\}$ plus an identity times a real-valued number $a>0$ as
${\cal P}_{d,+a}(\mathbf{W}_{1},\ldots,\mathbf{W}_{N}).$ (40)
Appendix D Further results on our method
D.1 Derivation
Consider the linearized model
${\bm{y}}_{t}={f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{L}}_{t}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t-1|t-1})+\tilde{\mathbf{H}}_{t}({\bm{h}}_{t}-{\bm{h}}_{t-1|t-1})+{\bm{e}}_{t},$ (41)
where ${\boldsymbol{\ell}}_{t-1|t-1}$ and ${\bm{h}}_{t-1|t-1}$ are given.
We start the algorithm by initialising the beliefs about the last layer parameters ${\boldsymbol{\ell}}$ and the hidden layer parameters ${\bm{h}}$. We set
$\mathbb{E}[{\boldsymbol{\ell}}_{0}]$ $={\boldsymbol{\ell}}_{0|0},$ $\mathbb{E}[{\bm{h}}_{0}]$ $={\bm{h}}_{0|0},$ ${\rm Var}({\boldsymbol{\ell}}_{0})$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0},$ ${\rm Var}({\bm{h}}_{0})$ $=\mathbf{C}_{0}^{\intercal}\,\mathbf{C}_{0},$ (42)
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. As mentioned above, the notation ${\bm{\theta}}_{t|k}$ (for $k,t\in\mathbb{N}$) stands for the best linear unbiased predictor (BLUP) of ${\bm{\theta}}_{t}$ given the information up to time $k$. Next, we assume that for $t\geq 1$, the latent last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\$ (43)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables333 i.e., ${\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},i},{\bm{u}}_{{\bm{h}},j})=\boldsymbol{0}$ for all $i,j\in\{1,\ldots,T\}$. with dynamics covariance matrices ${\rm Var}({\bm{u}}_{{\boldsymbol{\ell}},t})=\mathbf{Q}_{{\boldsymbol{\ell}},t}\in{\mathcal{M}^{+}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and ${\rm Var}({\bm{u}}_{{\bm{h}},t})=\mathbf{Q}_{{\bm{h}},t}\in{\mathcal{M}^{+}_{{D_{{\bm{h}}}}}(\mathbb{R})}$. For simplicity and to exploit efficient linear algebra techniques for belief updates, we assume $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$. A simple choice is to set $\mathbf{Q}_{{\bm{h}},t}=0\mathbf{I}$ to avoid forgetting past data; see e.g., [18].
Proposition D.1 (Predict step).
With the SSM assumption (43), the approximate covariance ${\rm Var}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and the approximate covariance ${\rm Var}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}$, the predict step becomes
${\bm{h}}_{t|t-1}$ $={\bm{h}}_{t-1|t-1},$ ${\boldsymbol{\ell}}_{t|t-1}$ $={\boldsymbol{\ell}}_{t-1|t-1},$ (44)
$\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},$ $\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}$ $=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{{\bm{h}},t}\,\mathbf{I}_{D_{{\bm{h}}}}.$
The proof of Proposition D.1 is in Appendix F.1.
Proposition D.2 (Variance of the innovation).
The upper Cholesky decomposition of innovation variance takes the form
$\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$ (45)
where $\tilde{\mathbf{L}}_{t}$ and $\tilde{\mathbf{H}}_{t}$ are defined in Section 3.2.
The proof of Proposition D.2 is in Appendix F.2.
Next, our update step for the BLUP resembles an EKF update, while the EVC update approximates a block-diagonal covariance: the last layer block is full-rank, and the hidden layer blocks are low-rank. To maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers. The Cholesky factor is taken from a surrogate matrix that avoids computing a full ${\mathcal{M}_{{D_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ matrix by neglecting the artificial dynamics covariance $q_{{\boldsymbol{\ell}},t}$. The low-rank component for hidden layers is approximated via a two-step process: (i) a fast surrogate predicted covariance, and (ii) a rank ${D_{{\bm{h}}}}$ projection error. Like the Cholesky factor, the surrogate covariance in (i) reduces the computational cost of dynamics noise, which does not reflect the system’s true dynamics. However, we retain some information from the dynamics noise to update all hidden layer parameters. Step (ii) enables a fast update rule, crucial for overparameterized neural networks [38]. We detail this step in the following proposition.
Proposition D.3 (Kalman gain, update BLUP, and update EVC).
Define the gain matrices
$\mathbf{K}_{{\bm{h}},t}^{\intercal}=\mathbf{V}_{{\bm{h}},t}\mathbf{C}_{t-1}\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\mathbf{V}_{{\bm{h}},t},\quad\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}=\mathbf{V}_{{\boldsymbol{\ell}},t}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\mathbf{V}_{{\boldsymbol{\ell}},t},$ (46)
where $\mathbf{V}_{{\bm{h}},t}=\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\tilde{\mathbf{H}}_{t},\quad$ and $\mathbf{V}_{{\boldsymbol{\ell}},t}=\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\tilde{\mathbf{L}}_{t}$. The updated BLUP for the last layer parameters and hidden layer parameters are
${\bm{h}}_{t|t}={\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\,\boldsymbol{\epsilon}_{t},\quad{\boldsymbol{\ell}}_{t|t}={\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\epsilon}_{t}.$ (47)
The approximate posterior covariance $\hat{\boldsymbol{\Sigma}}_{t|t}$ is the best block-diagonal approximation (in Frobenius norm) of the surrogate matrix $\tilde{\boldsymbol{\Sigma}}_{t|t}$ given by
$\scriptsize\left[\begin{smallmatrix}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\hat{\boldsymbol{\Sigma}}_{t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}-\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}-\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right],$
that has full rank for the last layer and rank ${d_{{\bm{h}}}}$ for the hidden layer. This results in a Cholesky factor for the last-layer covariance given by
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right),$ (48)
and a low-rank factor for the hidden layers given by
$\mathbf{C}_{t}={\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal},\;\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right),$ (49)
The proof is in Appendix F.3.
D.2 Computational complexity
We now compare the computational complexity of the algorithms we employ. Recall that ${D_{{\boldsymbol{\ell}}}}$, ${D_{{\bm{h}}}}$, and ${D_{\bm{y}}}$ are the number of parameters in the last layer, hidden layers, and the size of the output layer, respectively. As before, ${D_{\bm{\theta}}}={D_{{\boldsymbol{\ell}}}}+{D_{{\bm{h}}}}$. The dimensions ${d_{{\boldsymbol{\ell}}}},{d_{{\bm{h}}}}$, with ${d_{{\boldsymbol{\ell}}}}\ll{D_{{\boldsymbol{\ell}}}}$ and ${d_{{\bm{h}}}}\ll{D_{{\bm{h}}}}$, are the low-dimensional subspaces for the last and hidden layers, respectively, and $d={d_{{\boldsymbol{\ell}}}}+{d_{{\bm{h}}}}$.
The computational complexity of our algorithm when using full-rank in the last layer is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$. For low-dimensional outputs, this is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. For high-dimensional outputs, we may choose to use a low-rank approximation for the last layer (which we call LoLoFi) this reduces the cost to $O({D_{{\boldsymbol{\ell}}}}\,({d_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$, which is linear in ${D_{\bm{\theta}}}$ The primary computational bottlenecks in our method are the calculations for the Kalman gain of the hidden and last layers, with costs $O(2\,{D_{{\bm{h}}}}\,{D_{\bm{y}}}^{2})$ and $O(2\,{D_{{\boldsymbol{\ell}}}}\,{D_{\bm{y}}}^{2})$ respectively. This efficiency arises because $\mathbf{S}_{t}^{1/2}$ is upper triangular, allowing the system $\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2}\,\mathbf{V}=\tilde{\mathbf{J}}$ for $\mathbf{V}$ to be solved in $O(2\,{D_{{\bm{h}}}}\,{D_{\bm{y}}}^{2})$ as opposed to $O({D_{{\bm{h}}}}\,{D_{\bm{y}}}^{3})$. Additionally, approximating the covariance matrix via truncated SVD incurs a cost of $O({D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2}+({d_{{\bm{h}}}}+{D_{\bm{y}}})^{3}+{d_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})\,{D_{{\bm{h}}}})$, as detailed in Figure 8.6.1 in [25]. Finally, the corresponding cost for the last layer is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2})$.
Among related methods, the closest to ours in terms of computational costs are variational Bayes approaches such as the Slang method [46], the L-RVGA method [36], and the LoFi method [10]. In particular, LoFi uses the linearized Gaussian updates (similar to the ones in this paper), but approximates the precision matrix using a diagonal-plus-low rank (DLR) form. The appeal for DLR precision matrices is twofold. First, they enable the application of the Woodbury identity, leading to a predict step of cost $O({D_{\bm{\theta}}}\,d+d^{3})$ and an update step of $O({D_{\bm{\theta}}}\,(d+{D_{\bm{y}}})^{2})=O({D_{{\boldsymbol{\ell}}}}\,(d+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,(d+{D_{\bm{y}}})^{2})$. Second, incorporating positive diagonal terms ensures the matrix is positive definite, so that a valid posterior Gaussian density is defined.
Although HiLoFi and LoFi have the same asymptotic complexity, LoFi incurs additional practical overhead due to three key operations absent in HiLoFi: (1) the inversion of a $(d+{D_{\bm{y}}})$ rectangular matrix, (2) the inversion of a $d$ rectangular matrix, and (3) the Cholesky decomposition of a $d$ rectangular matrix. These operations increase the actual per-step computational time, which in our example, scales at about one second per additional rank. This behaviour is shown in the empirical comparison across HiLoFi, LRKF, and LoFi in the online classification setting (Figure 6, Appendix G.1).
Appendix E Algorithms
E.1 HiLoFi
See section D.1 for the derivation and algorithm 2 for the code.
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\mathbf{C}_{t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  // BLUP update step
15:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
16:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
17:  // EVC low-rank and Cholesky update step
18:  $\mathbf{C}_{t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
19:  $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\leftarrow{\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2},\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
20:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}\right)$
Algorithm 2 Single step of HiLoFi for online learning.
E.2 LoLoFi
Algorithm 3 shows a single step of LoLoFi. The low-rank factor for last layers $\mathbf{C}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}_{{d_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and the low-rank factor for the hidden layers $\mathbf{C}_{{\bm{h}},0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ are chosen at initialization.
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $(q_{{\boldsymbol{\ell}},t},q_{{\bm{h}},t})$ dynamics covariance for last layer and hidden layers
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $\left({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},\mathbf{C}_{{\boldsymbol{\ell}},t-1},\mathbf{C}_{{\bm{h}},t-1}\right)$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$
3:  $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
4:  $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$
5:  // innovation (one-step-ahead error) and Cholesky innovation variance
6:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
7:  $\mathbf{S}_{t}^{1/2}\leftarrow{\cal Q}_{R}\left({\mathbf{C}}_{{\boldsymbol{\ell}},t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{{\bm{h}},t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$
8:  // gain for hidden layers
9:  $\mathbf{V}_{{\bm{h}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{H}}_{t}$
10:  $\mathbf{K}_{{\bm{h}},t}^{\intercal}\leftarrow\mathbf{V}_{{\bm{h}},t}\,\mathbf{C}_{{\bm{h}},t-1}\,\mathbf{C}_{{\bm{h}},t-1}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{V}_{{\bm{h}},t}$
11:  // gain for last layer
12:  $\mathbf{V}_{{\boldsymbol{\ell}},t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\tilde{\mathbf{L}}_{t}$
13:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
14:  $\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\leftarrow\mathbf{V}_{{\boldsymbol{\ell}},t}\,\mathbf{C}_{{\boldsymbol{\ell}},t-1}\,\mathbf{C}_{{\boldsymbol{\ell}},t-1}^{\intercal}+q_{{\boldsymbol{\ell}},t}\,\mathbf{V}_{{\boldsymbol{\ell}},t}$
15:  // mean update step
16:  ${\bm{h}}_{t|t}\leftarrow{\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\boldsymbol{\epsilon}_{t}$
17:  ${\boldsymbol{\ell}}_{t|t}\leftarrow{\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\boldsymbol{\epsilon}_{t}$
18:  // low-rank updates
19:  ${\mathbf{C}}_{{\boldsymbol{\ell}},t}\leftarrow{\cal P}_{{d_{{\boldsymbol{\ell}}}},+q_{{\boldsymbol{\ell}},t}}\left(\mathbf{C}_{{\boldsymbol{\ell}},t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right)$
20:  ${\mathbf{C}}_{{\bm{h}},t}\leftarrow{\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{{\bm{h}},t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)$
21:  return  $\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{{\boldsymbol{\ell}},t},\mathbf{C}_{{\bm{h}},t}\right)$
Algorithm 3 Single step of LoLoFi for online learning for $t\geq 1$.
E.3 Low-rank Kalman filter (LRKF)
Inspired by [60], we derive a novel version of the low-rank Kalman filter which is (i) more computationally efficient, (ii) simple to implement, and (iii) provably optimal (in terms of Frobenius norm). We call this the LRKF method. Instead of keeping track of $\boldsymbol{\Sigma}_{t|t}$ or $\mathbf{S}_{t}$, we keep track of a low-rank factor for the covariance, and compute the Cholesky factor for the innovation variance $\mathbf{S}_{t}^{1/2}$. Algorithm 4 summarizes the predict and update steps for LRKF. Next we turn to the derivation.
We assume the model parameters ${\bm{\theta}}$ and the observations ${\bm{y}}$ follow
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (50)
${\bm{y}}_{t}$ $=\mathbf{H}_{t}\,{\bm{\theta}}_{t}+{\bm{e}}_{t},$
where ${\rm Var}({\bm{e}}_{t})=\mathbf{R}_{t}$ is known and $\mathbf{H}_{t}$ is the known projection matrix. Both, ${\bm{e}}_{t}$ and ${\bm{u}}_{t}$ are zero-mean random vectors.
In the Kalman filter update equations, the so-called posterior mean and covariances are given by
$\boldsymbol{\mu}_{t}$ $=\boldsymbol{\mu}_{t-1}+\mathbf{K}_{t}\,({\bm{y}}_{t}-\hat{{\bm{y}}}_{t})\,,$ (51)
$\boldsymbol{\Sigma}_{t}$ $={\rm Var}({\bm{\theta}}_{t}-\boldsymbol{\mu}_{t}),$
where $\boldsymbol{\mu}_{0}$ is given,
$\hat{{\bm{y}}}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\mu}_{t|t-1}=\mathbf{H}_{t}\,\boldsymbol{\mu}_{t-1|t-1},$ (52)
$\mathbf{K}_{t}$ $={\rm Cov}({\bm{\theta}}_{t},\boldsymbol{\epsilon}_{t}){\rm Var}(\boldsymbol{\epsilon}_{t})^{-1}=\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}\mathbf{S}_{t}^{-1},$
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-\mathbf{H}_{t}\,\boldsymbol{\mu}_{t|t-1},$
$\mathbf{S}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t},$
$\mathbf{R}_{t}$ $={\rm Var}({\bm{e}}_{t}).$
The memory requirements of the posterior covariance is $O(D^{2})$, with $D$ number of parameters. This makes it unfeasible to store for moderately-sized neural networks.
Thus, to reduce the computational cost, we maintain the best rank $d$ approximation of the covariance matrix by
$\boldsymbol{\Sigma}_{t|t}\approx\mathbf{W}_{t}^{\intercal}\,\mathbf{W}_{t},$ (53)
where $\mathbf{W}_{t}\in\mathbb{R}^{d\times D}$ is the best rank-$d$ approximation to $\boldsymbol{\Sigma}_{t}$ (in a Frobenius norm sense). Furthermore, to maintain a numerically stable method, we work with the Cholesky of the innovation variance. In effect, this is a variant square-root low-rank Kalman filter method. We explain the details of this method below.
The predict step equations are given by
$\boldsymbol{\mu}_{t|t-1}$ $=\boldsymbol{\mu}_{t-1},$ (54)
$\boldsymbol{\Sigma}_{t|t-1}$ $=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\mathbf{I}.$ (55)
Proposition E.1 (Kalman gain and innovations).
The variance $\mathbf{S}_{t}$ of the innovation is
$\mathbf{S}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$ (56)
$=\mathbf{H}_{t}\,(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\mathbf{I})\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\mathbf{H}_{t}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal}+q_{t}\,\mathbf{H}_{t}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\begin{bmatrix}\mathbf{H}_{t}\,\mathbf{C}_{t}^{\intercal}&\sqrt{q_{t}}\,\mathbf{H}_{t}&\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal}\\ \sqrt{q_{t}}\mathbf{H}_{t}^{\intercal}\\ \mathbf{R}_{t}^{1/2}\end{bmatrix}$
$=\mathbf{S}_{t}^{\intercal/2}\mathbf{S}_{t}^{1/2}\,,$
with $\mathbf{S}_{t}^{1/2}$ given by
$\mathbf{S}_{t}={\cal Q}_{R}(\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal},\,\sqrt{q_{t}}\,\mathbf{H}_{t}^{\intercal}\,,\,\mathbf{R}_{t}^{1/2}),$ (57)
where ${\cal Q}_{R}$ returns the $R$ matrix from the QR decomposition row-stacked arguments and $\mathbf{R}_{t}^{1/2}$ is the upper-triangular Cholesky decomposition of $\mathbf{R}_{t}$. The Kalman gain is given by
$\mathbf{K}_{t}$ $=\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1}$ (58)
$=\left(\mathbf{S}_{t}^{-1}\,\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\right)^{\intercal}$
$=\left(\mathbf{S}_{t}^{-1}\,\mathbf{S}_{t}^{-\intercal}\,\mathbf{H}_{t}\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{{D_{\bm{\theta}}}}\right)\right)^{\intercal}.$
Proposition E.2 (Update step).
The updated mean is
$\boldsymbol{\mu}_{t}=\boldsymbol{\mu}_{t|t-1}+\mathbf{K}_{t}({\bm{y}}_{t}-\hat{{\bm{y}}}_{t}).$ (59)
The update covariance $\boldsymbol{\Sigma}_{t|t}$, approximated through a two-step procedure: first, a surrogate covariance that ignores the cross-term effect of $\mathbf{K}_{t}\,\mathbf{H}_{t}$ on the artificial noise covariance $q_{t}$, and second, a best rank-$d$ approximation (in Frobenious norm) to the surrogate matrix takes the low-rank form
$\hat{\boldsymbol{\Sigma}}_{t|t}=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t},$ (60)
with
$\mathbf{C}_{t}={\cal P}_{d}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)\in{\mathcal{M}_{d\times{D_{\bm{\theta}}}}(\mathbb{R})},$ (61)
the low-rank (rectangular decomposition) matrix.
Proof.
The updated covariance mean follows directly from (13).
The updated covariance is
$\boldsymbol{\Sigma}_{t}$ $=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\boldsymbol{\Sigma}_{t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\mathbf{K}_{t}^{\intercal}$ (62)
$=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}\right)\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}$
$=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}\right)\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}$
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}.\end{bmatrix}+q_{t}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}.$
Next, we consider the surrogate covariance
$\tilde{\boldsymbol{\Sigma}}_{t}$ $=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}.\end{bmatrix}+q_{t}\mathbf{I}_{D_{\bm{\theta}}}$ (63)
$=\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{t}\mathbf{I}_{D_{\bm{\theta}}},$
where
$\tilde{\mathbf{C}}_{t}=\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\end{bmatrix}\in\mathbb{R}^{(d+o)\times D}.$ (64)
Next, the best rank-$d$ matrix is
$\hat{\boldsymbol{\Sigma}}_{t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ (65)
$=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$
$=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t},$
where
$\mathbf{C}_{t}={\cal P}_{d,+q_{t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)\in{\mathcal{M}_{d\times{D_{\bm{\theta}}}}(\mathbb{R})}$ (66)
is the best $d$-dimensional low-rank matrix given the stacked matrices $\tilde{\mathbf{C}}_{t}$ and the dynamics covariance $q_{t}$.
Proposition E.3.
The per-step error induced by the approximation of the covariance at time $t$ is bounded by
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq q_{t}\,\left(2\,\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}+\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}^{2}\right)+\sqrt{\sum_{k=d+1}^{D_{\bm{\theta}}}\lambda_{k}^{2}},$ (67)
where $\{\lambda_{k}\}_{k=d+1}^{D_{\bm{\theta}}}$ are the bottom $({D_{\bm{\theta}}}-d)$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Proof.
We seek to bound
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F},$ (68)
where $\boldsymbol{\Sigma}_{t|t}$ is given by (62) and $\hat{\boldsymbol{\Sigma}}_{t|t}$ is given by (65). We first note that
$\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}=\left(\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}\right)+\left(\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\right),$ (69)
where $\tilde{\boldsymbol{\Sigma}}_{t|t}$ is the surrogate covariance matrix (63). Then,
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq\|\underbrace{\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm surr}}\|_{\rm F}+\|\underbrace{\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm proj}}\|_{\rm F}$ (70)
The norm for $\mathbf{E}_{\rm proj}$ follows directly from the definition of the low-rank approximation. Let $\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$ be the SVD decomposition of $\tilde{\boldsymbol{\Sigma}}_{t|t}$ and let $\mathbf{U}\,\mathbf{S}_{:d}\,\mathbf{V}$ be the SVD decomposition of $\hat{\boldsymbol{\Sigma}}_{t|t}$. Here, $\mathbf{S}={\rm diag}(\lambda_{1},\ldots,\lambda_{D_{\bm{\theta}}})$ and $\mathbf{S}_{:d}={\rm diag}(\lambda_{1},\ldots,\lambda_{d},0,\ldots,0)$ are the singular values (eigenvalues) of the matrices $\tilde{\boldsymbol{\Sigma}}_{t|t}$ and $\hat{\boldsymbol{\Sigma}}_{t|t}$ respectively, ordered in descending order. Then,
$\|\mathbf{E}_{\rm proj}\|_{\rm F}$ $=\|\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}-\mathbf{U}\,\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\|_{\rm F}$ (71)
$=\|\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\|_{\rm F}$
$=\sqrt{{\rm Tr}[\left(\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\right)\left(\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\right)^{\intercal}]}$
$=\sqrt{{\rm Tr}[\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\mathbf{V}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{U}^{\intercal}]}$
$=\sqrt{{\rm Tr}[\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\mathbf{V}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{U}^{\intercal}\mathbf{U}]}$
$=\sqrt{{\rm Tr}[\left(\mathbf{S}-\mathbf{S}_{:d}\right)^{2}]}$
$=\sqrt{\sum_{k=d+1}^{D_{\bm{\theta}}}\lambda_{k}^{2}},$
where $\{\lambda_{k}\}_{k=d+1}^{D_{\bm{\theta}}}$ are the bottom $({D_{\bm{\theta}}}-d)$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Next, an upper bound for the norm $\mathbf{E}_{\rm surr}$ is as follows
$\|\mathbf{E}_{\rm surr}\|_{\rm F}$ $=\|\boldsymbol{\Sigma}_{t|t}-\tilde{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}$ (72)
$=\|q_{t}\,(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t})\,(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t})-q_{t}\,\mathbf{I}\|_{\rm F}$
$=q_{t}\,\|(\mathbf{K}_{t}\,\mathbf{H}_{t})\,(\mathbf{K}_{t}\,\mathbf{H}_{t})^{\intercal}-(\mathbf{K}_{t}\,\mathbf{H}_{t})-(\mathbf{K}_{t}\,\mathbf{H}_{t})^{\intercal}\|_{\rm F}$
$\leq q_{t}\left(\|(\mathbf{K}_{t}\,\mathbf{H}_{t})\|_{\rm F}^{2}+2\,\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}\right).$
0:  $\mathbf{R}_{t}$ // measurement variance
0:  $q_{t}$ // dynamics covariance
0:  $({\bm{y}}_{t},{\bm{x}}_{t})$ // observation and input
0:  $(\boldsymbol{\mu}_{t-1},\mathbf{C}_{t-1})$ // previous belief
1:  // predict step
2:  $\hat{{\bm{y}}}_{t}\leftarrow h({\bm{\theta}}_{t},{\bm{x}}_{t})$
3:  $\mathbf{H}_{t}\leftarrow\nabla_{\bm{\theta}}h(\boldsymbol{\mu}_{t-1},{\bm{x}}_{t})$
4:  // innovation and (Cholesky) innovation variance
5:  $\boldsymbol{\epsilon}_{t}\leftarrow{\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$
6:  $\mathbf{S}_{t}^{1/2}\leftarrow{\cal Q}_{R}(\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal},\sqrt{q_{t}}\,\mathbf{H}_{t},\,\mathbf{R}_{t}^{1/2})$
7:  // gain matrix
8:  $\mathbf{V}_{t}\leftarrow\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\mathbf{H}_{t}$
9:  $\mathbf{K}_{t}^{\intercal}\leftarrow\mathbf{V}_{t}\,\mathbf{C}_{t-1}\,\mathbf{C}_{t-1}^{\intercal}+q_{t}\,\mathbf{V}_{t}$ // low-rank update
10:  // mean and low-rank factor updates
11:  $\boldsymbol{\mu}_{t}\leftarrow\boldsymbol{\mu}_{t-1}+\mathbf{K}_{t}\,\boldsymbol{\epsilon}_{t}$
12:  $\mathbf{C}_{t}\leftarrow{\cal P}_{d}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)$
13:  return  $(\boldsymbol{\mu}_{t},\mathbf{C}_{t})$
Algorithm 4 Predict and update steps in the low-rank Kalman filter (LRKF)
E.4 Replay-buffer variational Bayesian last layer (OnVBLL)
The VBLL method has explicit posterior predictive
$p({\bm{y}}\,|\,\boldsymbol{\ell},{\bm{h}},\boldsymbol{\Sigma},\mathbf{R})={\cal N}\left({\bm{y}}\,|\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}}),\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}})\,\boldsymbol{\Sigma}\,\psi({\bm{h}},{\bm{x}})^{\intercal}\,\boldsymbol{\ell}+\mathbf{R}\right).$
Building on [27, 7], we observe that for any given $q(\bar{{\boldsymbol{\ell}}})={\cal N}(\bar{{\boldsymbol{\ell}}}\,|\,{\boldsymbol{\ell}},\boldsymbol{\Sigma})$, a lower-bound for the marginal log-likelihood of a single datapoint is given by
$\log p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},{\bm{h}},\mathbf{R})$ $\geq\mathbb{E}_{q(\bar{{\boldsymbol{\ell}}})}[\log p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\bar{{\boldsymbol{\ell}}},{\bm{h}},\mathbf{R})]$ (73)
$=\log{\cal N}\left({\bm{y}}_{t}\,|\,\boldsymbol{\ell}^{\intercal}\phi({\bm{h}},{\bm{x}}_{t}),\mathbf{R}\right)-\frac{1}{2}\phi({\bm{h}},{\bm{x}}_{t})\boldsymbol{\Sigma}\,\phi({\bm{h}},{\bm{x}}_{t})\operatorname{Tr}\left(\mathbf{R}^{-1}\right)$
$=:-{\cal L}_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma}).$
Following [34, 33, 31] a generalized posterior with loss function ${\cal L}$ is given by
$q_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})\propto q_{t-1}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})\,\exp(-{\cal L}_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})).$ (74)
Estimation of the posterior mean (MAP filtering) involves estimating
${\boldsymbol{\ell}}_{t},{\bm{h}}_{t},\mathbf{R}_{t},\boldsymbol{\Sigma}_{t}=\operatorname*{arg\,max}_{{\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma}}q_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})$ (75)
which can be done implicitly through adaptive optimization methods as shown in [3]. Next, to improve the performance, we consider a replay-buffer which has been shown to be much more efficient than doing fully-online SGD [2, 11, 40]. This assumption breaks with the fully-online assumption, but it is a good contender in sequential decision making problems in stationary environments.
Appendix F Proofs
F.1 Proof of Proposition D.1
Proof.
Following Proposition B.2, the predicted mean takes the form
${\bm{\theta}}_{t|t-1}={\bm{\theta}}_{t-1|t-1}=\begin{bmatrix}{\boldsymbol{\ell}}_{t-1|t-1}\\ {\bm{h}}_{t-1|t-1}\end{bmatrix}.$ (76)
Next, the predicted posterior covariance takes the form
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t},$ (77)
where
$\boldsymbol{\Sigma}_{t-1|t-1}$ (78)
$\quad={\rm Var}({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-1|t-1})$
$\quad=\begin{bmatrix}{\rm Var}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&{\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})\\ {\rm Cov}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1},{\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&{\rm Var}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1}).\end{bmatrix}$
$\quad=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1}\\ &{\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})\\ {\rm Cov}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1},{\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&\mathbf{W}_{t-1}^{\intercal}\,\mathbf{W}_{t-1}\end{bmatrix}.$
Next,
${\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})$ (79)
$\quad={\rm Cov}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1})$
$\quad={\rm Cov}\left({\boldsymbol{\ell}}_{0}+\sum_{\tau=1}^{t-1}{\bm{u}}_{{\boldsymbol{\ell}},\tau},\,{\bm{h}}_{0}+\sum_{\tau=1}^{t-1}{\bm{u}}_{{\bm{h}},\tau}\right)$
$\quad={\rm Cov}({\boldsymbol{\ell}}_{0},{\bm{h}}_{0})+\sum_{\tau=1}^{t-1}{\rm Cov}({\boldsymbol{\ell}}_{0},{\bm{u}}_{{\bm{h}},\tau})+\sum_{\tau=1}^{t-1}{\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},\tau},{\bm{h}}_{0})+\sum_{\tau=1}^{t-1}\sum_{\tau^{\prime}=1}^{-1}{\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},\tau},{\bm{u}}_{{\bm{h}},\tau^{\prime}})$
$\quad=\boldsymbol{0},$
which follows from the assumptions about the SSM. Then
$\boldsymbol{\Sigma}_{t-1|t-1}={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1},\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}).$ (80)
and
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t}$ (81)
$={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}+q_{\boldsymbol{\ell}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{\bm{h}}\,\mathbf{I}_{D_{{\bm{h}}}}).$
Given that $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}$ is approximated at $t-1$ through $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and $\boldsymbol{\Sigma}_{{\bm{h}},t-1|t-1}$ is approximated at $t-1$ through $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$, we obtain
$\boldsymbol{\Sigma}_{t|t-1,{\boldsymbol{\ell}}}$ $\approx\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}+q_{\boldsymbol{\ell}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},$ (82)
$\boldsymbol{\Sigma}_{t|t-1,{\bm{h}}}$ $\approx\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{\bm{h}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}}.$
F.2 Proof of Proposition D.2
Proof.
By definition of ${\bm{y}}_{t}$ and Proposition B.2, the innovation takes the form
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$ (83)
$=\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1})+{\bm{e}}_{t}$
$=\tilde{\mathbf{L}}_{t}\,({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})+\tilde{\mathbf{H}}_{t}\,({\bm{h}}_{t}-{\bm{h}}_{t|t-1})+{\bm{e}}_{t}.$
Next, the variance of the innovation is
$\mathbf{S}_{t}$ $={\rm Var}(\boldsymbol{\epsilon}_{t})$ (84)
$={\rm Var}\left(\tilde{\mathbf{L}}_{t}\,({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})+\tilde{\mathbf{H}}_{t}\,({\bm{h}}_{t}-{\bm{h}}_{t|t-1})+{\bm{e}}_{t}\right)$
$=\tilde{\mathbf{L}}_{t}\,{\rm Var}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,{\rm Var}({\bm{h}}_{t}-{\bm{h}}_{t|t-1})\tilde{\mathbf{H}}_{t}^{\intercal}+{\rm Var}({\bm{e}}_{t})$
$=\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\tilde{\mathbf{L}}_{t}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal}+q_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\,\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\tilde{\mathbf{H}}_{t}^{\intercal}+q_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\,\tilde{\mathbf{H}}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\begin{bmatrix}\tilde{\mathbf{L}}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{\intercal/2}&\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}\,\mathbf{C}_{t-1}^{\intercal}&\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}&\mathbf{R}^{\intercal/2}\,\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{L}}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\\ \sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal}\\ \mathbf{C}_{t-1}\,\tilde{\mathbf{H}}^{\intercal}\\ \sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal}\\ \mathbf{R}_{t}^{1/2}\end{bmatrix}$
$=\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2},$
with
$\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right).$ (85)
F.3 Proof of Proposition D.3
Proof.
First, by Proposition B.2, the Kalman gain matrix is given by
$\mathbf{K}_{t}=\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1}.$ (86)
Next, following Proposition D.1 and Proposition D.2, we rewrite the Kalman gain as
$\mathbf{K}_{t}$ $={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1})\begin{bmatrix}\tilde{\mathbf{L}}_{t}\\ \tilde{\mathbf{H}}_{t}\end{bmatrix}\mathbf{S}_{t}^{-1}$ (87)
$={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1})\begin{bmatrix}\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \tilde{\mathbf{H}}_{t}\,\mathbf{S}_{t}^{-1}\end{bmatrix}$
$=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}\,\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \end{bmatrix}$
$=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}}_{t}}\end{bmatrix},$
where
$\mathbf{K}_{{\boldsymbol{\ell}},t}$ $:=\Big{(}\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\Big{)}^{\intercal},$ (88)
$\mathbf{K}_{{\bm{h}},t}$ $=\Big{(}\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{H}}_{t}\,\mathbf{W}_{t-1}\,\mathbf{W}_{t-1}^{\intercal}\Big{)}^{\intercal}.$
From the above and Proposition D.1 the updated covariance matrix takes the form
$\boldsymbol{\Sigma}_{t|t}\\ =(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})\,\boldsymbol{\Sigma}_{t|t-1}\,(\mathbf{I}-\mathbf{K}_{t}\tilde{\mathbf{J}}_{t})^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\,\mathbf{K}_{t}^{\intercal},$ (89)
where
$\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t}$ (90)
$\quad=\begin{bmatrix}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}&\boldsymbol{0}\\ \boldsymbol{0}&\mathbf{I}_{D_{{\bm{h}}}}\end{bmatrix}-\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}},t}\\ \end{bmatrix}\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$
$\quad=\begin{bmatrix}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}&\boldsymbol{0}\\ \boldsymbol{0}&\mathbf{I}_{D_{{\bm{h}}}}\end{bmatrix}-\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{H}}_{t}\\ \mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\end{bmatrix}$
$\quad=\begin{bmatrix}\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{H}}_{t}\\ \mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\end{bmatrix},$
the predicted error covariance $\boldsymbol{\Sigma}_{t|t-1}$ is
$\boldsymbol{\Sigma}_{t|t-1}=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\end{bmatrix},$ (91)
and
$\mathbf{K}_{t}\,\mathbf{R}_{t}\,\mathbf{K}_{t}^{\intercal}$ $=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}},t}\end{bmatrix}\,\mathbf{R}_{t}\,\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}$ (92)
$=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \end{bmatrix}.$
Thus,
$\boldsymbol{\Sigma}_{t|t}$ (93)
$\quad={\tiny\begin{bmatrix}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}}.$
Next, we consider the surrogate covariance matrix (which modifies the block-diagonal entries)
$\tilde{\boldsymbol{\Sigma}}_{t|t}=$ (94)
$\left[\begin{smallmatrix}\tiny(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \tiny\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}+\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right].$
It then follows that
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t},\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}},\boldsymbol{\Sigma}_{{\bm{h}}}:{\rm rank}(\boldsymbol{\Sigma}_{{\bm{h}}})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{\Sigma}_{{\bm{h}}}\end{bmatrix}\right\|_{\rm F}^{2}$ (95)
$=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}},\boldsymbol{\Sigma}_{{\bm{h}}}:{\rm rank}(\boldsymbol{\Sigma}_{{\bm{h}}})=d}\|\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}-\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}}\|_{\rm F}^{2}+\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\boldsymbol{\Sigma}_{{\bm{h}}}\|_{\rm F}^{2}+2\,\|\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}^{2},$
with
$\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ $=(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal},$ (96)
$\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}.$
Then, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}=\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ and $\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ is given by the first ${d_{{\bm{h}}}}$ principal components of $\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$.
Next, we find the update rule for the Cholesky and low-rank factors. First, the posterior covariance for the last layer takes the form
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ $=\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}$ (97)
$=\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}$
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}^{\intercal/2}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}^{1/2}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\end{bmatrix},$
so that
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}={\cal Q}_{R}\left(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}^{1/2}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right).$ (98)
Second, the (surrogate) posterior covariance for the hidden layers (lower right block-diagonal) is
$\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}$ (99)
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}$
$=\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}.$
Finally, by Corollary C.3, the best rank-$d$ low-rank factor is
$\mathbf{C}_{t}={\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)\in{\mathcal{M}_{({d_{{\bm{h}}}}+{D_{\bm{y}}})\times{D_{{\bm{h}}}}}(\mathbb{R})}.$ (100)
So that
$\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}.$ (101)
F.4 Proof of Proposition 4.1
Proof.
We seek to bound
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}.$ (102)
Following the arguments from Proposition E.3, we obtain
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq\|\underbrace{\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm surr}}\|_{\rm F}+\|\underbrace{\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm proj}}\|_{\rm F},$ (103)
where
$\mathbf{E}_{\rm surr}=q_{{\bm{h}},t}\left(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2}\right)+q_{{\boldsymbol{\ell}},t}\,\left\|\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\right\|_{\rm F}$ (104)
and
$\mathbf{E}_{\rm proj}=\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\|_{\rm F}+2\,\|\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}$ (105)
Given that $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$ is the best rank-$d$ approximation of $\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$, it follows from a similar derivation to (71) that
$\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\|_{\rm F}=\sqrt{\sum_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}\lambda_{k}^{2}},$ (106)
where $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are bottom $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$. ∎
Appendix G Additional experimental details
G.1 MNIST for classification and sequential decision making
In this section, we provide more detail on the setup and results of the classification problem from Section 5.2.
The results in Section 5.2 consider the LeNet5 convolutional neural network (CNN) architecture [39]. In this architecture, the last layer is $80$-dimensional and the output layer is $10$-dimensional, which corresponds to each of the $10$ possible classes. Applying HiLoFi in this setting would need the storage and update of an $800\times 800$ covariance matrix and $800$-dimensional mean vector. Sequential update of the Cholesky representation of this covariance matrix, although feasible, takes most of the computational cost in a single step of HiLoFi. Given this, here we consider LoLoFi to offset the computational cost of the layer layer. Another reason to prefer LoLoFi over HiLoFi in this setting is that the rank of the last layer is $800$-dimensional, whereas the rank for the hidden layers is $50$ dimensional. This corresponds to an overparametrized linear regression, which has been shown to have some pathologies in the offline setting [63].
G.1.1 Mnist for online classification
In this experiment, we consider the problem online learning and classification of the MNIST dataset using the LeNet5 CNN. We use the optimizers detailed in Section 5.2. We take the data $\mathcal{D}_{1:T}$ with $T=60,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where ${\bm{x}}_{t}$ is a $(28\times 28\times 1)$ array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot-encoded vector such that $({\bm{y}}_{t})_{i}=1$ if ${\bm{x}}_{t}$ represents the digit $i$ and 00 otherwise. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$.
Having the BLUP over parameters ${\bm{\theta}}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$, the linearized model has mean and variance
${\bm{m}}({\bm{\theta}}_{t},{\bm{x}}_{t})$ $={\rm softmax}({f}({\bm{\theta}}_{t},{\bm{x}})),$ (107)
$\bar{{\bm{m}}}_{t}$ $={\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})+\nabla_{\bm{\theta}}{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}})({\bm{\theta}}-{\bm{m}}_{t-1}),$
$\bar{{\bm{s}}}_{t}$ $={\rm diag}({\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t}))+{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})\,{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})^{\intercal}+\varepsilon\,\mathbf{I}_{D_{\bm{y}}}.$
This moment-matched linearization corresponds, in the Bayesian setting, to a Gaussian likelihood whose first and second moments match that of a Multinomial distribution. This representation was introduced in the context of online learning in [48].
In (107), ${\rm diag}({\bm{v}})$ is a function that takes as input the vector ${\bm{v}}\in\mathbb{R}^{m}$ and outputs a diagonal matrix with entries ${\rm diag}({\bm{v}})_{i,j}={\bm{v}}_{i}\,\delta(i-j)$, and the term $\varepsilon>0$ is a small constant that ensures non-zero variance.
Figure 6 shows the one-step-ahead accuracy of the last $10,000$ images and the running time of processing all $60,000$ images for HiLoFi, LoLoFi, LRKF, and LoFi as a function of their rank. For LoLoFi, we fix the rank of the last layer to be $50$ and vary the rank of the hidden layers.
Figure 6: Results of MNIST for online classification. Left panel: accuracy as a function of rank. Right panel: running time as a function of rank.
We observe that for low-ranks up to dimension $20$, LoFi is faster than both LoLoFi and LRKF. However, the running time of LoFi increases much more rapidly (as a function of rank) than either LoLoFi and LRKF. This is a consequence of the computational costs associated with points (1,2,3) in Appendix D.2. Next, we observe that the running time of LoLoFi is sightly above LRKF for varying rank; this is because of the double approximation of the covariance matrix: one for the hidden layers and another one for the last layer. Finally, we observe that HiLoFi is the method with highest computational cost in this experiment. This is due the high-dimensionality of the last-layer, relative to the rest of the methods.
Lastly, we observe that all methods increase their performance as a function of rank, albeit marginally. The method that most benefit from an increase in performance in HiLoFi. We hypothesize that this is due to the total rank of the hidden layer, relative to the rank of the last layer, which in this experiment is $800$.
G.1.2 Error analysis for LRKF
The top panel of Figure 7 shows the rolling mean for the two sources of approximation error in the covariance matrix for the LRKF method (detailed in Proposition E.3), as well as the rolling one-step-ahead accuracy.
Figure 7: Top: upper bound error for the surrogate error and the low-rank approximation. Bottom: $5,000$-step rolling one-step-ahead accuracy for the MNIST dataset.
We observe that, on average, both sources of error decrease over time and the accuacy of LRKF improves over time. Here, the only sources of error are the surrogate error and the low-rank error. This is because LRKF does not distinguish between last-layer and hidden-layer parameters.
G.1.3 MNIST as multi-armed bandit
Here, we present further results from Section 5.2. More precisely, we study the multi-armed bandit approach to solving the MNIST classification problem.
Choice of hyperparameters.
The hyperparameters for this experiment were chosen as follows: we set $\varepsilon=10^{-4}$ for all methods. For Adamw and muon (used in $\epsilon$-greedy and in conjunction with LLL), we use a learning rate of $10^{-4}$. For Adamw, we take 5 inner iterations and a buffer size of one and for muon, we take 1 inner iteration and a buffer size of one. Next, HiLoFi considers a rank of $50$ for the hidden parameters, we take $q_{{\bm{h}},t}=10^{-6}$ and $q_{{\boldsymbol{\ell}}}=10^{-6}$. The initial covariances $\boldsymbol{\Sigma}_{{\bm{h}},0}$ and $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},0}$ are both initialized as identity times a factor of $10^{-1}$. Similarly, LoLoFi is initialized as with HiLoFi, but with rank in the last-layer to be $100$. For LRKF, we consider a rank of $50$, $\boldsymbol{\Sigma}_{0}$ is initialized as the identity matrix, and $q_{t}=10^{-6}$. Finally, for LoFi, we also considered rank $50$; then, following the experiments in [10], we considered the initial covariance to be $a\mathbf{I}$, with $a=\exp(-8)$ (this corresponds to low-rank comprised of a matrix of zeros and diagonal terms, all set to $a$); lastly, the dynamics covariance is $q_{t}=0$.
All initial model parameters from the neural network are shared across methods and trials.
Time-reward analysis.
Figure 8 shows the total running time to run the ten trials in the $x$-axis, the mean (relative to oracle) cumulative reward and two standard deviations in the $y$-axis. Here, we consider $\epsilon$-greedy variants for HiLoFi, LoLoFi, and LRKF. Next, TS with LLL and Adamw optimizer is denoted adamw-TS, TS with LLL and muon optimizer is denoted muon-TS. $\epsilon$-greedy with Adamw optimizer is denoted adamw-$\epsilon$, and $\epsilon$-greedy with muon optimizer is denoted muon-$\epsilon$.
Figure 8: Total running time for experiments ($x$-axis) versus mean relative cumulative reward ($y$-axis).
We observe Adamw and muon are the methods with lowest performance and lowest relative cumulative reward for either $\epsilon$-greedy or TS using LLL. Next, we observe that LRKF is faster than LoLoFi, and LoLoFi is faster than HiLoFi. This result is expected and is explained in Appendix D.2. We also observe that LoFi with $\epsilon$-greedy matches LoLoFi in time, LoLoFi outperforms in mean accuracy and has lower variance. We conjecture that this result is due to the explicit modelling of the last layer happening for LoLoFi.
Finally, we observe that for all methods, $\epsilon$-greedy is faster than TS. This is because $\epsilon$-greedy only requires the evaluation of the posterior predictive mean, whereas TS samples from the posterior predictive, which in turn requires building the posterior predictive covariance shown in (4).
G.2 Bandits as recommendation systems
Further results from Section 5.3
Choice of hyperparameters.
In this experiment, all agents shared a rank of $20$. The choice of hyperparameters for LLL, LoFi, VBLL, and OnVBLL were chosen following analysis of performance on the first $10,000$ observations of the datasets. In particular, we tried Bayesian optimization techniques, but found that manual tuning of hyperparameters on the first $10,000$ observations yielded higher overall performance.
For OnVBLL, we considered a buffer size of $10$, a regularization weight of $1.0$, and $50$ inner iterations. For VBLL, we considered a buffer size of $1000$, a regularization weight of $10^{-3}$, and $100$ inner iterations.
Next, for HiLoFi, we take $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0.0$ and set the initial covariances to be the identity. Next, for LRKF, we take $q_{t}=0$ and identity matrix as initial covariance.
Further results.
Figure 9 shows the daily cumulative reward for all methods. To produce this plot, we sum the cumulative rewards for all users at the end of each day and then perform a cumulative sum over the end-of-day reward.
Figure 9: Daily cumulative reward.
We observe that HiLoFi obtains the highest daily cumulative reward. The method with second highest daily cumulative reward is LRKF, closely followed by LoFi and VBLL. However, as shown in Figure 4. LRKF is more than an order of magnitude faster than VBLL. From Figure 9 and Figure 9, we observe that the top performing methods in this experiment are HiLoFi and LRKF. However, LRKF is slightly faster than HiLoFi relative to all other methods.
G.3 Bayesian optimization
Further results from Section 5.4
Choice of hyperparameters.
We fix hyperparameters across methods to ensure fairness and reproducibility. For rank-based methods, we use a rank of 50. Methods requiring buffers (e.g., LLL, OnVBLL) use a FIFO replay buffer of size 20, with 50 inner iterations per update. VBLL uses the full dataset and 100 iterations per step.
Where applicable, aleatoric uncertainty is discarded ($\mathbf{R}_{t}=0$) to isolate epistemic effects. We consider equal learning rates across optimizer-based methods (e.g., LLL, OnVBLL) at $10^{-4}$. In filtering-based methods (e.g., HiLoFi, LoFi), this learning rate corresponds to the initial hidden-layer covariance; we set the final-layer prior variance to 1, emphasising epistemic modelling in the output layer. Hyperparameters for VBLL (Wishart scale, regularization weight) follow prior work [7].
Per-step results.
Figure 10 shows the median performance and the interquartile range of the best value found by each method for all test datasets.
Figure 10: Bayesian optimization benchmark on stationary environments.
We observe that VBLL, OnVBLL, and HiLoFi are among the top performers for all datasets. Next, LLL is the least performant method for the low-dimensional datasets, but among the top performers for Nndraw50 and the top perfomer in Nndraw200.
Next, Figure 11 shows the median performance and the interquartile range of the best value found by each method for all test datasets using expected improvement.
Figure 11: Bayesian optimization benchmark on stationary environments. Query points are chosen using expected improvement.
G.4 In-between uncertainty
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [23] after a single pass of the data. We consider the one-dimensional dataset introduced in [61] (Figure 1). For this experiment, we consider a four hidden-layer MLP with 128 units per layer and ELU activation function.
Result for HiLoFi.
We take $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$, initial covariances for last and hidden layers to be identity times $1/2$, and $\mathbf{R}_{t}=0.0$, which corresponds to having no aleatoric uncertainty in the data generating process.
Figure 12 shows the posterior predicted mean surrounded by the two-standard deviation posterior predictive for various ranks in the hidden layer.
Figure 12: In-between uncertainty of the posterior predictive induced by HiLoFi as a function of rank.
We observe that a rank of $1$ is not able to capture any uncertainty around the region with no observations; however, as we increase the rank, we observe that the posterior predictive becomes less and less confident about the true value of the mean on regions without data.
Next, Figure 13 shows the evolution of the posterior predictive mean and the posterior predictive variance as a function of the number of seen observations.
Figure 13: In-between uncertainty of the posterior predictive induced by HiLoFi as a function of seen observations.
We observe that the uncertainty around the posterior predictive mean is wide (covering the limit from $-10$ to $10$) and starts to decrease after $10$ observations. By $30$ observations, the uncertainty is narrower over regions where data has been observed, and by $120$ observations, most of the posterior predictive uncertainty is on regions where no data has been observed.
Result for LRKF
We repeat the experiment above for LRKF. Figure 14 shows the posterior predicted mean surrounded by the two-standard deviation posterior predictive for various ranks.
Figure 14: In-between uncertainty of the posterior predictive induced by LRKF as a function of rank.
In contrast to HiLoFi, we observe that LRKF requires a much higher rank to capture a reasonable level of in-between uncertainty. The panel for rank $10$ is empty because the posterior predictive (with $\mathbf{R}_{t}=0$) is psd, so it is not defined.
An important distinction between LRKF and HiLoFi is the HiLoFi considers a full-rank covariance matrix in the last layer. Despite this, LRKF with rank $200$ (which shows some notion of uncertainty) requires around $3$ times more coefficients in the covariance than HiLoFi with rank $50$ in the hidden layers.
We show the evolution of the posterior predictive mean and variance with rank $200$ of LRKF as a function of seen observations in Figure 15.
Figure 15: In-between uncertainty of the posterior predictive induced by LRKF as a function of seen observations.
We see that the evolution of the posterior predictive resembles that of HiLoFi, but at the cost of $3$ times more number of coefficients in the covariance.
Result for VBLL.
We contrast the result of HiLoFi with the offline VBLL method. For VBLL, the data is presented all at once and we perform full-batch gradient descent using Adamw with learning rate $10^{-3}$. Figure 16 shows the posterior predictive mean and two standard deviations around the posterior mean, as well as the loss curve for various epochs.
We observe that VBLL takes around $10^{4}$ epochs to converge, whereas HiLoFi takes only one. We emphasize that each epoch in VBLL considers all datapoints at once, so the gradient update is more informative, whereas for HiLoFi, the method only gets to observe the data once and in a sequence. This highlights the efficiency of our method to produce approximate posterior predictives that can be used online.
Figure 16: In-between uncertainty of the posterior predictive induced by VBLL as a function of number of epochs.