### v2

# Martingale Posterior Neural Networks for Fast Sequential Decision Making

Abstract
We introduce scalable algorithms for online learning of neural network parameters and Bayesian sequential decision making. Unlike classical Bayesian neural networks, which induce predictive uncertainty through a posterior over model parameters, our methods adopt a predictive-first perspective based on martingale posteriors. In particular, we work directly with the one-step-ahead posterior predictive, which we parameterize with a neural network and update sequentially with incoming observations. This decouples Bayesian decision-making from parameter-space inference: we sample from the posterior predictive for decision making, and update the parameters of the posterior predictive via fast, frequentist Kalman-filter-like recursions. Our algorithms operate in a fully online, replay-free setting, providing principled uncertainty quantification without costly posterior sampling. Empirically, they achieve competitive performance–speed trade-offs in non-stationary contextual bandits and Bayesian optimization, offering 10–100 times faster inference than classical Thompson sampling while maintaining comparable or superior decision performance.
1 Introduction
In various sequential decision-making problems, such as Bayesian optimization and contextual bandits, uncertainty quantification and efficient belief updates are essential for balancing exploration and exploitation. A prominent Bayesian approach is Thompson sampling (TS) [52], which samples from the posterior over model parameters and evaluates the corresponding predictive function (e.g., reward or surrogate) for decision-making.
Various works combine Bayesian uncertainty with the expressiveness of neural networks for sequential decision making [5, 51]. These methods, often called Bayesian neural networks (BNNs), typically approximate the posterior over model parameters (e.g., via variational inference) and then use this posterior for downstream tasks. However, BNNs often rely on misspecified priors and likelihoods [33], which often degrade predictive performance [60] and result in slower inference than non-Bayesian alternatives [35].
These limitations are especially problematic in online settings where updates and decisions must be made quickly, such as in recommender systems, adaptive control, financial forecasting, and large-scale Bayesian optimization [55, 43, 45, 6, 59].
To address these problems, we present a class of algorithms for Bayesian sequential decision making inspired by the predictive-first philosophy of the martingale posterior framework [20, 26]. Our approach operates in a fully online setting: the one-step-ahead posterior predictive is parameterized by a neural network, updated directly from real observations via frequentist Kalman filter-style recursions, and sampled once per step for decision making. This avoids the need for costly posterior inference, while still supporting uncertainty and sampling-based sequential decision making.
The parameters defining the predictive density are updated with a single-pass frequentist Kalman filter-style methods, which resemble online natural gradient steps [48] and require no replay buffer (copies of past observations, which one typically uses for multiple inner-iterations). We explore three structured covariance strategies: LRKF applies low-rank updates to all layers; HiLoFi applies a full-rank update to the last layer and a low-rank update to the hidden layers; LoLoFi applies low-rank updates to both. HiLoFi and LoLoFi are inspired by last-layer Bayesian methods [25], and offer different tradeoffs between computational efficiency and expressiveness.
Decision making is performed by sampling directly from the posterior predictive, resulting in 10 to 100 times faster inference than that of classical TS, even when compared with structured posteriors that use diagonal plus low-rank covariances [8]. Our algorithms scale to million-parameter networks, maintain constant-time updates, and operate in fully online, streaming-compatible settings. In addition, our methodology supports any acquisition strategy, such as classical TS or expected improvement.
In summary, our probabilistic method for training neural networks enjoys the following properties: (i) efficient closed-form updates (no Monte Carlo sampling required); (ii) sample efficiency through low-rank Kalman filter-like updates which do not require the specification of a posterior density; (iii) closed-form uncertainty-aware predictions via the posterior predictive.
We evaluate the performance of our methods across a range of sequential decision-making tasks. In non-stationary neural contextual bandits, our methods achieve the highest performance at the lowest computational cost compared to that of baselines. In stationary settings such as Bayesian optimization, our approach matches the performance of replay-buffer methods while achieving Pareto-efficient tradeoffs between runtime and performance. Our code is available at https://github.com/gerdm/martingale-posterior-neural-networks.
A complete summary of the notation used throughout the paper is in Appendix A.
2 Problem statement
We consider a sequential-decision-making agent that at time $t$ observes a context-action pair ${\bm{x}}_{t}=({\bm{c}}_{t},{\bm{a}}_{t})$ with context ${\bm{c}}_{t}\in\mathbb{R}^{{D_{c}}}$ (possibly empty) and action ${\bm{a}}_{t}\in{\cal A}\subseteq\mathbb{R}^{D_{a}}$. The environment returns a reward (or observation) according to
${\bm{y}}_{t}\sim p_{\rm env}(\cdot\,|\,{\bm{x}}_{t}),\qquad{\bm{y}}_{t}\in\mathbb{R}^{{D_{\bm{y}}}},$ (1)
which depends only on the context-action pair ${\bm{x}}_{t}$. We use the notation $y_{t}$ when the outcome is scalar (${D_{\bm{y}}}=1$) and ${\bm{y}}_{t}$ when it is vector-valued (${D_{\bm{y}}}>1$).
This formulation covers, e.g., Bayesian optimization (${\bm{c}}_{t}$ empty and ${\bm{a}}_{t}$ the query location), multi-armed bandits (${\bm{c}}_{t}$ empty and ${\bm{a}}_{t}$ the chosen arm), and contextual bandits (${\bm{c}}_{t}$ non-empty and ${\bm{a}}_{t}$ the chosen arm). It also serves as a proxy representation for multi-armed bandits by letting the components of ${\bm{y}}_{t}$ encode per-arm signals or auxiliary outcomes.
We write $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ for a datapoint and $\mathcal{D}_{1:t}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{t}\}$ for the dataset. Here, we assume that $p_{\rm env}$ is unknown, but we approximate it through the observation model
${\bm{y}}_{t}=f({\bm{\theta}}_{t},{\bm{x}}_{t})+{\bm{e}}_{t},$ (2)
with ${\bm{e}}_{t}$ a zero-mean random vector with covariance matrix $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$, $f$ a neural network, and ${\bm{\theta}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ unknown time-varying model parameters.
The agent maintains a belief state ${\bm{b}}_{t}$ that summarizes its knowledge of the environment after observing $\mathcal{D}_{1:t}$. This belief defines a parametric approximation to the one-step-ahead posterior predictive
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1})\approx p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t}).$ (3)
Our objective is to maintain ${\bm{b}}_{t}$ via efficient, single-pass frequentist updates and to use the predictive $p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})$ for probabilistic, sample-based sequential decision making. In this sense, we combine frequentist recursive parameter estimation with Bayesian decision making via the posterior predictive.
3 Background
3.1 Tools for sequential decision making
A central challenge in sequential decision making is the exploration-exploitation tradeoff, where an agent balances whether to use its current knowledge of the environment (exploitation) or to acquire new information to improve future decisions (exploration) [3].
For example, when solving Bayesian neural contextual bandit problems [51] with the classical Thompson sampling (TS) algorithm [52], an agent samples a parameter vector from the posterior distribution (or an approximation [50]), evaluates the reward function under this sampled parameter for each action, and selects the action with the highest sampled reward.
Here, we propose a novel approach that avoids sampling from (high-dimensional) parameter posteriors often found in Bayesian neural networks. Instead, we work directly with the posterior predictive distribution defined by the current belief state ${\bm{b}}_{t}$. For each candidate action, we sample a possible outcome from (3) and then choose the action with the highest sampled reward. The belief ${\bm{b}}_{t}$ is subsequently updated via frequentist recursions (Section 3.2).
Table 1 highlights the key differences between our predictive sampling approach and classical TS. The essential distinction is whether uncertainty is represented in parameter space (TS) or directly in outcome space (our approach).
Step Predictive sampling (Our approach) Classical TS
Sample Posterior predictive at each action Posterior over model parameters
Evaluate N/A Sampled parameters on reward function
Select action Argmax over sampled rewards Argmax over sampled rewards
Get reward $y_{t+1}\sim p_{\rm env}(\cdot\mid{\bm{x}}_{t+1})$ $y_{t+1}\sim p_{\rm env}(\cdot\mid{\bm{x}}_{t+1})$
Update belief Frequentist update (e.g., Algorithm 2) Bayesian posterior update over parameters
Table 1: Comparison between Bayesian predictive sampling (our approach) and classical Thompson sampling. Differences highlighted in bold.
Algorithm 1 shows our predictive sampling procedure for contextual bandits. At each decision-making step, the agent samples rewards from the predictive distribution for each action, selects the action with the highest sampled reward, observes the reward from the environment, and updates its belief state.
[…]
4:for all ${\bm{a}}\in{\cal A}$ do
5:   $\hat{{\bm{x}}}\leftarrow({\bm{c}}_{t+1},\,{\bm{a}})$
[…]
7:end for
[…]
12:Return $({\bm{b}}_{t+1},y_{t+1})$ // belief and reward
Algorithm 1 Predictive sampling for sequential decision making in contextual bandits.
3.2 Extended Kalman filtering for online learning and sequential decision making
To maintain and update the belief state ${\bm{b}}_{t}$, we adapt ideas from the frequentist perspective of the extended Kalman filter (EKF). Here, ${\bm{b}}_{t}=({\bm{\theta}}_{t|t},\,\boldsymbol{\Sigma}_{t})$, ${\bm{\theta}}_{t|t}=\mathbf{A}_{t}\,{\bm{y}}_{1:t}$, $\mathbf{A}_{t}=\operatorname*{arg\,min}_{\mathbf{A}}\mathbb{E}[\|{\bm{\theta}}_{t}-\mathbf{A}\,{\bm{y}}_{1:t}\|_{2}^{2}]$ is the best linear unbiased predictor (BLUP) of ${\bm{\theta}}_{t}$, and $\boldsymbol{\Sigma}_{t}={\rm Var}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|t})$ is the error-variance-covariance (EVC) matrix of the estimate. The initial belief state ${\bm{b}}_{0}=({\bm{\theta}}_{0|0},\,\boldsymbol{\Sigma}_{0})$ is given.
Next, we partition the neural network parameters into last-layer weights and hidden-layer weights: ${\bm{\theta}}_{t}=({\boldsymbol{\ell}}_{t},{\bm{h}}_{t})$, with ${\boldsymbol{\ell}}_{t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ (last layer) and ${\bm{h}}_{t}\in\mathbb{R}^{D_{{\bm{h}}}}$ (hidden layers) so that ${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t})$, and ${f}({\boldsymbol{\ell}}_{t},{\bm{h}}_{t},{\bm{x}}_{t})={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})\in\mathbb{R}$, where $\phi({\bm{x}}_{t},{\bm{h}}_{t})$ is the hidden-layer feature map of the neural network.
To obtain closed-form updates, we linearize the neural network around the previous parameter estimate ${\bm{\theta}}_{t-1|t-1}=({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1})$, and make a random-walk assumption for the model parameters ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}+{\bm{u}}_{t}$, with ${\rm Var}({\bm{u}}_{t})=\mathbf{Q}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{\theta}}}}(\mathbb{R})}$. This assumption prevents uncertainty collapse, adaptation in non-stationary environments, and numerical stability (see e.g., [46] or Section 2.6.4 in [13]).
With the above assumptions, we obtain the linearized state-space model
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (4)
${\bm{y}}_{t}$ $={f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{L}}_{t}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t-1|t-1})+\tilde{\mathbf{H}}_{t}({\bm{h}}_{t}-{\bm{h}}_{t-1|t-1})+{\bm{e}}_{t},$ (5)
where ${\bm{e}}_{t}$ is a zero-mean random variable with ${\rm Var}[{\bm{e}}_{t}]=\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$, $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})\in{\mathcal{M}_{1\times 2}(\mathbb{R})}$, and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$ are the Jacobians of the neural network w.r.t. the last and hidden layers respectively.
Updates.
The state-space model defined by (4)-(5) allows for recursive estimation of ${\bm{b}}_{t}$ via Kalman-like updates. See Appendix B for details.
Posterior predictive model.
At each timestep the one-step-ahead posterior predictive induced by the state-space model (4)-(5), given the belief state ${\bm{b}}_{t}$ and under the assumption $p({\bm{e}}_{t})={\cal N}({\bm{e}}_{t}\mid\boldsymbol{0},\,\mathbf{R}_{t})$ is
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})=\mathcal{N}\,\left({\bm{y}}_{t+1}\,\middle|\,{f}({\bm{\theta}}_{t|t},{\bm{x}}_{t+1}),\;\underbrace{\tilde{\mathbf{J}}_{t+1}\,\boldsymbol{\Sigma}_{t}\,\tilde{\mathbf{J}}_{t+1}^{\intercal}}_{\text{epistemic}}+\underbrace{\mathbf{R}_{t}}_{\text{aleatoric}}\right),$ (6)
where $\tilde{\mathbf{J}}_{t}=\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$. Here, the covariance of the posterior predictive decomposes into epistemic uncertainty (from parameter uncertainty under the linearization) and aleatoric uncertainty (from observation noise $\mathbf{R}_{t}$). The condition $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$ ensures that (6) defines a proper Gaussian density.
3.3 Martingale posteriors
Our notion of Bayesian uncertainty is inspired by the martingale posterior framework [20], which interprets Bayesian uncertainty through missing data [26]. In the original setting, one imagines imputing an infinite sequence of future observations; statistics computed on these extended datasets form a martingale, and under mild conditions they converge.
However, in sequential decision making, however, given the context-action pair ${\bm{x}}_{t+1}$, the only missing data, is the next observation ${\bm{y}}_{t+1}$ from the environment. In our case, uncertainty is quantified to guide decisions: ${\bm{y}}_{t+1}$ (or a partial observation of it, in the bandit setting) is always revealed before the next step. Thus, rather than simulating an infinite sequence of future steps, we focus entirely on the one-step-ahead posterior predictive $p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\mid{\bm{x}}_{t+1})$.
4 Method
The standard EKF strategy for online learning presented in Section 3.2 does not scale to large neural networks because the memory cost is $O({D_{\bm{\theta}}}^{2})$ and the compute cost is $O({D_{\bm{\theta}}}^{3})$. A number of approaches have been proposed to tackle this issue (see Section 6 for details). Our method leverages the frequentist perspective of filtering to derive low-rank updates and the Bayesian interpretation of filtering to make sequential decisions through the posterior predictive.
4.1 Algorithm
Consider the linearized model (5), with initial conditions
$\mathbb{E}[{\boldsymbol{\ell}}_{0}]$ $={\boldsymbol{\ell}}_{0|0},$ $\mathbb{E}[{\bm{h}}_{0}]$ $={\bm{h}}_{0|0},$ ${\rm Var}({\boldsymbol{\ell}}_{0})$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0},$ ${\rm Var}({\bm{h}}_{0})$ $=\mathbf{C}_{0}^{\intercal}\,\mathbf{C}_{0},$ (7)
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. Next, assume that for $t\geq 1$, the last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\\$ (8)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables with covariance matrices $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$, with $q_{{\boldsymbol{\ell}},t}\geq 0$ and $q_{{\bm{h}},t}\geq 0$. Our method maintains a belief state for the state-space model represented by ${\bm{b}}_{t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t})$, where ${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ is the estimate for model parameters in the last layer, ${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ is the estimate for model parameters in the hidden layers, $\hat{\boldsymbol{\Sigma}}_{t-1}^{1/2}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ is the Cholesky EVC factor for the covariance of the last layer, and $\mathbf{C}_{t}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ is the low-rank EVC factor for the hidden layers.
Updates.
Given ${\bm{b}}_{t-1}$, the updated ${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t})$ follows directly from the EKF equations which takes the form ${\bm{\theta}}_{t|t}={\bm{\theta}}_{t-1|t-1}+\mathbf{K}_{t}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}={\bm{y}}_{t}-\hat{{\bm{y}}}_{t}$ is the error or innovation term, and $\mathbf{K}_{t}$ is the gain matrix. Because of the block-diagonal assumption, it can be shown that $\mathbf{K}_{t}$ is decoupled into last-layer-only and a hidden-layers-only submatrices.
Next, applying the EKF EVC update to ${\bm{b}}_{t-1}$ yields the dense matrix $\boldsymbol{\Sigma}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$ which couples the effects from the last and hidden layers. To obtain a block-diagonal matrix with components $(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t})$, we consider a two-stage approximation. The first stage replaces $\boldsymbol{\Sigma}_{t}$ with a surrogate covariance matrix $\tilde{\boldsymbol{\Sigma}}_{t}$ that avoids the interaction between the hidden layer parameters and the synthetic dynamics $q_{{\bm{h}},t}$; the second stage approximates the surrogate matrix with the best block-diagonal approximation whose first block (for last-layer parameters) are of rank $D_{{\boldsymbol{\ell}}}$} and the second block (for hidden-layer parameters) are rank-$d_{{\bm{h}}}$}. Finally, to maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers, i.e., $\hat{\boldsymbol{\Sigma}}_{t}^{1/2}={\rm diag}(\mathbf{C}_{t},\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2})$. In this sense, estimation of the factors for the last-layer and the hidden layer follows the sequence
$\boldsymbol{\Sigma}_{t}\to\tilde{\boldsymbol{\Sigma}}_{t}\to\hat{\boldsymbol{\Sigma}}_{t}\to\hat{\boldsymbol{\Sigma}}_{t}^{1/2}={\rm diag}(\mathbf{C}_{t},\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}).$
Algorithm 2 shows a single step of HiLoFi. The derivation is in Appendix D.1.
[…]
5:// predict step
[…]
9:// innovation (one-step-ahead error) and Cholesky innovation variance
[…]
12:// gain for hidden layers
[…]
15:// gain for last layer
[…]
18:// mean update step
[…]
21:// EVC low-rank and Cholesky update step
[…]
24:Return ${\bm{b}}_{t}=\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2},\mathbf{C}_{t}\right)$ // updated belief
Algorithm 2 Single update step of HiLoFi. The notation ${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ denotes the Cholesky factorization of the matrix $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor; see Appendix C.1 for details. The function ${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})$ returns a $d\times D$ matrix which is the best rank-$d$ approximation of the matrix $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$; see Appendix C.2 for details.
The following proposition shows the upper bound of the the per-step error incurred by HiLoFi.
Proposition 4.1 (Covariance approximation error).
At time $t$, the per-step approximation error of HiLoFi under the linearized SSM (5) is bounded by
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq q_{{\bm{h}},t}\,\mathbf{E}_{{\bm{h}},{\rm surr}}+q_{{\boldsymbol{\ell}},t}\,\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}+\sqrt{\sum_{k={d_{{\bm{h}}}}+1}^{{D_{{\bm{h}}}}}\lambda_{k}^{2}}+2\|\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}^{2},$ (9)
where $\mathbf{E}_{{\bm{h}},{\rm surr}}=(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2})$, $\mathbf{E}_{{\boldsymbol{\ell}},{\rm surr}}=\|(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\,(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}\|_{\rm F}$, and $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are smallest $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Remark 4.2.
Proposition 4.1 shows that the per-step error in the HiLoFi approximation is bounded by: (i) the neglected terms by each surrogate matrix, (ii) the low-rank approximation for the covariance of the hidden layers, and (iii) the cross-terms from the blocks in the off-diagonal Note, however, that the bound does not account for error due to linearization of the neural network. One could minimize the above lower bound by setting $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$. However, this may result in lower per-step performance. See Appendix G.1.2 for an example of the main sources of error. See Appendix E.4 for a proof.
Posterior predictive model.
After every update, the posterior predictive model used for sequential decision making is
$p_{{\bm{b}}_{t}}({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1})={\cal N}\left({\bm{y}}_{t}\mid{f}({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},{\bm{x}}_{t+1}),\;\;\tilde{\mathbf{L}}_{t+1}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t+1}^{\intercal}+\tilde{\mathbf{H}}_{t+1}\,\boldsymbol{\Sigma}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t+1}^{\intercal}+\mathbf{R}_{t+1}\right).$ (10)
The computational complexity of HiLoFi is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$. A full breakdown of the computational complexity and its relationship to other methods is in Appendix D.2.
Variants.
Whenever the output layer or the input to the last layer is high dimensional, we perform a low-rank approximation to the last-layer to reduce computational costs. We call this variant LoLoFi. Its computational complexity is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. See Appendix F.1 for details.
A purely low-rank version of our method, which we call LRKF, is discussed in Appendix F.2. This method uses a single low rank covariance matrix approximation for all parameters in the neural network. An error analysis of LRKF in the linear setting is presented in Appendix F.2.1. In the experiments, we find that this approach performs worse than that of HiLoFi, which is not surprising. It also has lower performance than LoLoFi (at least on the contextual bandit problem in section˜5.2), presumably because modeling the last layer separately allows us to use a different subspace for the low rank approximation of the output layer weights.
5 Experiments
In this section, we evaluate our method. First, we consider a one-dimensional example “in-between” uncertainty, to illustrate the behavior of the method. We then compare its performance to that of other methods across the following sequential decision-making tasks:
  1. (I)
    MNIST for classification as a bandit problem: This is an online high-dimensional classification problem, which we study as a contextual bandit problem.
  2. (II)
    Recommender systems: This is a challenging real-world problem with non-stationary data. Here, tackling exploration-exploitation tradeoff is key.
  3. (III)
    Bayesian optimization: The goal of this task is to find the maximum of an unknown blackbox function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$. This is a static inference problem where sample efficiency and uncertainty quantification are crucial.
For each of the above tasks, we assess the various methods on predictive performance and wall-clock time. All experiments were run on a TPU v4-8. An additional experiment demonstrating the scalability of LRKF to multi-million-parameter neural networks for online classification is provided in Appendix G.2.
We consider the following variants of our framework. High-rank low-rank filter (HiLoFi): Our main method that uses full rank for last layer and low rank for hidden layers. Low-rank low-rank filter (LoLoFi): A version of our method that uses low-rank matrices for last and hidden layers. We only consider LoLoFi in task (II), due to the relatively large dimension of the last layer. Low-rank square-root Kalman Filter (LRKF): This is a special case of our method, described in Appendix F.2, that uses a single low-rank covariance for all the parameters. This illustrates the gains from modeling the final layer separately.
As existing baselines, we consider the following. Diagonal plus low-rank precision (LoFi) [8]: a fully online method that models the precision matrix using a low-rank plus diagonal matrix, and uses the same linearization scheme as us. Gaussian processes (GP) [62]: a standard approach to modeling predictive uncertainty. We only include GPs in the low-dimensional tasks in (III) because of scalability limitations. Variational Bayesian last layer (VBLL) [5]: a partially Bayesian method that requires access to the full dataset (or a replay buffer), and performs multiple optimization steps per update. See Appendix F.3 for details. Online variational Bayesian last-layer (OnVBLL): a straightforward modification of the VBLL method with a first-in-first-out (FIFO) buffer and no regularization. Last-layer Laplace approximation with FIFO buffer (LLL) [11]: a method that uses a Laplace approximation on the last layer, with a FIFO replay buffer and multiple inner updates per time step. Unless otherwise stated, we train the parameters of the neural network with the AdamW optimization algorithm [44].
5.1 In-between uncertainty
Problem description.
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [21] after a single pass of the data. This concept refers to how well a model captures uncertainty in regions of input space that lie between, or away from, observed data. Intuitively, we expect uncertainty to be higher in such unexplored areas. This behavior is important for Bayesian decision-making problems for effictive balancing of exploration and exploitation. We consider the one-dimensional dataset introduced in [58].
Model.
We use an MLP with 4 hidden layers and 128 units per layer, and we employ an ELU activation function.
Results.
Figure 1 shows the evolution of the posterior predictive mean and the posterior predictive variance as a function of the number of seen observations. We see that HiLoFi behaves in an intuitively sensible way, showing more uncertainty away from the data, just like a Gaussian process. See Appendix G.5 for the performance of other methods (specifically LRKF and VBLL) on this problem, as well an ablation study that illustrates the effect of changing rank.
Figure 1: In-between uncertainty induced by HiLoFi as a function of the processed observations.
5.2 MNIST
Problem description.
We consider the MNIST contextual bandit task introduced by [49], where the agent is presented with an image and must choose one of ten possible classes as an action, and then gets a binary feedback, based on whether its prediction was correct or not; we refer to this as the incomplete information case. An additional experiment on online classification is in Appendix G.1.1.
Model.
The predictive model (for the 10 label logits or the per-action rewards) is represented using a modified LeNet5 CNN architecture [39] with ELU activation function.
Exploration Methods.
We consider two exploration strategies: Classical TS and our proposed predictive-sampling (PS) as outlined in Table 1. Additional results comparison TS, PS, and $\epsilon$-greedy are in Appendix G.1.3.
Inference methods.
We compare LLL, LRKF, LoFi, and HiLoFi. For LLL, we use the online last-layer variant from [11], without hyperparameter re-optimization at each timestep. Choice for hyperparameters are detailed in Appendix G.1.3.
Bandit results.
Figure 2 shows the average cumulative reward over 10 runs of the contextual bandit version of MNIST. We observe that the top three performing methods are HiLoFi, LoLoFi, and LRKF. This outcome is expected because HiLoFi employs a full-rank covariance matrix in the last layer, which is better than a low-rank approximation of the covariance matrix associated with the last layer (LoLoFi), or a low-rank covariance to model all dependencies across layers (LRKF). For further results showing regret and comparison using $\varepsilon$-greedy, see Appendix G.1.3.
Figure 2: Cumulative average reward for the bandit MNIST problem.
Next, Table 2 reports the average cumulative reward and total runtime (over 10 trials) for each method, comparing posterior predictive sampling (pBayes) with Thompson sampling (TS).
Agent Reward (pBayes) Reward (TS) Time (pBayes) Time (TS)
HiLoFi 31,167.7 8.44
LoLoFi 30,038.5 3.67
LRKF 28,354.9 2.51
LoFi 23,184.5 23,304.3 4.76 38.90
LLL 13,922.3 7,201.9 1.36 191.55
Table 2: Performance comparison of Bayesian decision-making strategies: posterior predictive sampling (pBayes) vs. Thompson sampling (TS). Rewards are cumulative and averaged over 10 runs; time is in minutes.
Overall, pBayes is consistently faster across all methods. For the two agents where both approaches are applicable (LoFi and LLL), the average cumulative reward is nearly identical for LoFi, while pBayes clearly outperforms TS for LLL. However, TS incurs a much higher runtime due to the need to sample repeatedly from the high-dimensional posterior over model parameters, whereas pBayes only requires sampling in the comparatively low-dimensional outcome space (one dimension per arm/class). Finally, as expected, HiLoFi achieves the best performance, followed by LoLoFi and then LRKF. This performance hierarchy comes at a cost in runtime: HiLoFi is slower than LoLoFi, which in turn is slower than LRKF.
5.3 Recommender systems
Problem description.
We study the performance of the methods on the Kuairec dataset of [22], which is derived from a real-world, non-stationary, video recommender system. This dataset is used to study non-stationary contextual bandits in [65].
Dataset.
We group rows in the dataset (for a given user) in blocks of the 10 next videos that the user saw. For the features, we consider the $\log(1+x)$ transform of like count, share count, play count, comment count, complete play count, follow count, reply comment count, and download count. For the target variable, we use the $\log(1+x)$ transform of the watch ratio $0,\infty)$.
Model.
The reward model is a neural network with an embedding layer (one per video/arm), and a dense layer for all features. We join the embedding layer and the dense layer, and we then consider a three-hidden-layer neural network (50 units per layer, ELU activations), and linear output unit.
Algorithm.
We perform sequential Bayesian inference using the relevant method, and then use TS to choose the action at each step (see Section 3.1). The choice of hyperparameters is in G.3.
Results.
Figure 3 shows the average daily reward (left panel) and the running time of each method (right panel). We observe that HiLoFi has the highest average daily reward with second lowest running time. On average, LRKF and LoFi have similar performance, however, LoFi has a higher running time. Next, VBLL has higher average daily reward than OnVBLL, but it is slower. For an in-depth analysis of the results see Appendix G.3.
Figure 3: Recommender system results. Left: average daily reward. Right: running time.
5.4 Bayesian optimization
Problem description.
Following [5], we evaluate the competing methods on classical Bayesian optimization (BO) tasks, where the goal is to maximize an unknown function $g:[0,1]^{D_{\bm{x}}}\to\mathbb{R}$ with a surrogate model trained sequentially on past observations.
Model.
We employ a three-layer MLP (180 units per layer, ELU activations) as the surrogate for all neural-based methods. For the GP baseline, we use a Matérn-5/2 kernel with lengthscale 0.1 and a 20-point buffer to reduce computational cost. While GP performs better when its hyperparameters are tuned during training [28], this incurs much higher computational costs, so we do not pursue it here. The choice of hyperparameters is detailed in Appendix G.4.
Datasets.
We consider seven benchmark functions commonly used in BO [5]: Ackley (2D, 5D, 10D, 50D), Branin (2D), Hartmann (6D), and DrawNN (50D and 200D). These span a range of dimensionalities and multi-modalities, and are standard in evaluating global optimal performance. For all methods, except for DrawNN, we sample a function from the posterior predictive and evaluate it on a fixed set of candidate points generated using a Sobol sequence [54] to find its maximum. Instead, for Drawnn functions, we use projected gradient descent, implemented with the Jaxopt library [4], to optimize the sampled function directly over the continuous domain $[0,1]$.
Algorithm.
We perform sequential Bayesian inference using the relevant method, and then choose the next query point at each step using TS (see Section 3.1). Figure 12 in the Appendix shows the results using the expected improvement algorithm [29].
Results.
Figure 4 (left) shows the final best value on the vertical axis (higher is better), and the amount of run time on the horizontal axis (lower is better). The number of steps for each dataset are in the x-axis in Figure 11 in Appendix G.4. Figure 4 (right) shows the tradeoff between performance (measured by rank) and compute time. Our method is on the Pareto frontier for a range of tradeoffs between time and performance; furthermore, across tasks, the performance and compute time of HiLoFi dominates (better performance and less compute time) those from competing methods except from VBLL (higher performance) and LRKF (lower compute time). Thus, HiLoFi offers strong tradeoffs between compute time and performance. For further analyses and results see Appendix G.4.
Figure 4: Left panel: Performance across all BO benchmark functions. For time, lower is better. For performance, higher is better. The dashed lines correspond to the results for our method, so methods that are above and to the left are better. Right panel: Time versus performance tradeoff plots.
6 Related work
We briefly review deterministic approaches for approximate online parameter inference. Several methods maintain diagonal plus low-rank (DLR) structures for efficient covariance updates, including L-RVGA [36], SLANG [47], and LoFi [8]. Low-rank Kalman filtering has also been explored in state-space models [53], though with stronger restrictions on the rank relative to measurement dimension. Beyond low-rank, FOO-VB [64] exploits Kronecker-structured approximations, while offline methods explore similar ideas for scaling BNNs [41, 56, 34]. Another direction builds on the lottery ticket hypothesis [42, 38], restricting learning to sparse or low-dimensional subspaces. Examples include the subspace neural bandit [15] and PULSE [7], which pre-train projection matrices offline before online adaptation.
7 Conclusion and future work
We introduced a predictive-first approach for efficient online training of neural networks, based on linearization and structured low-rank approximations of error covariances. Our three variants achieve strong performance-runtime trade-offs, with HiLoFi particularly effective in high-dimensional, non-stationary settings. Limitations include the lack of fixed-lag smoothing, sensitivity to linearization and outliers [14, 37], and reliance on hyperparameters. Future work includes extending our approach to fully-online reinforcement learning [18].
We view our contributions as primarily methodological, with no particular ethical concerns.
Acknowledgements
We thank the TPU Research Cloud program for providing the compute resources used in our experiments, and James Harrison for helpful comments.
Appendix
Appendix A Notation
Symbol Description
$\|\cdot\|_{2}$ $\ell_{2}$ norm.
$\|\cdot\|_{F}$ Frobenius norm.
${D_{\bm{y}}},{D_{\bm{x}}}\in\mathbb{N}$ Dimension of observations / inputs.
${D_{{\bm{h}}}},{D_{{\boldsymbol{\ell}}}}\in\mathbb{N}$ Number of parameters for the hidden layer / last layer.
${D_{\bm{\theta}}}={D_{{\bm{h}}}}+{D_{{\boldsymbol{\ell}}}}$ Total number of neural network parameters.
${\cal N}({\bm{x}}\,|\,{\bm{m}},\mathbf{S})$ Multivariate Gaussian density evaluated at $\bm{x}$} with mean $\bm{m}$} and covariance matrix $\mathbf{S}$.
$p_{\rm env}({\bm{y}}\,|\,{\bm{x}})$ Unknown data-generating process for $\bm{y}$} conditioned on $\bm{x}$}
$p_{{\bm{b}}}({\bm{y}}\,|\,{\bm{x}})$ Posterior predictive model for $\bm{y}$}, conditioned on $\bm{x}$}, and parameterized by $\bm{b}$}.
$\mathcal{M}_{m\times n}(\mathbb{R})$} Space of $m$-by-$n$ real-valued matrices.
${\mathcal{M}^{++}_{m}(\mathbb{R})}$ Space of $m$-dimensional positive definite (pd) matrices.
${\mathcal{M}^{+}_{m}(\mathbb{R})}$ Space of $m$-dimensional positive semidefinite (psd) matrices.
$\mathbf{I}_{D}$ $D$-dimensional identity matrix.
$\mathbf{S}^{1/2}$ Upper-triangular Cholesky decomposition of the pd matrix $\mathbf{S}$.
${\bm{x}}_{t}=({\bm{c}}_{t},{\bm{a}}_{t})$ Inputs with context ${\bm{c}}_{t}\in\mathbb{R}^{{D_{c}}}$ (possibly empty) and action ${\bm{a}}_{t}\in{\cal A}\subseteq\mathbb{R}^{D_{a}}$.
$y_{t}\in\mathbb{R}$ Scalar measurements, rewards, or observations obtained at timestep $t$.
${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ Measurements, rewards, or observations obtained at timestep $t$.
$\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$ Datapoint at time $t$.
${\bm{u}}_{1:t}=({\bm{u}}_{1},\ldots,{\bm{u}}_{t})$ Time-ordered collection of vectors ${\bm{u}}_{k}\in\mathbb{R}^{D_{{\bm{u}}}}$.
$\mathcal{D}_{1:t}=(\mathcal{D}_{1},\ldots,\mathcal{D}_{t})$ Dataset at time $t$.
${\boldsymbol{\ell}}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ Last layer parameters.
${\bm{h}}\in\mathbb{R}^{D_{{\bm{h}}}}$ Hidden layer parameters.
${\bm{\theta}}=({\boldsymbol{\ell}},{\bm{h}})\in\mathbb{R}^{D_{\bm{\theta}}}$ Collection of all neural network parameters.
${f}({\bm{\theta}},{\bm{x}})={f}({\boldsymbol{\ell}},{\bm{h}},{\bm{x}})\in\mathbb{R}^{D_{\bm{y}}}$ Neural network output with parameters $\bm{\theta}$} and inputs $\bm{x}$}.
${\bm{e}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ Zero-mean noise with covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$.
${\bm{u}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ Zero-mean dynamics noise with variance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$.
$q_{{\boldsymbol{\ell}},t}\geq 0$ Last-layer dynamics.
$q_{{\bm{h}},t}\geq 0$ Hidden-layer dynamics.
${\bm{u}}_{t,{\boldsymbol{\ell}}}\in\mathbb{R}^{{D_{{\boldsymbol{\ell}}}}}$ Zero-mean dynamics noise of last layer parameters with variance $q_{{\boldsymbol{\ell}},t}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$.
${\bm{u}}_{t,{\bm{h}}}\in\mathbb{R}^{D_{{\bm{h}}}}$ Zero-mean dynamics noise of hidden-layer parameters with variance $q_{{\bm{h}},t}\,\mathbf{I}_{D_{{\bm{h}}}}$.
${\boldsymbol{\ell}}_{t|t}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$ Frequentist estimate for ${\boldsymbol{\ell}}_{t}$ given $\mathcal{D}_{1:t}$.
${\bm{h}}_{t|t}\in\mathbb{R}^{D_{{\bm{h}}}}$ Frequentist estimate for ${\bm{h}}_{t}$ given $\mathcal{D}_{1:t}$.
${\bm{\theta}}_{t|t}=({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t})\in\mathbb{R}^{D_{\bm{\theta}}}$ Frequentist estimate for ${\bm{\theta}}_{t}$ given $\mathcal{D}_{1:t}$.
$\boldsymbol{\Sigma}_{t}={\rm Var}({\bm{\theta}}_{t}-{\bm{\theta}}_{t|t})$ Error variance-covariance matrix.
$\tilde{\boldsymbol{\Sigma}}_{t}$ Algorithm-dependent surrogate covariance matrix to $\boldsymbol{\Sigma}_{t}$.
$\hat{\boldsymbol{\Sigma}}_{t}=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ Best rank-$d$ approximation to the surrogate covariance matrix $\tilde{\boldsymbol{\Sigma}}_{t}$.
$\mathbf{C}_{t}\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Rectangular matrix $(d\times D)$ matrix such that $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}=\hat{\boldsymbol{\Sigma}}_{t}$
$\tilde{\mathbf{L}}_{t+1}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t},\,{\bm{x}}_{t+1})$ Jacobian of neural network w.r.t parameters in last layer.
$\tilde{\mathbf{H}}_{t+1}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t|t},\,{\bm{h}}_{t|t},\,{\bm{x}}_{t+1})$ Jacobian of neural network w.r.t. parameter in hidden layers.
${\cal Q}_{R}(\mathbf{B}_{1},\ldots,\mathbf{B}_{k})$ Cholesky factorization of $\sum_{k=1}^{K}\mathbf{B}_{k}^{\intercal}\,\mathbf{B}_{k}$, where each $\mathbf{B}_{k}$ is an upper-triangular Cholesky factor (see Appendix C.1).
${\cal P}_{d}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}$, with $\mathbf{A}_{k}\in{\mathcal{M}_{d_{k}\times D}(\mathbb{R})}$, $d_{k}\leq D$.
${\cal P}_{d,+a}(\mathbf{A}_{1},\ldots,\mathbf{A}_{K})\in{\mathcal{M}_{d\times D}(\mathbb{R})}$ Best rank-$d$ approximation of $\sum_{k=1}^{K}\mathbf{A}_{k}^{\intercal}\,\mathbf{A}_{k}+a\,\mathbf{I}_{D}$, with $a>0$ (see Appendix C.2).
Appendix B Extended Kalman filtering for online learning
Here, we explain in more detail the background of the EKF for online learning presented in Section 3.2. To make this section self-contained, we repeat parts of the material introduced in that section. Let us assume that ${\bm{y}}_{t}={\boldsymbol{\ell}}_{t}^{\intercal}\,\phi({\bm{x}}_{t},{\bm{h}}_{t})+{\bm{e}}_{t}$, where ${\bm{e}}_{t}$ is a zero-mean random variable with observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. Given a starting vector ${\bm{\theta}}_{0}$, we assume that model parameters ${\bm{\theta}}_{t}=({\boldsymbol{\ell}}_{t},{\bm{h}}_{t})\in\mathbb{R}^{D_{\bm{\theta}}}$ and observations ${\bm{y}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ evolve according to the state-space model
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (11)
${\bm{y}}_{t}$ $={f}({\bm{\theta}}_{t},\,{\bm{x}}_{t})+{\bm{e}}_{t},$
where ${\bm{u}}_{t}\in\mathbb{R}^{D_{\bm{\theta}}}$ is a zero-mean random vector with known dynamics covariance $\mathbf{Q}_{t}\in{\mathcal{M}^{+}_{{D_{\bm{\theta}}}}(\mathbb{R})}$, and ${\bm{e}}_{t}\in\mathbb{R}^{D_{\bm{y}}}$ is a zero-mean random vector with known observation covariance $\mathbf{R}_{t}\in{\mathcal{M}^{++}_{{D_{\bm{y}}}}(\mathbb{R})}$. As before, we assume ${\rm Cov}({\bm{\theta}}_{0},{\bm{e}}_{t})=\boldsymbol{0}$ for all $t\in\mathbb{N}$ and ${\rm Cov}({\bm{u}}_{t},{\bm{e}}_{k})=\boldsymbol{0}$ for all $t,k\in\mathbb{N}$.
We now consider a first-order approximation of f{f} around ${\bm{\theta}}_{t-1|t-1}$, that is,
${\bm{y}}_{t}\approx{f}({\bm{\theta}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}\,({\bm{\theta}}_{t}-{\bm{\theta}}_{t-1|t-1})+{\bm{e}}_{t}\,$, (12)
where ${\bm{\theta}}_{t-1|t-1}=\begin{bmatrix}{\boldsymbol{\ell}}_{t-1|t-1}^{\intercal}&{\bm{h}}_{t-1|t-1}^{\intercal}\end{bmatrix}^{\intercal}$ is given, and $\tilde{\mathbf{J}}_{t}=\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$ with $\tilde{\mathbf{L}}_{t}=\nabla_{{\boldsymbol{\ell}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$ and $\tilde{\mathbf{H}}_{t}=\nabla_{{\bm{h}}}{f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})$.
Following a frequentist approach to the Kalman filter [27] and given the starting latent random vector ${\bm{\theta}}_{0|0}:={\bm{\theta}}_{0}$, we seek to obtain sequential updates for the best ($L^{2}$) linear estimate of the expected value of ${\bm{\theta}}_{t}$ given data ${\bm{y}}_{1:t}$. We formalize this in the following proposition.
Proposition B.1.
Let $k,t\in\mathbb{N}$ and let ${\bm{y}}_{t}$, ${\bm{\theta}}_{t}$ follow the SSM (12). The solution to the optimization problem
[…]
is the matrix $\mathbf{A}_{t|k}^{\star}$ given by
$\mathbf{A}^{\star}_{t|k}={\rm Cov}({\bm{\theta}}_{t},{\bm{y}}_{1:k})\,{\rm Var}({\bm{y}}_{1:k})^{-1}.$ (14)
The best linear unbiased predictor (BLUP) for model parameters ${\bm{\theta}}_{t}$, given observations ${\bm{y}}_{1:k}$ is
${\bm{\theta}}_{t|k}$ $=\mathbf{A}_{t|k}^{\star}\,{\bm{y}}_{1:k},$ (15)
and the error variance covariance (EVC) matrix is defined as
$\boldsymbol{\Sigma}_{t|k}$ $={\rm Var}({\bm{\theta}}_{t})-\mathbf{A}_{t|k}^{\star}\,{\rm Var}({\bm{y}}_{1:k})\,\mathbf{A}_{t|k}^{\star\intercal}.$ (16)
Proposition B.2.
Under the SSM (12), the BLUP and the EVC can be written in the form of Kalman filtering predict and update equations. Here, the predict equations are
${\bm{\theta}}_{t|t-1}$ $={\bm{\theta}}_{t-1|t-1},$ (17)
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t},$
and the update equations are
${\bm{\theta}}_{t|t}$ $={\bm{\theta}}_{t|t-1}+\mathbf{K}_{t}\,\boldsymbol{\epsilon}_{t},$ (18)
$\boldsymbol{\Sigma}_{t|t}$ $=(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})\,\boldsymbol{\Sigma}_{t|t-1}(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\mathbf{K}_{t}^{\intercal},$
with
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\bm{\theta}}_{t-1|t-1},{\bm{x}}_{t}),$ (19)
$\mathbf{K}_{t}$ $=\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1},$
$\mathbf{S}_{t}$ $=\tilde{\mathbf{J}}_{t}\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}+\mathbf{R}_{t}.$
Proof.
The proof follows from the linear form of (12), Lemmas 2.1 to 2.3, and Theorem 4.2 in [19].
Remark B.3.
The update equations in (18) are recursive and characterize estimate of the unknown latent state and the estimated error estimation through the first two moments. Furthermore, the update for the covariance in (18) is in Joseph form [63] and is known to be numerically stable. In our method, we leverage these two facts to target a low-rank EVC through a second-step optimization that has numerically-stable updates.
Corollary B.4 (Kalman filter as a Bayesian posterior).
Assume a Gaussian density for the initial parameters $p({\bm{\theta}}_{0})={\cal N}({\bm{\theta}}_{0}\,|\,{\bm{\theta}}_{0|0},\boldsymbol{\Sigma}_{0|0})$ and for the noise term $p({\bm{e}}_{t})={\cal N}({\bm{e}}_{t}\,|\,\boldsymbol{0},\,\mathbf{R}_{t})$. The BLUP ${\bm{\theta}}_{t|t}$ and the EVC $\boldsymbol{\Sigma}_{t|t}$ are the parameters of the Gaussian that characterize the posterior of model parameters ${\bm{\theta}}_{t}$ given ${\bm{y}}_{1:t}$, more precisely,
$p({\bm{\theta}}_{t}\,|\,{\bm{y}}_{1:t})={\cal N}({\bm{\theta}}_{t}\,|\,{\bm{\theta}}_{t|t},\boldsymbol{\Sigma}_{t|t}).$ (20)
As a consequence, the posterior predictive for the next observation in the sequence is
$p({\bm{y}}_{t}\,|\,{\bm{y}}_{1:t-1})$ $=\int{\cal N}\left({\bm{y}}_{t}\,|\,{f}({\bm{\theta}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1}),\mathbf{R}_{t}\right)p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}$ (21)
$={\cal N}({\bm{y}}_{t}\,|\,{f}({\bm{\theta}}_{t-1|t-1},{\bm{x}}_{t}),\;\;\tilde{\mathbf{J}}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\tilde{\mathbf{J}}_{t}^{\intercal}+\ \mathbf{R}_{t}).$
Appendix C Linear algebra results
C.1 Sum of Cholesky matrices
Here, we generalize the result found in [57] to any set of $K>2$ matrices, which we use to derive the method LRKF, HiLoFi, and LoLoFi.
Proposition C.1 (QR of sum of Cholesky matrices).
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,I\}$ be a collection of $(D\times D)$ positive definite matrices and let $\mathbf{A}_{i}^{1/2}$ be the upper-triangular Cholesky decomposition of $\mathbf{A}_{i}$. Let $\mathbf{M}$ be the stacked $(D\,I\times D)$ matrix given by
$\mathbf{M}=\begin{bmatrix}\mathbf{A}^{1/2}_{1}\\ \mathbf{A}^{1/2}_{2}\\ \vdots\\ \mathbf{A}^{1/2}_{I}\end{bmatrix}$. (22)
It follows that
$\mathbf{R}^{\intercal}\,\mathbf{R}=\sum_{i=1}^{I}\mathbf{A}_{i}\,.$ (23)
where $\mathbf{R}$ is R component in the QR decomposition of $\mathbf{M}$.
Proof.
We see that
$\mathbf{M}^{\intercal}\,\mathbf{M}=\sum_{i=1}^{I}\mathbf{A}_{i}\,.$ (24)
Consider the QR decomposition of $\mathbf{M}$
$\mathbf{M}=\mathbf{Q}\mathbf{R},$ (25)
with $\mathbf{Q}$ an orthogonal matrix and $\mathbf{R}$ and upper-triangular matrix.
Then,
$\sum_{i=1}^{I}\mathbf{A}_{i}=\mathbf{M}^{\intercal}\,\mathbf{M}=\mathbf{R}^{\intercal}\,\mathbf{Q}^{\intercal}\,\mathbf{Q}\,\mathbf{R}=\mathbf{R}^{\intercal}\,\mathbf{R}.$ (26)
As a consequence, we note that the square root of the matrix $\mathbf{A}_{1}+\dots+\mathbf{A}_{I}$ is the upper-triangular matrix in the QR decomposition of $\mathbf{M}$. ∎
Remark
In what follows, we denote the result of obtaining the $\mathbf{R}$ matrix of the QR decomposition of the row-stacked matrices $\mathbf{A}^{1/2}_{i}$ for $i\in\{1,\dots,I\}$ by
${\cal Q}_{R}(\mathbf{A}^{1/2}_{1},\dots,\mathbf{A}^{1/2}_{I}).$ (27)
C.2 Singular value decomposition given sum of low-rank-matrices
Proposition C.2 (SVD of sum of low-rank matrices).
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,N\}$ be ${\mathcal{M}^{++}_{D}(\mathbb{R})}$ positive semi-definite matrices such that $\mathbf{A}_{i}=\mathbf{W}_{i}^{\intercal}\,\mathbf{W}_{i}$, where $\mathbf{W}_{i}\in\mathbb{R}^{d\times D}$ and $d\ll D$. The best rank-$d$ approximation of the sum is
$\sum_{i=N}^{I}\mathbf{A}_{i}\approx\mathbf{J}^{\intercal}\,\mathbf{J}\,$, (28)
where $\mathbf{J}=\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D}$, $\mathbf{S}_{:d}$ are the top $d$ singular values and $\mathbf{V}$ contains the right singular vectors of the stacked $(d\,N\times D)$ matrix given by
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\ \mathbf{W}_{2}\\ \vdots\\ \mathbf{W}_{N}\end{bmatrix}$. (29)
Proof.
Let $\mathbf{A}_{i}$ for $i\in\{1,2,\dots,N\}$ be ${\mathcal{M}^{++}_{D}(\mathbb{R})}$ matrices with $\mathbf{A}_{i}=\mathbf{W}_{i}^{\intercal}\mathbf{W}_{i}$, $\mathbf{W}_{i}\in\mathbb{R}^{d\times D}$, and $d\ll D$. Form the matrix $\mathbf{N}$ given by
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\[6.0pt] \mathbf{W}_{2}\\[3.0pt] \vdots\\[3.0pt] \mathbf{W}_{N}\end{bmatrix}=\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal},$ (30)
where $\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$ is its full singular value decomposition, with $\mathbf{U}\in\mathbb{R}^{dN\times dN}$, $\mathbf{S}\in\mathbb{R}^{dN\times D}$, and $\mathbf{V}\in\mathbb{R}^{D\times D}$. Then
$\sum_{i=1}^{N}\mathbf{A}_{i}$ $=\sum_{i=1}^{N}\mathbf{W}_{i}^{\intercal}\mathbf{W}_{i}=\mathbf{N}^{\intercal}\mathbf{N}$ (31)
$=(\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal})^{\intercal}\,(\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal})$
$=\mathbf{V}\,\mathbf{S}^{\intercal}\,\mathbf{U}^{\intercal}\,\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$
$=\mathbf{V}\,\mathbf{S}^{2}\,\mathbf{V}^{\intercal}.$
Hence $\sum_{i}\mathbf{A}_{i}$ has eigenvectors $\mathbf{V}$ and eigenvalues given by the diagonal entries of $\mathbf{S}^{2}$. By the symmetric‐matrix form of the Eckart–Young theorem [17, 10], its best rank-$d$ approximation in Frobenius norm is
$\mathbf{V}\,{\rm diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2},0,\dots)\,\mathbf{V}^{\intercal}.$ (32)
Writing $\mathbf{S}_{:d}={\rm diag}(\sigma_{1},\dots,\sigma_{d})\in\mathbb{R}^{d\times D}$, one checks
$\mathbf{V}\,{\rm diag}(\sigma_{1}^{2},\dots,\sigma_{d}^{2},0,\dots)\,\mathbf{V}^{\intercal}=\bigl(\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\bigr)^{T}\,\bigl(\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\bigr)=\mathbf{J}^{\intercal}\,\mathbf{J},$ (33)
where
$\mathbf{J}$ $=\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D},$ (34)
$\mathbf{S}_{:d}$ $={\rm diag}\left(\sigma_{1},\dots,\sigma_{d}\right),$
and $\{\sigma_{k}\}$ are the singular vectors of $\sum_{n=1}^{N}\mathbf{A}_{n}$. ∎
Takeaway.
Proposition C.2 shows that to form the best rank-$d$ approximation of $\sum_{i=1}^{N}\mathbf{A}_{i}$ one does not need to assemble or carry out the SVD of the full $D\times D$ sum. Instead, one stacks the low-rank factors into
$\mathbf{N}=\begin{bmatrix}\mathbf{W}_{1}\\ \vdots\\ \mathbf{W}_{N}\end{bmatrix}\in\mathbb{R}^{(dN)\times D},$ (35)
and compute a reduced SVD (or symmetric eigendecomposition) of the small $(dN)\times(dN)$ Gram matrix $\mathbf{N}\mathbf{N}^{\top}$. This costs
$\mathcal{O}((dN)^{2}D+(dN)^{3})\approx\mathcal{O}(d^{2}N^{2}D)\text{ when }dN\ll D,$ (36)
versus the $\mathcal{O}(D^{3})$ required for a full SVD on the $D\times D$ sum. Moreover, memory and computation scale with $d$ and $dN$ rather than with the large dimension $D$ because all operations involve only the $(dN)\times D$ matrix $\mathbf{N}$ and the $d\times D$ factor $\mathbf{J}$.
In what follows, we denote the result of obtaining the best rank-$d$ approximation matrix of the SVD decomposition of the row-stacked matrices $\mathbf{W}^{1/2}_{i}$ for $i\in\{1,\dots,N\}$ by
${\cal P}_{d}(\mathbf{W}_{1},\ldots,\mathbf{W}_{N}).$ (37)
Corollary C.3.
If the sum of matrices is of the form
$\sum_{n=1}^{N}\mathbf{A}_{n}+a\,\mathbf{I}_{D}$ (38)
with $a>0$, then the best rank-$d$ approximation is of the form
$\sum_{n=1}^{N}\mathbf{A}_{n}+a\,\mathbf{I}_{D}\approx\mathbf{J}_{+q}^{\intercal}\,\mathbf{J}_{+q}$ (39)
where
$\mathbf{J}_{+q}$ $=\mathbf{S}_{:d,+q}\,\mathbf{V}^{\intercal}\in\mathbb{R}^{d\times D},$ (40)
$\mathbf{S}_{:d,+q}$ $={\rm diag}\left(\sqrt{\sigma_{1}^{2}+q},\dots,\sqrt{\sigma_{d}^{2}+q}\right),$
and $\{\sigma_{k}\}_{k=1}^{d}$ are the top-$d$ singular values of $\sum_{n=1}^{N}\mathbf{A}_{n}$.
Proof.
From (31) in Proposition C.2 we obtain
$\sum_{i=1}^{N}\mathbf{A}_{i}+a\,\mathbf{I}=\mathbf{V}\,\mathbf{S}^{2}\,\mathbf{V}^{\intercal}+a\mathbf{I}=\mathbf{V}\,(\mathbf{S}^{2}+a\mathbf{I}_{D})\,\mathbf{V}^{\intercal}=\mathbf{V}\,\mathbf{S}_{+a}^{2}\,\mathbf{V}^{\intercal},$ (41)
where
$\mathbf{S}_{+a}^{2}={\rm diag}(\sigma_{1}^{2}+a,\ldots,\sigma_{D}^{2}+a).$ (42)
Let
$\mathbf{S}_{+a}={\rm diag}\left(\sqrt{\sigma_{1}^{2}+a},\ldots,\sqrt{\sigma_{D}^{2}+a}\right),$ (43)
then
$\sum_{n=1}^{N}\mathbf{A}_{n}$ $=\mathbf{V}\,\mathbf{S}_{+a}^{2}\,\mathbf{V}^{\intercal}$ (44)
$=(\mathbf{V}\,\mathbf{S}_{+a}\mathbf{V}^{\intercal})^{\intercal}\,(\mathbf{V}\,\mathbf{S}_{+a}\mathbf{V}^{\intercal})$
$=\mathbf{J}_{+a}^{\intercal}\,\mathbf{J}_{+a}.$
Takeaway.
If the sum of matrices can be written as the sum of multiplied low-rank factors plus an additional identity matrix times a constant, then, computation of the best rank $d$-matrix in low-rank factor can be obtained by performing SVD over the low-rank factors and modifying the singular values to include the term $a$.
In what follows, we denote the result of obtaining the best rank-$d$ approximation matrix of the SVD decomposition of the row-stacked matrices $\mathbf{W}^{1/2}_{i}$ for $i\in\{1,\dots,N\}$ plus an identity times a real-valued number $a>0$ as
${\cal P}_{d,+a}(\mathbf{W}_{1},\ldots,\mathbf{W}_{N}).$ (45)
Appendix D HiLoFi— further details
D.1 Derivation of HiLoFi
Consider the linearized model
${\bm{y}}_{t}={f}({\boldsymbol{\ell}}_{t-1|t-1},\,{\bm{h}}_{t-1|t-1},\,{\bm{x}}_{t})+\tilde{\mathbf{L}}_{t}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t-1|t-1})+\tilde{\mathbf{H}}_{t}({\bm{h}}_{t}-{\bm{h}}_{t-1|t-1})+{\bm{e}}_{t},$ (46)
where ${\boldsymbol{\ell}}_{t-1|t-1}$ and ${\bm{h}}_{t-1|t-1}$ are given.
[…]
$\mathbb{E}[{\boldsymbol{\ell}}_{0}]$ $={\boldsymbol{\ell}}_{0|0},$ $\mathbb{E}[{\bm{h}}_{0}]$ $={\bm{h}}_{0|0},$ ${\rm Var}({\boldsymbol{\ell}}_{0})$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0},$ ${\rm Var}({\bm{h}}_{0})$ $=\mathbf{C}_{0}^{\intercal}\,\mathbf{C}_{0},$ (47)
for known $\mathbf{C}_{0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$, ${\bm{h}}_{0|0}\in\mathbb{R}^{D_{{\bm{h}}}}$, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}^{++}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$, and ${\boldsymbol{\ell}}_{0|0}\in\mathbb{R}^{D_{{\boldsymbol{\ell}}}}$. Next, we assume that for $t\geq 1$, the latent last layer parameters ${\boldsymbol{\ell}}_{1:t}$ and the hidden layer parameters ${\bm{h}}_{1:t}$ follow the dynamics
${\boldsymbol{\ell}}_{t}={\boldsymbol{\ell}}_{t-1}+{\bm{u}}_{{\boldsymbol{\ell}},t},\qquad{\bm{h}}_{t}={\bm{h}}_{t-1}+{\bm{u}}_{{\bm{h}},t},\\$ (48)
where ${\bm{u}}_{{\boldsymbol{\ell}},1:t}$ and ${\bm{u}}_{{\bm{h}},1:t}$ are zero-mean independent noise variables111i.e., ${\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},i},{\bm{u}}_{{\bm{h}},j})=\boldsymbol{0}$ for all $i,j\in\{1,\ldots,T\}$. with dynamics covariance matrices ${\rm Var}({\bm{u}}_{{\boldsymbol{\ell}},t})=\mathbf{Q}_{{\boldsymbol{\ell}},t}\in{\mathcal{M}^{+}_{{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and ${\rm Var}({\bm{u}}_{{\bm{h}},t})=\mathbf{Q}_{{\bm{h}},t}\in{\mathcal{M}^{+}_{{D_{{\bm{h}}}}}(\mathbb{R})}$. For simplicity and to exploit efficient linear algebra techniques for belief updates, we assume $\mathbf{Q}_{{\boldsymbol{\ell}},t}=q_{{\boldsymbol{\ell}},t}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}$ and $\mathbf{Q}_{{\bm{h}},t}=q_{{\bm{h}},t}\mathbf{I}_{D_{{\bm{h}}}}$. A simple choice is to set $\mathbf{Q}_{{\bm{h}},t}=0\mathbf{I}$ to avoid forgetting past data; see e.g., [16].
Proposition D.1 (Predict step).
With the SSM assumption (48), the approximate covariance ${\rm Var}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and the approximate covariance ${\rm Var}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}$, the predict step becomes
${\bm{h}}_{t|t-1}$ $={\bm{h}}_{t-1|t-1},$ ${\boldsymbol{\ell}}_{t|t-1}$ $={\boldsymbol{\ell}}_{t-1|t-1},$ (49)
$\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}$ $=\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}+q_{{\boldsymbol{\ell}},t}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},$ $\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}$ $=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{{\bm{h}},t}\,\mathbf{I}_{D_{{\bm{h}}}}.$
The proof of Proposition D.1 is in Appendix E.1.
Proposition D.2 (Variance of the innovation).
The upper Cholesky decomposition of innovation variance takes the form
$\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right),$ (50)
where $\tilde{\mathbf{L}}_{t}$ and $\tilde{\mathbf{H}}_{t}$ are defined in Section 3.2.
The proof of Proposition D.2 is in Appendix E.2.
Next, our update step for the BLUP resembles an EKF update, while the EVC update approximates a block-diagonal covariance: the last layer block is full-rank, and the hidden layer blocks are low-rank. To maintain numerical stability and low-memory updates, we track the Cholesky factor for the last layer and a low-rank factor for the hidden layers. The Cholesky factor is taken from a surrogate matrix that avoids computing a full $\mathcal{M}_{{D_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})$} matrix by neglecting the artificial dynamics covariance $q_{{\boldsymbol{\ell}},t}$. The low-rank component for hidden layers is approximated via a two-step process: (i) a fast surrogate predicted covariance, and (ii) a rank $D_{{\bm{h}}}$} projection error. Like the Cholesky factor, the surrogate covariance in (i) reduces the computational cost of dynamics noise, which does not reflect the system’s true dynamics. However, we retain some information from the dynamics noise to update all hidden layer parameters. Step (ii) enables a fast update rule, crucial for overparameterized neural networks [38]. We detail this step in the following proposition.
Proposition D.3 (Kalman gain, update BLUP, and update EVC).
Define the gain matrices
$\mathbf{K}_{{\bm{h}},t}^{\intercal}=\mathbf{V}_{{\bm{h}},t}\mathbf{C}_{t-1}\mathbf{C}_{t-1}^{\intercal}+q_{{\bm{h}},t}\mathbf{V}_{{\bm{h}},t},\quad\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}=\mathbf{V}_{{\boldsymbol{\ell}},t}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}+q_{{\boldsymbol{\ell}},t}\mathbf{V}_{{\boldsymbol{\ell}},t},$ (51)
where $\mathbf{V}_{{\bm{h}},t}=\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\tilde{\mathbf{H}}_{t},\quad$ and $\mathbf{V}_{{\boldsymbol{\ell}},t}=\mathbf{S}_{t}^{-1/2}\,\mathbf{S}_{t}^{-\intercal/2}\,\tilde{\mathbf{L}}_{t}$. The updated BLUP for the last layer parameters and hidden layer parameters are
${\bm{h}}_{t|t}={\bm{h}}_{t-1|t-1}+\mathbf{K}_{{\bm{h}},t}\,\boldsymbol{\epsilon}_{t},\quad{\boldsymbol{\ell}}_{t|t}={\boldsymbol{\ell}}_{t-1|t-1}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\boldsymbol{\epsilon}_{t}.$ (52)
The approximate posterior covariance $\hat{\boldsymbol{\Sigma}}_{t|t}$ is the best block-diagonal approximation (in Frobenius norm) of the surrogate matrix $\tilde{\boldsymbol{\Sigma}}_{t|t}$ given by
$\scriptsize\left[\begin{smallmatrix}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\hat{\boldsymbol{\Sigma}}_{t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}-\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}-\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right],$
that has full rank for the last layer and rank $d_{{\bm{h}}}$} for the hidden layer. This results in a Cholesky factor for the last-layer covariance given by
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}={\cal Q}_{R}\left(\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right),$ (53)
and a low-rank factor for the hidden layers given by
$\mathbf{C}_{t}={\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal},\;\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right),$ (54)
The proof is in Appendix E.3.
D.2 Computational complexity
We now compare the computational complexity of the algorithms we employ. Recall that $D_{{\boldsymbol{\ell}}}$}, $D_{{\bm{h}}}$}, and $D_{\bm{y}}$} are the number of parameters in the last layer, hidden layers, and the size of the output layer, respectively. As before, ${D_{\bm{\theta}}}={D_{{\boldsymbol{\ell}}}}+{D_{{\bm{h}}}}$. The dimensions ${d_{{\boldsymbol{\ell}}}},{d_{{\bm{h}}}}$, with ${d_{{\boldsymbol{\ell}}}}\ll{D_{{\boldsymbol{\ell}}}}$ and ${d_{{\bm{h}}}}\ll{D_{{\bm{h}}}}$, are the low-dimensional subspaces for the last and hidden layers, respectively, and $d={d_{{\boldsymbol{\ell}}}}+{d_{{\bm{h}}}}$.
The computational complexity of our algorithm when using full-rank in the last layer is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$. For low-dimensional outputs, this is $O({D_{{\boldsymbol{\ell}}}}^{3}+{D_{{\bm{h}}}}{d_{{\bm{h}}}}^{2})$. For high-dimensional outputs, we may choose to use a low-rank approximation for the last layer (which we call LoLoFi) this reduces the cost to $O({D_{{\boldsymbol{\ell}}}}\,({d_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2})$, which is linear in $D_{\bm{\theta}}$} The primary computational bottlenecks in our method are the calculations for the Kalman gain of the hidden and last layers, with costs $O(2\,{D_{{\bm{h}}}}\,{D_{\bm{y}}}^{2})$ and $O(2\,{D_{{\boldsymbol{\ell}}}}\,{D_{\bm{y}}}^{2})$ respectively. This efficiency arises because $\mathbf{S}_{t}^{1/2}$ is upper triangular, allowing the system $\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2}\,\mathbf{V}=\tilde{\mathbf{J}}$ for $\mathbf{V}$ to be solved in $O(2\,{D_{{\bm{h}}}}\,{D_{\bm{y}}}^{2})$ as opposed to $O({D_{{\bm{h}}}}\,{D_{\bm{y}}}^{3})$. Additionally, approximating the covariance matrix via truncated SVD incurs a cost of $O({D_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})^{2}+({d_{{\bm{h}}}}+{D_{\bm{y}}})^{3}+{d_{{\bm{h}}}}\,({d_{{\bm{h}}}}+{D_{\bm{y}}})\,{D_{{\bm{h}}}})$, as detailed in Figure 8.6.1 in [23]. Finally, the corresponding cost for the last layer is $O({D_{{\boldsymbol{\ell}}}}\,({D_{{\boldsymbol{\ell}}}}+{D_{\bm{y}}})^{2})$.
Among related methods, the closest to ours in terms of computational costs are variational Bayes approaches such as the Slang method [47], the L-RVGA method [36], and the LoFi method [8]. In particular, LoFi uses the linearized Gaussian updates (similar to the ones in this paper), but approximates the precision matrix using a diagonal-plus-low rank (DLR) form. The appeal for DLR precision matrices is twofold. First, they enable the application of the Woodbury identity, leading to a predict step of cost $O({D_{\bm{\theta}}}\,d+d^{3})$ and an update step of $O({D_{\bm{\theta}}}\,(d+{D_{\bm{y}}})^{2})=O({D_{{\boldsymbol{\ell}}}}\,(d+{D_{\bm{y}}})^{2}+{D_{{\bm{h}}}}\,(d+{D_{\bm{y}}})^{2})$. Second, incorporating positive diagonal terms ensures the matrix is positive definite, so that a valid posterior Gaussian density is defined.
Although HiLoFi and LoFi have the same asymptotic complexity, LoFi incurs additional practical overhead due to three key operations absent in HiLoFi: (1) the inversion of a $(d+{D_{\bm{y}}})$ rectangular matrix, (2) the inversion of a $d$ rectangular matrix, and (3) the Cholesky decomposition of a $d$ rectangular matrix. These operations increase the actual per-step computational time, which in our example, scales at about one second per additional rank. This behaviour is shown in the empirical comparison across HiLoFi, LRKF, and LoFi in the online classification setting (Figure 6, Appendix G.1).
Appendix E Proofs
E.1 Proof of Proposition D.1
Proof.
Following Proposition B.2, the predicted mean takes the form
${\bm{\theta}}_{t|t-1}={\bm{\theta}}_{t-1|t-1}=\begin{bmatrix}{\boldsymbol{\ell}}_{t-1|t-1}\\ {\bm{h}}_{t-1|t-1}\end{bmatrix}$. (55)
Next, the predicted posterior covariance takes the form
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t},$ (56)
where
$\boldsymbol{\Sigma}_{t-1|t-1}$ (57)
$\quad={\rm Var}({\bm{\theta}}_{t-1}-{\bm{\theta}}_{t-1|t-1})$
$\quad=\begin{bmatrix}{\rm Var}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&{\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})\\ {\rm Cov}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1},{\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&{\rm Var}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1}).\end{bmatrix}$
$\quad=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1}\\ &{\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})\\ {\rm Cov}({\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1},{\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1})&\mathbf{W}_{t-1}^{\intercal}\,\mathbf{W}_{t-1}\end{bmatrix}$.
Next,
${\rm Cov}({\boldsymbol{\ell}}_{t-1}-{\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1}-{\bm{h}}_{t-1|t-1})$ (58)
$\quad={\rm Cov}({\boldsymbol{\ell}}_{t-1},{\bm{h}}_{t-1})$
$\quad={\rm Cov}\left({\boldsymbol{\ell}}_{0}+\sum_{\tau=1}^{t-1}{\bm{u}}_{{\boldsymbol{\ell}},\tau},\,{\bm{h}}_{0}+\sum_{\tau=1}^{t-1}{\bm{u}}_{{\bm{h}},\tau}\right)$
$\quad={\rm Cov}({\boldsymbol{\ell}}_{0},{\bm{h}}_{0})+\sum_{\tau=1}^{t-1}{\rm Cov}({\boldsymbol{\ell}}_{0},{\bm{u}}_{{\bm{h}},\tau})+\sum_{\tau=1}^{t-1}{\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},\tau},{\bm{h}}_{0})+\sum_{\tau=1}^{t-1}\sum_{\tau^{\prime}=1}^{-1}{\rm Cov}({\bm{u}}_{{\boldsymbol{\ell}},\tau},{\bm{u}}_{{\bm{h}},\tau^{\prime}})$
$\quad=\boldsymbol{0},$
which follows from the assumptions about the SSM. Then
$\boldsymbol{\Sigma}_{t-1|t-1}={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1},\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}).$ (59)
and
$\boldsymbol{\Sigma}_{t|t-1}$ $=\boldsymbol{\Sigma}_{t-1|t-1}+\mathbf{Q}_{t}$ (60)
$={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}+q_{\boldsymbol{\ell}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{\bm{h}}\,\mathbf{I}_{D_{{\bm{h}}}}).$
Given that $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}$ is approximated at t1t-1 through $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}$ and $\boldsymbol{\Sigma}_{{\bm{h}},t-1|t-1}$ is approximated at t1t-1 through $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$, we obtain
$\boldsymbol{\Sigma}_{t|t-1,{\boldsymbol{\ell}}}$ $\approx\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}+q_{\boldsymbol{\ell}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}},$ (61)
$\boldsymbol{\Sigma}_{t|t-1,{\bm{h}}}$ $\approx\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{\bm{h}}\,\mathbf{I}_{D_{{\boldsymbol{\ell}}}}.$
E.2 Proof of Proposition D.2
Proof.
By definition of ${\bm{y}}_{t}$ and Proposition B.2, the innovation takes the form
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-{f}({\boldsymbol{\ell}}_{t-1|t-1},{\bm{h}}_{t-1|t-1},{\bm{x}}_{t})$ (62)
$=\tilde{\mathbf{J}}_{t}(\,{\bm{\theta}}_{t}-{\bm{\theta}}_{t|t-1})+{\bm{e}}_{t}$
$=\tilde{\mathbf{L}}_{t}\,({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})+\tilde{\mathbf{H}}_{t}\,({\bm{h}}_{t}-{\bm{h}}_{t|t-1})+{\bm{e}}_{t}.$
Next, the variance of the innovation is
$\mathbf{S}_{t}$ $={\rm Var}(\boldsymbol{\epsilon}_{t})$ (63)
$={\rm Var}\left(\tilde{\mathbf{L}}_{t}\,({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})+\tilde{\mathbf{H}}_{t}\,({\bm{h}}_{t}-{\bm{h}}_{t|t-1})+{\bm{e}}_{t}\right)$
$=\tilde{\mathbf{L}}_{t}\,{\rm Var}({\boldsymbol{\ell}}_{t}-{\boldsymbol{\ell}}_{t|t-1})\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,{\rm Var}({\bm{h}}_{t}-{\bm{h}}_{t|t-1})\tilde{\mathbf{H}}_{t}^{\intercal}+{\rm Var}({\bm{e}}_{t})$
$=\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\tilde{\mathbf{L}}_{t}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}\,\tilde{\mathbf{L}}_{t}^{\intercal}+q_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\,\tilde{\mathbf{L}}_{t}^{\intercal}+\tilde{\mathbf{H}}_{t}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\tilde{\mathbf{H}}_{t}^{\intercal}+q_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\,\tilde{\mathbf{H}}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\begin{bmatrix}\tilde{\mathbf{L}}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{\intercal/2}&\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}\,\mathbf{C}_{t-1}^{\intercal}&\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}&\mathbf{R}^{\intercal/2}\,\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{L}}\,\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t-1}^{1/2}\\ \sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal}\\ \mathbf{C}_{t-1}\,\tilde{\mathbf{H}}^{\intercal}\\ \sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal}\\ \mathbf{R}_{t}^{1/2}\end{bmatrix}$
$=\mathbf{S}_{t}^{\intercal/2}\,\mathbf{S}_{t}^{1/2},$
with
$\mathbf{S}_{t}^{1/2}={\cal Q}_{R}\left(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t}^{1/2}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\sqrt{q_{{\boldsymbol{\ell}},t}}\,\tilde{\mathbf{L}}_{t}^{\intercal},\,\mathbf{C}_{t-1}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\sqrt{q_{{\bm{h}},t}}\,\tilde{\mathbf{H}}_{t}^{\intercal},\,\mathbf{R}_{t}^{1/2}\right).$ (64)
E.3 Proof of Proposition D.3
Proof.
First, by Proposition B.2, the Kalman gain matrix is given by
$\mathbf{K}_{t}=\boldsymbol{\Sigma}_{t|t-1}\,\tilde{\mathbf{J}}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1}.$ (65)
Next, following Proposition D.1 and Proposition D.2, we rewrite the Kalman gain as
$\mathbf{K}_{t}$ $={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1})\begin{bmatrix}\tilde{\mathbf{L}}_{t}\\ \tilde{\mathbf{H}}_{t}\end{bmatrix}\mathbf{S}_{t}^{-1}$ (66)
$={\rm diag}(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1})\begin{bmatrix}\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \tilde{\mathbf{H}}_{t}\,\mathbf{S}_{t}^{-1}\end{bmatrix}$
$=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}\,\tilde{\mathbf{L}}_{t}\,\mathbf{S}_{t}^{-1}\\ \end{bmatrix}$
$=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}}_{t}}\end{bmatrix}$,
where
$\mathbf{K}_{{\boldsymbol{\ell}},t}$ $=\Big(\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{L}}_{t}\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\Big)^{\intercal},$ (67)
$\mathbf{K}_{{\bm{h}},t}$ $=\Big(\mathbf{L}_{t}^{-1}\,\mathbf{L}_{t}^{-\intercal}\tilde{\mathbf{H}}_{t}\,\mathbf{W}_{t-1}\,\mathbf{W}_{t-1}^{\intercal}\Big)^{\intercal}.$
From the above and Proposition D.1 the updated covariance matrix takes the form
$\boldsymbol{\Sigma}_{t|t}\\ =(\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t})\,\boldsymbol{\Sigma}_{t|t-1}\,(\mathbf{I}-\mathbf{K}_{t}\tilde{\mathbf{J}}_{t})^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\,\mathbf{K}_{t}^{\intercal},$ (68)
where
$\mathbf{I}-\mathbf{K}_{t}\,\tilde{\mathbf{J}}_{t}$ (69)
$\quad=\begin{bmatrix}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}&\boldsymbol{0}\\ \boldsymbol{0}&\mathbf{I}_{D_{{\bm{h}}}}\end{bmatrix}-\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}},t}\\ \end{bmatrix}\begin{bmatrix}\tilde{\mathbf{L}}_{t}&\tilde{\mathbf{H}}_{t}\end{bmatrix}$
$\quad=\begin{bmatrix}\mathbf{I}_{D_{{\boldsymbol{\ell}}}}&\boldsymbol{0}\\ \boldsymbol{0}&\mathbf{I}_{D_{{\bm{h}}}}\end{bmatrix}-\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{H}}_{t}\\ \mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\end{bmatrix}$
$\quad=\begin{bmatrix}\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{H}}_{t}\\ \mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{L}}_{t}&\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\end{bmatrix}$,
the predicted error covariance $\boldsymbol{\Sigma}_{t|t-1}$ is
$\boldsymbol{\Sigma}_{t|t-1}=\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}\end{bmatrix}$, (70)
and
$\mathbf{K}_{t}\,\mathbf{R}_{t}\,\mathbf{K}_{t}^{\intercal}$ $=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\\ \mathbf{K}_{{\bm{h}},t}\end{bmatrix}\,\mathbf{R}_{t}\,\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}$ (71)
$=\begin{bmatrix}\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \end{bmatrix}$.
Thus,
$\boldsymbol{\Sigma}_{t|t}$ (72)
$\quad={\tiny\begin{bmatrix}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})\boldsymbol{\Sigma}_{{\bm{h}},t|t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}}$.
Next, we consider the following surrogate covariance matrix which modifies the second block-diagonal entry
$\tilde{\boldsymbol{\Sigma}}_{t|t}=$ (73)
$\left[\begin{smallmatrix}\tiny(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny\mathbf{K}_{{\boldsymbol{\ell}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}\\ \tiny\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}&\tiny(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\mathbf{R}_{t}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}\end{smallmatrix}\right].$
It then follows that
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t},\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}},\boldsymbol{\Sigma}_{{\bm{h}}}:{\rm rank}(\boldsymbol{\Sigma}_{{\bm{h}}})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t|t}-\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{\Sigma}_{{\bm{h}}}\end{bmatrix}\right\|_{\rm F}^{2}$ (74)
$=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}},\boldsymbol{\Sigma}_{{\bm{h}}}:{\rm rank}(\boldsymbol{\Sigma}_{{\bm{h}}})=d}\|\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}-\boldsymbol{\Sigma}_{{\boldsymbol{\ell}}}\|_{\rm F}^{2}+\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\boldsymbol{\Sigma}_{{\bm{h}}}\|_{\rm F}^{2}+2\,\|\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}^{2},$
with
$\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ $=(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}(\mathbf{I}_{{D_{{\boldsymbol{\ell}}}}}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t})^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal},$ (75)
$\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}(\mathbf{I}_{{D_{{\bm{h}}}}}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t})^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\mathbf{I}_{{D_{{\bm{h}}}}}.$
Then, $\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}=\tilde{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ and $\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ is given by the first $d_{{\bm{h}}}$} principal components of $\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$.
Next, we find the update rule for the Cholesky and low-rank factors. First, the EVC for the last layer takes the form
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}$ $=\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}$ (76)
$=\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}+\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}$
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}^{\intercal/2}&\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t|t-1}^{1/2}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\end{bmatrix}$,
so that
$\hat{\boldsymbol{\Sigma}}_{{\boldsymbol{\ell}},t}^{1/2}={\cal Q}_{R}\left(\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},t-1|t-1}^{1/2}\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\boldsymbol{\ell}},t}^{\intercal}\right).$ (77)
Then, the EVC covariance for the hidden layers (lower right block-diagonal) is
$\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$ $=\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal}+\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}$ (78)
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{{\bm{h}},t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\end{bmatrix}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}$
$=\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{{\bm{h}},t}\,\mathbf{I}_{{D_{{\bm{h}}}}}.$
Finally, by Corollary C.3, the best rank-$d$ low-rank factor is
$\mathbf{C}_{t}={\cal P}_{{d_{{\bm{h}}}},+q_{{\bm{h}},t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{{\bm{h}},t}\,\tilde{\mathbf{H}}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{{\bm{h}},t}^{\intercal}\right)\in{\mathcal{M}_{({d_{{\bm{h}}}}+{D_{\bm{y}}})\times{D_{{\bm{h}}}}}(\mathbb{R})}.$ (79)
So that
$\hat{\boldsymbol{\Sigma}}_{{\bm{h}},t}=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}.$ (80)
E.4 Proof of Proposition 4.1
Proof.
We seek to bound
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}.$ (81)
Following the arguments from Proposition F.3, we obtain
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq\|\underbrace{\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm surr}}\|_{\rm F}+\|\underbrace{\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm proj}}\|_{\rm F},$ (82)
where
$\mathbf{E}_{\rm surr}=q_{{\bm{h}},t}\left(2\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}+\|\mathbf{K}_{{\bm{h}},t}\tilde{\mathbf{H}}_{t}\|_{\rm F}^{2}\right)+q_{{\boldsymbol{\ell}},t}\,\left\|\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)\,\left(\mathbf{I}-\mathbf{K}_{{\boldsymbol{\ell}},t}\,\tilde{\mathbf{L}}_{t}\right)^{\intercal}\right\|_{\rm F}$ (83)
and
$\mathbf{E}_{\rm proj}=\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\|_{\rm F}+2\,\|\mathbf{K}_{{\boldsymbol{\ell}},t}\,\mathbf{R}_{t}\,\mathbf{K}_{{\bm{h}},t}^{\intercal}\|_{\rm F}$ (84)
Given that $\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}$ is the best rank-$d$ approximation of $\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}$, it follows from a similar derivation to (107) that
$\|\tilde{\boldsymbol{\Sigma}}_{{\bm{h}},t}-\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\|_{\rm F}=\sqrt{\sum_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}\lambda_{k}^{2}},$ (85)
where $\{\lambda_{k}\}_{k={d_{{\bm{h}}}}+1}^{D_{{\bm{h}}}}$ are bottom $({D_{{\bm{h}}}}-{d_{{\bm{h}}}})$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$. ∎
Appendix F Further algorithms
F.1 LoLoFi
Algorithm 3 shows a single step of LoLoFi. The low-rank factor for last layers $\mathbf{C}_{{\boldsymbol{\ell}},0}\in{\mathcal{M}_{{d_{{\boldsymbol{\ell}}}}\times{D_{{\boldsymbol{\ell}}}}}(\mathbb{R})}$ and the low-rank factor for the hidden layers $\mathbf{C}_{{\bm{h}},0}\in{\mathcal{M}_{{d_{{\bm{h}}}}\times{D_{{\bm{h}}}}}(\mathbb{R})}$ are chosen at initialization.
[…]
5:// predict step
[…]
9:// innovation(one-step-ahead error) and Cholesky innovation variance
[…]
12:// gain for hidden layers
[…]
15:// gain for last layer
[…]
19:// mean update step
[…]
22:// low-rank updates
[…]
25:Return ${\bm{b}}_{t}=\left({\boldsymbol{\ell}}_{t|t},{\bm{h}}_{t|t},\mathbf{C}_{{\boldsymbol{\ell}},t},\mathbf{C}_{{\bm{h}},t}\right)$
Algorithm 3 Single step of LoLoFi for online learning for $t\geq 1$.
F.2 Low-rank Kalman filter (LRKF)
Inspired by [57], we derive a novel version of the low-rank Kalman filter which is (i) more computationally efficient, (ii) simple to implement, and (iii) provably optimal (in terms of Frobenius norm). We call this the LRKF method. Instead of keeping track of $\boldsymbol{\Sigma}_{t|t}$ or $\mathbf{S}_{t}$, we keep track of a low-rank factor for the covariance, and compute the Cholesky factor for the innovation variance $\mathbf{S}_{t}^{1/2}$. Algorithm 4 summarizes the predict and update steps for LRKF. Next we turn to the derivation.
We assume the model parameters $\bm{\theta}$} and the observations $\bm{y}$} follow
${\bm{\theta}}_{t}$ $={\bm{\theta}}_{t-1}+{\bm{u}}_{t},$ (86)
${\bm{y}}_{t}$ $=\mathbf{H}_{t}\,{\bm{\theta}}_{t}+{\bm{e}}_{t},$
where ${\rm Var}({\bm{e}}_{t})=\mathbf{R}_{t}$ is known and $\mathbf{H}_{t}$ is the known projection matrix. Both, ${\bm{e}}_{t}$ and ${\bm{u}}_{t}$ are zero-mean random vectors.
In the Kalman filter update equations, the so-called posterior mean and covariances are given by
$\boldsymbol{\mu}_{t}$ $=\boldsymbol{\mu}_{t-1}+\mathbf{K}_{t}\,({\bm{y}}_{t}-\hat{{\bm{y}}}_{t})\,$, (87)
$\boldsymbol{\Sigma}_{t}$ $={\rm Var}({\bm{\theta}}_{t}-\boldsymbol{\mu}_{t}),$
where $\boldsymbol{\mu}_{0}$ is given,
$\hat{{\bm{y}}}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\mu}_{t|t-1}=\mathbf{H}_{t}\,\boldsymbol{\mu}_{t-1|t-1},$ (88)
$\mathbf{K}_{t}$ $={\rm Cov}({\bm{\theta}}_{t},\boldsymbol{\epsilon}_{t}){\rm Var}(\boldsymbol{\epsilon}_{t})^{-1}=\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}\mathbf{S}_{t}^{-1},$
$\boldsymbol{\epsilon}_{t}$ $={\bm{y}}_{t}-\mathbf{H}_{t}\,\boldsymbol{\mu}_{t|t-1},$
$\mathbf{S}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t},$
$\mathbf{R}_{t}$ $={\rm Var}({\bm{e}}_{t}).$
The memory requirements of the posterior covariance is $O(D^{2})$, with $D$ number of parameters. This makes it unfeasible to store for moderately-sized neural networks.
Thus, to reduce the computational cost, we maintain the best rank $d$ approximation of the covariance matrix by
$\boldsymbol{\Sigma}_{t|t}\approx\mathbf{W}_{t}^{\intercal}\,\mathbf{W}_{t},$ (89)
where $\mathbf{W}_{t}\in\mathbb{R}^{d\times D}$ is the best rank-$d$ approximation to $\boldsymbol{\Sigma}_{t}$ (in a Frobenius norm sense). Furthermore, to maintain a numerically stable method, we work with the Cholesky of the innovation variance. In effect, this is a variant square-root low-rank Kalman filter method. We explain the details of this method below.
The predict step equations are given by
$\boldsymbol{\mu}_{t|t-1}$ $=\boldsymbol{\mu}_{t-1},$ (90)
$\boldsymbol{\Sigma}_{t|t-1}$ $=\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\mathbf{I}.$ (91)
Proposition F.1 (Kalman gain and innovations).
The variance $\mathbf{S}_{t}$ of the innovation is
$\mathbf{S}_{t}$ $=\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$ (92)
$=\mathbf{H}_{t}\,(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\mathbf{I})\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\mathbf{H}_{t}\,\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal}+q_{t}\,\mathbf{H}_{t}\,\mathbf{H}_{t}^{\intercal}+\mathbf{R}_{t}$
$=\begin{bmatrix}\mathbf{H}_{t}\,\mathbf{C}_{t}^{\intercal}&\sqrt{q_{t}}\,\mathbf{H}_{t}&\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal}\\ \sqrt{q_{t}}\mathbf{H}_{t}^{\intercal}\\ \mathbf{R}_{t}^{1/2}\end{bmatrix}$
$=\mathbf{S}_{t}^{\intercal/2}\mathbf{S}_{t}^{1/2}\,$,
with $\mathbf{S}_{t}^{1/2}$ given by
$\mathbf{S}_{t}={\cal Q}_{R}(\mathbf{C}_{t}\,\mathbf{H}_{t}^{\intercal},\,\sqrt{q_{t}}\,\mathbf{H}_{t}^{\intercal}\,,\,\mathbf{R}_{t}^{1/2}),$ (93)
where ${\cal Q}_{R}$ returns the $R$ matrix from the QR decomposition row-stacked arguments and $\mathbf{R}_{t}^{1/2}$ is the upper-triangular Cholesky decomposition of $\mathbf{R}_{t}$. The Kalman gain is given by
$\mathbf{K}_{t}$ $=\boldsymbol{\Sigma}_{t|t-1}\,\mathbf{H}_{t}^{\intercal}\,\mathbf{S}_{t}^{-1}$ (94)
$=\left(\mathbf{S}_{t}^{-1}\,\mathbf{H}_{t}\,\boldsymbol{\Sigma}_{t|t-1}\right)^{\intercal}$
$=\left(\mathbf{S}_{t}^{-1}\,\mathbf{S}_{t}^{-\intercal}\,\mathbf{H}_{t}\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{{D_{\bm{\theta}}}}\right)\right)^{\intercal}.$
Proposition F.2 (Update step).
The updated mean is
$\boldsymbol{\mu}_{t}=\boldsymbol{\mu}_{t|t-1}+\mathbf{K}_{t}({\bm{y}}_{t}-\hat{{\bm{y}}}_{t}).$ (95)
The update covariance $\boldsymbol{\Sigma}_{t|t}$, approximated through a two-step procedure: first, a surrogate covariance that ignores the cross-term effect of $\mathbf{K}_{t}\,\mathbf{H}_{t}$ on the artificial noise covariance $q_{t}$, and second, a best rank-$d$ approximation (in Frobenious norm) to the surrogate matrix takes the low-rank form
$\hat{\boldsymbol{\Sigma}}_{t|t}=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t},$ (96)
with
$\mathbf{C}_{t}={\cal P}_{d}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)\in{\mathcal{M}_{d\times{D_{\bm{\theta}}}}(\mathbb{R})},$ (97)
the low-rank (rectangular decomposition) matrix.
Proof.
The updated covariance mean follows directly from (18).
The updated covariance is
$\boldsymbol{\Sigma}_{t}$ $=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\boldsymbol{\Sigma}_{t|t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}\mathbf{K}_{t}^{\intercal}$ (98)
$=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}\right)\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}$
$=\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\left(\mathbf{C}_{t-1}^{\intercal}\,\mathbf{C}_{t-1}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}\right)\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}+\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}$
$=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}.\end{bmatrix}+q_{t}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}.$
Next, we consider the surrogate EVC matrix
$\tilde{\boldsymbol{\Sigma}}_{t}$ $=\begin{bmatrix}\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)\,\mathbf{C}_{t-1}^{\intercal}&\mathbf{K}_{t}\,\mathbf{R}_{t}^{\intercal/2}\end{bmatrix}\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}.\end{bmatrix}+q_{t}\mathbf{I}_{D_{\bm{\theta}}}$ (99)
$=\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{t}\mathbf{I}_{D_{\bm{\theta}}},$
where
$\tilde{\mathbf{C}}_{t}=\begin{bmatrix}\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal}\\ \mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\end{bmatrix}\in\mathbb{R}^{(d+o)\times D}.$ (100)
Next, the best rank-$d$ matrix is
$\hat{\boldsymbol{\Sigma}}_{t}$ $=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\boldsymbol{\Sigma}}_{t}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$ (101)
$=\operatorname*{arg\,min}_{\boldsymbol{\Sigma}:{\rm rank}(\boldsymbol{\Sigma})=d}\left\|\tilde{\mathbf{C}}_{t}^{\intercal}\,\tilde{\mathbf{C}}_{t}+q_{t}\,\mathbf{I}_{D_{\bm{\theta}}}-\boldsymbol{\Sigma}\right\|_{\rm F}^{2}$
$=\mathbf{C}_{t}^{\intercal}\,\mathbf{C}_{t},$
where
$\mathbf{C}_{t}={\cal P}_{d,+q_{t}}\left(\mathbf{C}_{t-1}\,\left(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t}\right)^{\intercal},\,\mathbf{R}_{t}^{1/2}\mathbf{K}_{t}^{\intercal}\right)\in{\mathcal{M}_{d\times{D_{\bm{\theta}}}}(\mathbb{R})}$ (102)
is the best $d$-dimensional low-rank matrix given the stacked matrices $\tilde{\mathbf{C}}_{t}$ and the dynamics covariance $q_{t}$.
Proposition F.3.
The per-step error induced by the approximation of the covariance at time $t$ is bounded by
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq q_{t}\,\left(2\,\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}+\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}^{2}\right)+\sqrt{\sum_{k=d+1}^{D_{\bm{\theta}}}\lambda_{k}^{2}},$ (103)
where $\{\lambda_{k}\}_{k=d+1}^{D_{\bm{\theta}}}$ are the bottom $({D_{\bm{\theta}}}-d)$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Proof.
We seek to bound
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F},$ (104)
where $\boldsymbol{\Sigma}_{t|t}$ is given by (98) and $\hat{\boldsymbol{\Sigma}}_{t|t}$ is given by (101). We first note that
$\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}=\left(\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}\right)+\left(\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\right),$ (105)
where $\tilde{\boldsymbol{\Sigma}}_{t|t}$ is the surrogate covariance matrix (99). Then,
$\|\boldsymbol{\Sigma}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}\leq\|\underbrace{\boldsymbol{\Sigma}-\tilde{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm surr}}\|_{\rm F}+\|\underbrace{\tilde{\boldsymbol{\Sigma}}_{t|t}-\hat{\boldsymbol{\Sigma}}_{t|t}}_{=\mathbf{E}_{\rm proj}}\|_{\rm F}$ (106)
The norm for $\mathbf{E}_{\rm proj}$ follows directly from the definition of the low-rank approximation. Let $\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}$ be the SVD decomposition of $\tilde{\boldsymbol{\Sigma}}_{t|t}$ and let $\mathbf{U}\,\mathbf{S}_{:d}\,\mathbf{V}$ be the SVD decomposition of $\hat{\boldsymbol{\Sigma}}_{t|t}$. Here, $\mathbf{S}={\rm diag}(\lambda_{1},\ldots,\lambda_{D_{\bm{\theta}}})$ and $\mathbf{S}_{:d}={\rm diag}(\lambda_{1},\ldots,\lambda_{d},0,\ldots,0)$ are the singular values (eigenvalues) of the matrices $\tilde{\boldsymbol{\Sigma}}_{t|t}$ and $\hat{\boldsymbol{\Sigma}}_{t|t}$ respectively, ordered in descending order. Then,
$\|\mathbf{E}_{\rm proj}\|_{\rm F}$ $=\|\mathbf{U}\,\mathbf{S}\,\mathbf{V}^{\intercal}-\mathbf{U}\,\mathbf{S}_{:d}\,\mathbf{V}^{\intercal}\|_{\rm F}$ (107)
$=\|\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\|_{\rm F}$
$=\sqrt{{\rm Tr}[\left(\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\right)\left(\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\right)^{\intercal}]}$
$=\sqrt{{\rm Tr}[\mathbf{U}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\mathbf{V}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{U}^{\intercal}]}$
$=\sqrt{{\rm Tr}[\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{V}^{\intercal}\mathbf{V}\left(\mathbf{S}-\mathbf{S}_{:d}\right)\mathbf{U}^{\intercal}\mathbf{U}]}$
$=\sqrt{{\rm Tr}[\left(\mathbf{S}-\mathbf{S}_{:d}\right)^{2}]}$
$=\sqrt{\sum_{k=d+1}^{D_{\bm{\theta}}}\lambda_{k}^{2}},$
where $\{\lambda_{k}\}_{k=d+1}^{D_{\bm{\theta}}}$ are the bottom $({D_{\bm{\theta}}}-d)$ eigenvalues of $\tilde{\boldsymbol{\Sigma}}_{t}$.
Next, an upper bound for the norm $\mathbf{E}_{\rm surr}$ is as follows
$\|\mathbf{E}_{\rm surr}\|_{\rm F}$ $=\|\boldsymbol{\Sigma}_{t|t}-\tilde{\boldsymbol{\Sigma}}_{t|t}\|_{\rm F}$ (108)
$=\|q_{t}\,(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t})\,(\mathbf{I}-\mathbf{K}_{t}\,\mathbf{H}_{t})-q_{t}\,\mathbf{I}\|_{\rm F}$
$=q_{t}\,\|(\mathbf{K}_{t}\,\mathbf{H}_{t})\,(\mathbf{K}_{t}\,\mathbf{H}_{t})^{\intercal}-(\mathbf{K}_{t}\,\mathbf{H}_{t})-(\mathbf{K}_{t}\,\mathbf{H}_{t})^{\intercal}\|_{\rm F}$
$\leq q_{t}\left(\|(\mathbf{K}_{t}\,\mathbf{H}_{t})\|_{\rm F}^{2}+2\,\|\mathbf{K}_{t}\,\mathbf{H}_{t}\|_{\rm F}\right).$
[…]
5:// predict step
[…]
8:// innovation and (Cholesky) innovation variance
[…]
11:// gain matrix
[…]
14:// mean and low-rank factor updates
[…]
17:Return ${\bm{b}}_{t}=(\boldsymbol{\mu}_{t},\mathbf{C}_{t})$ // updated belief
Algorithm 4 Predict and update steps in the low-rank Kalman filter (LRKF)
F.2.1 Error analysis for LRKF
Below, we analyze the single-step error incurred by LRKF without dynamics in the model parameters and scalar linear model, i.e., ${\bm{\theta}}_{t}={\bm{\theta}}_{t-1}=\ldots={\bm{\theta}}$ and $y_{t}={\bm{\theta}}^{\intercal}\,{\bm{x}}_{t}+{\bm{e}}_{t}$ with ${\rm Var}[e_{t}]=r^{2}$.
Proposition F.4.
Consider the linear model
$y_{t}={\bm{\theta}}^{\intercal}{\bm{x}}_{t}+e_{t},$
where $e_{t}$ is zero-mean with ${\rm Var}[e_{t}]=r^{2}$, and let $\boldsymbol{\Sigma}_{t-1}={\rm Var}({\bm{\theta}}-{\bm{\theta}}_{t-1|t-1})$. Let $\widehat{\boldsymbol{\Sigma}}_{t-1}$ denote the rank-$d$ covariance used by LRKF, and define
$\gamma_{t}\;=\;\min\!\bigl\{{\bm{x}}_{t}^{\intercal}\boldsymbol{\Sigma}_{t-1}{\bm{x}}_{t},\ {\bm{x}}_{t}^{\intercal}\widehat{\boldsymbol{\Sigma}}_{t-1}{\bm{x}}_{t}\bigr\}+r^{2}\quad(>0),\qquad\epsilon_{t}\;=\;y_{t}-{\bm{\theta}}_{t-1|t-1}^{\intercal}{\bm{x}}_{t}.$
Then the one-step difference between the BLUP (full-rank update) and the rank-$d$ LRKF update satisfies
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}\;\leq\;|\epsilon_{t}|\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}\left(\frac{1}{\gamma_{t}}\;+\;\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{2}}{\gamma_{t}^{2}}\right).$ (109)
Here $\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\geq\cdots\geq\sigma_{D}(\boldsymbol{\Sigma}_{t-1})$ are the singular values of $\boldsymbol{\Sigma}_{t-1}$, ${\bm{\theta}}_{t|t}$ is the BLUP, and $\hat{{\bm{\theta}}}_{t|t}$ is the rank-$d$ LRKF estimate.
In this setting, the surrogate covariance introduced by LRKF is unnecessary, and $\hat{\boldsymbol{\Sigma}}_{t}=\mathbf{C}_{t}^{\intercal}\mathbf{C}_{t}$ reduces to the best rank-$d$ approximation of the EVC matrix $\boldsymbol{\Sigma}_{t-1}={\rm Var}({\bm{\theta}}-{\bm{\theta}}_{t-1|t-1})$. As Proposition F.4 shows, a single step of LRKF deviates from the dense (i.e., full-KF) update by a factor governed entirely by the rank truncation. At time $t$, the residual $\epsilon_{t}$, the maximum eigenvalue $\sigma_{1}(\boldsymbol{\Sigma}_{t-1})$, and the feature ${\bm{x}}_{t}$ are all fixed. Consequently, the only way to tighten the bound and reduce the gap to the full update is to increase the chosen rank $d$.
Proof.
We first note that
${\bm{\theta}}_{t|t}$ $={\bm{\theta}}_{t-1|t-1}+\mathbf{K}_{t}\,\varepsilon_{t},$ (110)
$\hat{{\bm{\theta}}}_{t|t}$ $={\bm{\theta}}_{t-1|t-1}+\hat{\mathbf{K}}_{t}\,\varepsilon_{t},$
with
$\mathbf{K}_{t}=\frac{\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}}{S_{t}},$ $S_{t}={\bm{x}}_{t}^{\intercal}\,\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}+r^{2},$ (111)
$\hat{\mathbf{K}}_{t}=\frac{\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}}{\hat{S}_{t}},$ $\hat{S}_{t}={\bm{x}}_{t}^{\intercal}\,\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}+r^{2}.$
Then
$\mathbf{K}_{t}-\hat{\mathbf{K}}_{t}$ $=\frac{\boldsymbol{\Sigma}_{t-1}\,{\bm{x}}_{t}}{S_{t}}-\frac{\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}}{\hat{S}_{t}}$ (112)
$=\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}+\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}\!\left(\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right).$
Thus,
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}$ $=\|(\mathbf{K}_{t}-\hat{\mathbf{K}}_{t})\,\varepsilon_{t}\|_{2}\leq|\varepsilon_{t}|\,\Bigg(\underbrace{\left\|\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}\right\|_{2}}_{\mathrm{(I)}}+\underbrace{\left\|\hat{\boldsymbol{\Sigma}}_{t-1}\,{\bm{x}}_{t}\!\left(\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right)\right\|_{2}}_{\mathrm{(II)}}\Bigg).$ (113)
Bound for $\mathrm{(I)}$.
$\left\|\frac{(\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1})\,{\bm{x}}_{t}}{S_{t}}\right\|_{2}\leq\frac{\|\boldsymbol{\Sigma}_{t-1}-\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}}{|S_{t}|}\leq\frac{\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}}{\gamma_{t}},$
where we used Eckart–Young–Mirsky [17] and $|S_{t}|\geq\gamma_{t}:=\min\{{\bm{x}}_{t}^{\top}\boldsymbol{\Sigma}_{t-1}{\bm{x}}_{t},\ {\bm{x}}_{t}^{\top}\hat{\boldsymbol{\Sigma}}_{t-1}{\bm{x}}_{t}\}+r^{2}>0$.
Bound for $\mathrm{(II)}$.
Since
$\left|\frac{1}{S_{t}}-\frac{1}{\hat{S}_{t}}\right|=\frac{|\,\hat{S}_{t}-S_{t}\,|}{|S_{t}\,\hat{S}_{t}|}=\frac{|\,{\bm{x}}_{t}^{\top}(\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}){\bm{x}}_{t}\,|}{|S_{t}\,\hat{S}_{t}|}\leq\frac{\|{\bm{x}}_{t}\|_{2}^{2}\,\|\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}\|_{2}}{\gamma_{t}^{2}},$
we obtain
$\mathrm{(II)}\leq\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}\,\frac{\|{\bm{x}}_{t}\|_{2}^{2}\,\|\hat{\boldsymbol{\Sigma}}_{t-1}-\boldsymbol{\Sigma}_{t-1}\|_{2}}{\gamma_{t}^{2}}=\frac{\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}\,\|{\bm{x}}_{t}\|_{2}^{3}\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})}{\gamma_{t}^{2}}.$
Since $\|\hat{\boldsymbol{\Sigma}}_{t-1}\|_{2}=\sigma_{1}(\boldsymbol{\Sigma}_{t-1})$, this becomes
$\mathrm{(II)}\leq\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{3}\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})}{\gamma_{t}^{2}}.$
Plugging the bounds for $\mathrm{(I)}$ and $\mathrm{(II)}$ into (113) yields
$\|{\bm{\theta}}_{t|t}-\hat{{\bm{\theta}}}_{t|t}\|_{2}\leq|\varepsilon_{t}|\,\sigma_{d+1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}\left(\frac{1}{\gamma_{t}}+\frac{\sigma_{1}(\boldsymbol{\Sigma}_{t-1})\,\|{\bm{x}}_{t}\|_{2}^{2}}{\gamma_{t}^{2}}\right),$
as claimed. ∎
F.3 Replay-buffer variational Bayesian last layer (OnVBLL)
The VBLL method has explicit posterior predictive
$p({\bm{y}}\,|\,\boldsymbol{\ell},{\bm{h}},\boldsymbol{\Sigma},\mathbf{R})={\cal N}\left({\bm{y}}\,|\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}}+),\,\boldsymbol{\ell}^{\intercal}\,\psi({\bm{h}},{\bm{x}})\,\boldsymbol{\Sigma}\,\psi({\bm{h}},{\bm{x}})^{\intercal}\,\boldsymbol{\ell}+\mathbf{R}\right).$
Building on [25, 5], we observe that for any given $q(\bar{{\boldsymbol{\ell}}})={\cal N}(\bar{{\boldsymbol{\ell}}}\,|\,{\boldsymbol{\ell}},\boldsymbol{\Sigma})$, a lower-bound for the marginal log-likelihood of a single datapoint is given by
$\log p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},{\bm{h}},\mathbf{R})$ $\geq\mathbb{E}_{q(\bar{{\boldsymbol{\ell}}})}[\log p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\bar{{\boldsymbol{\ell}}},{\bm{h}},\mathbf{R})]$ (114)
$=\log{\cal N}\left({\bm{y}}_{t}\,|\,\boldsymbol{\ell}^{\intercal}\phi({\bm{h}},{\bm{x}}_{t}),\mathbf{R}\right)-\frac{1}{2}\phi({\bm{h}},{\bm{x}}_{t})\boldsymbol{\Sigma}\,\phi({\bm{h}},{\bm{x}}_{t})\operatorname{Tr}\left(\mathbf{R}^{-1}\right)$
$=:-{\cal L}_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma}).$
[…]
$q_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})\propto q_{t-1}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})\,\exp(-{\cal L}_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})).$ (115)
Estimation of the posterior mean (MAP filtering) involves estimating
${\boldsymbol{\ell}}_{t},{\bm{h}}_{t},\mathbf{R}_{t},\boldsymbol{\Sigma}_{t}=\operatorname*{arg\,max}_{{\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma}}q_{t}({\boldsymbol{\ell}},{\bm{h}},\mathbf{R},\boldsymbol{\Sigma})$ (116)
which can be done implicitly through adaptive optimization methods as shown in [2]. Next, to improve the performance, we consider a replay-buffer which has been shown to be much more efficient than doing fully-online SGD [1, 9, 40]. This assumption breaks with the fully-online assumption, but it is a good contender in sequential decision making problems in stationary environments.
Appendix G Additional experiments
G.1 Online classification and sequential decision making on the MNIST dataset
The results in Section 5.2 consider the LeNet5 convolutional neural network (CNN) architecture [39]. In this architecture, the last layer is $80$-dimensional and the output layer is $10$-dimensional, which corresponds to each of the $10$ possible classes. Applying HiLoFi in this setting would need the storage and update of an $800\times 800$ covariance matrix and $800$-dimensional mean vector. Sequential update of the Cholesky representation of this covariance matrix, although feasible, takes most of the computational cost in a single step of HiLoFi. Given this, here we consider LoLoFi to offset the computational cost of the layer layer. Another reason to prefer LoLoFi over HiLoFi in this setting is that the rank of the last layer is $800$-dimensional, whereas the rank for the hidden layers is $50$ dimensional. This corresponds to an overparametrized linear regression, which has been shown to have some pathologies in the offline setting [61].
As a simpler baseline, which does not model uncertainty, we consider Muon (muon) [31], which a special case of the Shampoo optimizer [24]. This is a quasi second-order optimization method that is scalable and shows strong empirical performance. We use the implementation in the Optax library [12]. Unlike the other methods, it only computes a point estimate, so we cannot use it for computing the posterior predictive. We use this method for the $\epsilon$-greedy bandit in Appendix G.1.3 and as an additional choice of optimizer in Section G.1.1.
G.1.1 Online classification on the MNIST dataset
In this experiment, we consider the problem online learning and classification on the MNIST dataset using the LeNet5 CNN. We use the optimizers detailed in Section 5.2. We take the data $\mathcal{D}_{1:T}$ with $T=60,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where ${\bm{x}}_{t}$ is a $(28\times 28\times 1)$ array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot-encoded vector such that $({\bm{y}}_{t})_{i}=1$ if ${\bm{x}}_{t}$ represents the digit $i$ and 0 otherwise. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$.
Having the BLUP over parameters ${\bm{\theta}}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$, the linearized model has mean and variance
${\bm{m}}({\bm{\theta}}_{t},{\bm{x}}_{t})$ $={\rm softmax}({f}({\bm{\theta}}_{t},{\bm{x}})),$ (117)
$\bar{{\bm{m}}}_{t}$ $={\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})+\nabla_{\bm{\theta}}{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}})({\bm{\theta}}-{\bm{m}}_{t-1}),$
$\bar{{\bm{s}}}_{t}$ $={\rm diag}({\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t}))+{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})\,{\bm{m}}({\bm{\theta}}_{t|t-1},{\bm{x}}_{t})^{\intercal}+\varepsilon\,\mathbf{I}_{D_{\bm{y}}}.$
This moment-matched linearization corresponds, in the Bayesian setting, to a Gaussian likelihood whose first and second moments match that of a Multinomial distribution. This representation was introduced in the context of online learning in [48].
In (117), ${\rm diag}({\bm{v}})$ is a function that takes as input the vector ${\bm{v}}\in\mathbb{R}^{m}$ and outputs a diagonal matrix with entries ${\rm diag}({\bm{v}})_{i,j}={\bm{v}}_{i}\,\delta(i-j)$, and the term $\varepsilon>0$ is a small constant that ensures non-zero variance.
Online learning results.
Figure 5 shows the 5000-step rolling mean of the one-step-ahead classification outcome (one for correct, zero for incorrect) using the various learning algorithms. We observe that there is no clear difference between the methods. This shows that one can tackle this complete information problem without needing to model uncertainty. This is not the case in the incomplete information case that we study in any of the experiments presented in Section 5.
Figure 5: Rolling one-step-ahead accuracy for the MNIST dataset.
Figure 6 shows the one-step-ahead accuracy of the last $10,000$ images and the running time of processing all $60,000$ images for HiLoFi, LoLoFi, LRKF, and LoFi as a function of their rank. For LoLoFi, we fix the rank of the last layer to be $50$ and vary the rank of the hidden layers.
Figure 6: Results of MNIST for online classification. Left panel: accuracy as a function of rank. Right panel: running time as a function of rank.
We observe that for low-ranks up to dimension $20$, LoFi is faster than both LoLoFi and LRKF. However, the running time of LoFi increases much more rapidly (as a function of rank) than either LoLoFi and LRKF. This is a consequence of the computational costs associated with points (1,2,3) in Appendix D.2. Next, we observe that the running time of LoLoFi is sightly above LRKF for varying rank; this is because of the double approximation of the covariance matrix: one for the hidden layers and another one for the last layer. Finally, we observe that HiLoFi is the method with highest computational cost in this experiment. This is due the high-dimensionality of the last-layer, relative to the rest of the methods.
Lastly, we observe that all methods increase their performance as a function of rank, albeit marginally. The method that most benefit from an increase in performance in HiLoFi. We hypothesize that this is due to the total rank of the hidden layer, relative to the rank of the last layer, which in this experiment is $800$.
G.1.2 Error analysis for LRKF
The top panel of Figure 7 shows the rolling mean for the two sources of approximation error in the covariance matrix for the LRKF method (detailed in Proposition F.3), as well as the rolling one-step-ahead accuracy.
Figure 7: Top: upper bound error for the surrogate error and the low-rank approximation. Bottom: $5,000$-step rolling one-step-ahead accuracy for the MNIST dataset.
We observe that, on average, both sources of error decrease over time and the accuacy of LRKF improves over time. Here, the only sources of error are the surrogate error and the low-rank error. This is because LRKF does not distinguish between last-layer and hidden-layer parameters.
G.1.3 MNIST as multi-armed bandit
Here, we present further results from Section 5.2 More precisely, we study the multi-armed bandit approach to solving the MNIST classification problem.
Choice of hyperparameters.
The hyperparameters for this experiment were chosen as follows: For Adamw and muon (used in $\varepsilon$-greedy and in conjunction with LLL), we use a learning rate of $10^{-4}$, and $\varepsilon=0.05$ For Adamw, we take 5 inner iterations and a buffer size of one and for muon, we take 1 inner iteration and a buffer size of one. Next, HiLoFi considers a rank of $50$ for the hidden parameters, we take $q_{{\bm{h}},t}=10^{-6}$ and $q_{{\boldsymbol{\ell}}}=10^{-6}$. The initial covariances $\boldsymbol{\Sigma}_{{\bm{h}},0}$ and $\boldsymbol{\Sigma}_{{\boldsymbol{\ell}},0}$ are both initialized as identity times a factor of $10^{-1}$. Similarly, LoLoFi is initialized as with HiLoFi, but with rank in the last-layer to be $100$. For LRKF, we consider a rank of $50$, $\boldsymbol{\Sigma}_{0}$ is initialized as the identity matrix, and $q_{t}=10^{-6}$. Finally, for LoFi, we also considered rank $50$; then, following the experiments in [8], we considered the initial covariance to be $a\mathbf{I}$, with $a=\exp(-8)$ (this corresponds to low-rank comprised of a matrix of zeros and diagonal terms, all set to $a$); lastly, the dynamics covariance is $q_{t}=0$.
All initial model parameters from the neural network are shared across methods and trials.
Regret analysis.
Below, we present results considering regret. Here, the reward is either $1$, if the classification is done correctly and 0 otherwise. Let $y_{t}\in\{0,1\}$ be the reward obtained at time $t$, and ${\bm{y}}_{t,a}$ be the value of arm $a$. Then the regret obtain at time $t$ is given by
$\sum_{\tau=1}^{t}(\max_{a}{\bm{y}}_{\tau,a}-y_{\tau})=t-\sum_{\tau=1}^{t}y_{\tau}.$
Figure 8 shows the average regret (across 10 trials) for HiLoFi, LoLoFi, LRKF, LoFi, and LLL. Here, we consider $\epsilon$-greedy variants for HiLoFi, LoLoFi, and LRKF. Next, TS with LLL and Adamw optimizer is denoted adamw-TS, TS with LLL and muon optimizer is denoted muon-TS. $\epsilon$-greedy with Adamw optimizer is the denoted adamw-eps, and $\epsilon$-greedy with muon optimizer is denoted muon-eps.
Figure 8: Total running time for experiments ($x$-axis) versus average regret across ten trials ($y$-axis).
Time-regret analysis.
Figure 9 shows the total running time to run the ten trials in the $x$-axis, the regret around one-standard deviation in the $y$-axis.
Figure 9: Total running time for experiments ($x$-axis) versus mean relative cumulative reward ($y$-axis).
We observe Adamw and muon are the methods with lowest performance and lowest relative cumulative reward for either $\epsilon$-greedy or TS using LLL. Next, we observe that LRKF is faster than LoLoFi, and LoLoFi is faster than HiLoFi. This result is expected and is explained in Appendix D.2. We also observe that LoFi with $\epsilon$-greedy matches LoLoFi with PS in time; however, LoLoFi has lower regret. We conjecture that this result is due to the explicit modeling of the last layer done for LoLoFi.
Finally, we observe that for all methods, $\epsilon$-greedy is faster than TS. This is because $\epsilon$-greedy only requires the evaluation of the posterior predictive mean, whereas TS samples from the posterior predictive, which in turn requires building the posterior predictive covariance shown in (10).
G.2 Online classification using LRKF on the CIFAR-10 dataset
In this Section, we study the ability of LRKF to scale to million-dimensional parameters for online classification on the CIFAR-10 dataset. We consider two VGG-style convolutional networks of different capacity. Both follow the pattern of stacked $3\times 3$ convolutions, ELU activations. We summarize the architectures below
Model Conv Blocks Dense Layers Params
VGG $64\times 2\rightarrow 128\times 2\rightarrow 256$ $256\rightarrow 256$ 1.7M
VGG+Block $64\times 2\rightarrow 128\times 2\rightarrow 256\rightarrow 512\times 2$ $256\rightarrow 256$ 4.7M
Table 3: Summary of the two VGG-style architectures used in our experiments. $C\times n$” denotes $n$ convolutional layers with $C$ channels.
We follow a similar setup as that of Appendix G.1.1. We take the data $\mathcal{D}_{1:T}$ with $T=10,000$ and $\mathcal{D}_{t}=({\bm{x}}_{t},{\bm{y}}_{t})$, where $\bm{x}$} is a $(32\times 32\times 3)$-dimensional array and ${\bm{y}}_{t}\in\{0,1\}^{10}$ a one-hot encoded vector such that ${\bm{y}}_{t}$. At every timestep $t=1,\ldots,T$ each agent is presented the image ${\bm{x}}_{t}$ which it has to classify. A predicted classification is made through the prediction ${\bm{y}}_{t|t-1}={f}(\boldsymbol{\mu}_{t|t-1},{\bm{x}}_{t})$ with $\boldsymbol{\mu}_{t|t-1}=\mathbb{E}[{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t-1}]$ and then updates its beliefs given the (true) reward ${\bm{y}}_{t}$. Updates are done as outlined in Appendix G.1.1.
Table 4 shows the one-step-ahead accuracy for LRKF for ranks $1$, 55, and $10$.
Model LRKF-1 LRKF-5 LRKF-10
VGG (1.7M) 0.3355 0.3384 0.3379
VGG+Block (4.7M) 0.3129 0.3141 0.3148
AdamW (baseline) VGG: 0.3442    VGG+Block: 0.3097
Table 4: Online classification results on CIFAR-10 using LRKF with varying ranks (1, 5, 10), compared against AdamW baselines. Lower values are better.
These results show that LRKF scales reliably to multi-million parameter networks while maintaining stable performance across different low-rank configurations. This highlights the broader applicability of our approach (including HiLoFi and LoLoFi) to large-scale architectures in vision and related domains where sequential decision making is needed.
G.3 Bandits as recommendation systems
Further results from Section 5.3
Choice of hyperparameters.
In this experiment, all agents shared a rank of $20$. The choice of hyperparameters for LLL, LoFi, VBLL, and OnVBLL were chosen following analysis of performance on the first $10,000$ observations of the datasets. In particular, we tried Bayesian optimization techniques, but found that manual tuning of hyperparameters on the first $10,000$ observations yielded higher overall performance.
For OnVBLL, we considered a buffer size of $10$, a regularization weight of $1.0$, and $50$ inner iterations. For VBLL, we considered a buffer size of $1000$, a regularization weight of $10^{-3}$, and $100$ inner iterations.
Next, for HiLoFi, we take $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0.0$ and set the initial covariances to be the identity. Next, for LRKF, we take $q_{t}=0$ and identity matrix as initial covariance.
Further results.
Figure 10 shows the daily cumulative reward for all methods. To produce this plot, we sum the cumulative rewards for all users at the end of each day and then perform a cumulative sum over the end-of-day reward.
Figure 10: Daily cumulative reward.
We observe that HiLoFi obtains the highest daily cumulative reward. The method with second highest daily cumulative reward is LRKF, closely followed by LoFi and VBLL. However, as shown in Figure 3. LRKF is more than an order of magnitude faster than VBLL. From Figure 10 and Figure 10, we observe that the top performing methods in this experiment are HiLoFi and LRKF. However, LRKF is slightly faster than HiLoFi relative to all other methods.
G.4 Bayesian optimization
Further results from Section 5.4
Choice of hyperparameters.
We fix hyperparameters across methods to ensure fairness and reproducibility. For rank-based methods, we use a rank of 50. Methods requiring buffers (e.g., LLL, OnVBLL) use a FIFO replay buffer of size 20, with 50 inner iterations per update. VBLL uses the full dataset and 100 iterations per step.
Where applicable, aleatoric uncertainty is discarded ($\mathbf{R}_{t}=0$) to isolate epistemic effects. We consider equal learning rates across optimizer-based methods (e.g., LLL, OnVBLL) at $10^{-4}$. In filtering-based methods (e.g., HiLoFi, LoFi), this learning rate corresponds to the initial hidden-layer covariance; we set the final-layer prior variance to 1, emphasising epistemic modelling in the output layer. Hyperparameters for VBLL (Wishart scale, regularization weight) follow prior work [5].
Per-step results.
Figure 11 shows the median performance and the interquartile range of the best value found by each method for all test datasets.
Figure 11: Bayesian optimization benchmark on stationary environments.
We observe that VBLL, OnVBLL, and HiLoFi are among the top performers for all datasets. Next, LLL is the least performant method for the low-dimensional datasets, but among the top performers for Nndraw50 and the top perfomer in Nndraw200.
Next, Figure 12 shows the median performance and the interquartile range of the best value found by each method for all test datasets using expected improvement.
Figure 12: Bayesian optimization benchmark on stationary environments. Query points are chosen using expected improvement.
G.4.1 NN Draw
Here, we provide further results for the performance of HiLoFi and LRKF for the Bayesian optimization problem on the DrawNN dataset.
Figure 13 reports the best value found over the course of the optimization as a function of iteration, for different ranks.
Figure 13: Solid lines show the median best value over 20 runs, and shaded regions denote the interquartile range.
We observe that increasing the rank generally improves performance, with higher ranks yielding sharper improvements early in the search that compound into stronger final results. This comes with a trade-off between computational cost and the ability to capture uncertainty more effectively, as illustrated in Table 5, which reports the final performance and running time across ranks.
Rank Time (s) Final $y_{\text{best}}$
1 26.226 656.558
10 38.4492 630.556
20 28.8612 670.913
50 30.6409 684.009
100 34.5302 782.205
110 35.1504 774.416
120 39.2540 807.298
130 38.2288 846.218
140 39.1417 848.219
150 40.7939 864.806
160 42.1907 911.116
170 44.5485 900.566
180 45.5640 864.260
190 47.1553 911.262
200 48.2127 921.062
Table 5: Performance metrics across different low-rank configurations.
We observe that both performance and running time generally increase with rank.
Finally, Figure 14 shows the best value obtained during optimization as a function of the number of observations, for different ranks using LRKF.
Figure 14: Bayesian optimization on the NNDraw dataset using LRKF. Solid lines show the median best value over 20 runs, and shaded regions denote the interquartile range.
Compared to HiLoFi, LRKF requires substantially higher ranks to achieve similar median performance. For instance, performance is comparable at rank 300 for LRKF (73s) and rank 200 for HiLoFi (under 50s). This demonstrates that explicitly modeling the last layer in full rank, as done in HiLoFi, yields a more favorable accuracy–efficiency trade-off.
G.5 In-between uncertainty
In this experiment, we test the ability of HiLoFi to capture the in-between uncertainty [21] after a single pass of the data. We consider the one-dimensional dataset introduced in [58] (Figure 1). For this experiment, we consider a four hidden-layer MLP with 128 units per layer and ELU activation function.
Result for HiLoFi.
We take $q_{{\bm{h}},t}=q_{{\boldsymbol{\ell}},t}=0$, initial covariances for last and hidden layers to be identity times $1/2$, and $\mathbf{R}_{t}=0.0$, which corresponds to having no aleatoric uncertainty in the data generating process.
Figure 15 shows the posterior predicted mean surrounded by the two-standard deviation posterior predictive for various ranks in the hidden layer.
Figure 15: In-between uncertainty of the posterior predictive induced by HiLoFi as a function of rank.
We observe that a rank of $1$ is not able to capture any uncertainty around the region with no observations; however, as we increase the rank, we observe that the posterior predictive becomes less and less confident about the true value of the mean on regions without data.
Next, Figure 16 shows the evolution of the posterior predictive mean and the posterior predictive variance as a function of the number of seen observations.
Figure 16: In-between uncertainty of the posterior predictive induced by HiLoFi as a function of seen observations.
We observe that the uncertainty around the posterior predictive mean is wide (covering the limit from 10-10 to $10$) and starts to decrease after $10$ observations. By $30$ observations, the uncertainty is narrower over regions where data has been observed, and by $120$ observations, most of the posterior predictive uncertainty is on regions where no data has been observed.
Result for LRKF
We repeat the experiment above for LRKF. Figure 17 shows the posterior predicted mean surrounded by the two-standard deviation posterior predictive for various ranks.
Figure 17: In-between uncertainty of the posterior predictive induced by LRKF as a function of rank.
In contrast to HiLoFi, we observe that LRKF requires a much higher rank to capture a reasonable level of in-between uncertainty. The panel for rank $10$ is empty because the posterior predictive (with $\mathbf{R}_{t}=0$) is psd, so it is not defined.
An important distinction between LRKF and HiLoFi is the HiLoFi considers a full-rank covariance matrix in the last layer. Despite this, LRKF with rank $200$ (which shows some notion of uncertainty) requires around $3$ times more coefficients in the covariance than HiLoFi with rank $50$ in the hidden layers.
We show the evolution of the posterior predictive mean and variance with rank $200$ of LRKF as a function of seen observations in Figure 18.
Figure 18: In-between uncertainty of the posterior predictive induced by LRKF as a function of seen observations.
We see that the evolution of the posterior predictive resembles that of HiLoFi, but at the cost of $3$ times more number of coefficients in the covariance.
Result for VBLL.
We contrast the result of HiLoFi with the offline VBLL method. For VBLL, the data is presented all at once and we perform full-batch gradient descent using Adamw with learning rate $10^{-3}$. Figure 19 shows the posterior predictive mean and two standard deviations around the posterior mean, as well as the loss curve for various epochs.
We observe that VBLL takes around $10^{4}$ epochs to converge, whereas HiLoFi takes only one. We emphasize that each epoch in VBLL considers all datapoints at once, so the gradient update is more informative, whereas for HiLoFi, the method only gets to observe the data once and in a sequence. This highlights the efficiency of our method to produce approximate posterior predictives that can be used online.
Figure 19: In-between uncertainty of the posterior predictive induced by VBLL as a function of number of epochs.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The frequentist and Bayesian approaches to filtering are introduced in Section 3.2 and detailed in Appendix B. Our method and its assumptions are detailed in Section 4. Mathematical proofs of our claism are collected in Appendix E. Experiments and comparisons to other methods validating the experimental claims are provided in Section 5. The motivation and contribution for our method is provided in Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations of the method are discussed in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The main assumptions of the methods we introduced are detailed in Section 3.2 and Appendix B. The proof of Proposition D.1 is in Appendix E.1, the proof of Proposition D.2 is in Appendix E.2, the proof of Proposition D.3 is in Appendix E.3, and the proof of Proposition 4.1 is in Appendix E.4. Further propositions and proofs for the LRKF are stated and proved in Appendix F.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: To facilitate reproducibility of the methods we present, Algorithm 4 details LRKF, Algorithm 2 details HiLoFi, and Algorithm 3 details LoLoFi. The experiments are detailed in Section 5, which mentions the dataset and the neural architecture used. Further information (such as hyperparameter selection) for all experiments is detailed in Appendix G. Furthermore, we provide code to reproduce the results in the anonymized link in the introduction.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All our datasets are open source and available at the mentioned sources in Section 5. Preprocessing of the data is detailed in Section 5 and Appendix G. Detailed pseudocode for the methods we introduce are detailed in Algorithm 4 for LRKF, Algorithm 2 for HiLoFi, and Algorithm 3 for LoLoFi. Furthermore, we provide code to reproduce the results in the anonymized link in the introduction.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The training and test details are in Section 5 and Appendix G.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Error bars are reported and explained for all experiments in Section 5 and Appendix G. For the experiment in Section 5.2, suitable error bars are deferred to Figure 9 in Appendix G.1.3 as including them in the main figure would have made it overly cluttered and difficult to interpret.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Compute workers are detailed in Section 5. An estimate of the running time to replicate the results with the compute used are provided in Section 5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have reviewed the NeurIPS Code of Ethics, and the research conducted in the paper conforms, in every respect, with these guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Potential societal impacts are acknowledged in Section 7.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The methods developed in this paper are general-purpose tools for online learning and sequential decision-making. We do not believe they pose a high risk of misuse, and therefore no specific safeguards were necessary.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All assets used, whether code, data, or models have been properly cited in the main text.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Assets are detailed in the introduction of the main text. We include all details to reproduce our methods and the licence is provided within the code. Consent was not required.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper did not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper did not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.